### v1

# On complexity of restricted fragments of Decision DNNF

Abstract
Decision dnnf (a.k.a. $\wedge_{d}$-fbdd) is an important special case of Decomposable Negation Normal Form (dnnf), a landmark knowledge compilation model. Like other known dnnf restrictions, Decision dnnf admits fpt sized representation of cnfs of bounded primal treewidth. However, unlike other restrictions, the complexity of representation for cnfs of bounded incidence treewidth is wide open.
In [3], we resolved this question for two restricted classes of Decision dnnf that we name $\wedge_{d}$-obdd and Structured Decision dnnf. In particular, we demonstrated that, while both these classes have fpt-sized representations for cnfs of bounded primal treewidth, they need xp-size for representation of cnfs of bounded incidence treewidth.
In the main part of this paper we carry out an in-depth study of the $\wedge_{d}$-obdd model. We formulate a generic methodology for proving lower bounds for the model. Using this methodology, we reestablish the xp lower bound provided in [3]. We also provide exponential separations between fbdd and $\wedge_{d}$-obdd and between $\wedge_{d}$-obdd and an ordinary obdd. The last separation is somewhat surprising since $\wedge_{d}$-fbdd can be quasipolynomially simulated by fbdd.
In the remaining part of the paper, we introduce a relaxed version of Structured Decision dnnf that we name Structured $\wedge_{d}$-fbdd. In particular, we explain why Decision dnnf is equivalent to $\wedge_{d}$-fbdd but their structured versions are distinct.) We demonstrate that this model is quite powerful for cnfs of bounded incidence treewidth: it has an fpt representation for cnfs that can be turned into ones of bounded primal treewidth by removal of a constant number of clauses (while for both $\wedge_{d}$-obdd and Structured Decision dnnf an xp lower bound is triggered by just two long clauses).
We conclude the paper with an extensive list of open questions.
1 Introduction
1.1 The main objective
Decomposable negation normal forms (dnnfs) [5] is a model for representation of Boolean functions that is of a landmark importance in the area of (propositional) Knowledge compilation (kc). The difference of the model from deMorgan circuits is that all the $\wedge$ gates are decomposable meaning, informally, that two distinct inputs of the gate cannot be reached by the same variable. We will refer to such gates as $\wedge_{d}$. Another way to define dnnfs is as a Non-deterministic read-once branching program ($1$-nbp) equipped with $\wedge_{d}$ gates. This latter view is useful for understanding that (from the univariate perspective, that is when the size bounds depend on the number $n$ of input variables) dnnfs are not very strong compared to $1$-nbps since the latter can quasipolynomially simulate the former [14]. In particular, a class of Boolean functions requiring exponentially large $1$-nbps for its presentation requires an exponential lower bound for representation by dnnfs.
The strength of $\wedge_{d}$ becomes distinctive when we consider dnnfs from the multivariate perspective. In particular, cnfs of primal treewidth at most $k$ have fpt presentation as a dnnf [5] but, in general requires an xp-sized representation as a $1$-nbp. [15]. The fpt upper bound as above turned out to be robust in the sense that it is preserved by all known restrictions of dnnfs that retain $\wedge_{d}$-gates. In light of this, it is natural to investigate if fpt upper bounds hold for the incidence treewidth.
Among all known dnnf restrictions using $\wedge_{d}$ gates, there are two that are minimal in the sense that every other restriction is a generalization of one of them. One such minimal restriction is a Sentential Decision Diagram [6] that admits fpt presentation for a much more general class that cnfs of bounded incidence treewidth [2]. The other restriction is Decision dnnf [8, 11] where the parameterized complexity of representation of cnfs of bounded incidence treewidth is an open question formally stated below.
Open question 1.
Are there a function $f$ and a constant $a$ so that a cnf of incidence treewidth at most $k$ can be represented as a Decision dnnf of size at most $f(k)\cdot n^{a}$?
To understand the reason of difficulty of Open Question 1, it is convenient to view Decision dnnfs as Free Binary Decision Diagrams (fbdds) equipped with $\wedge_{d}$ gates [1]. Throughout this paper, we will refer to this equivalent representation as $\wedge_{d}$-fbdd. This representation restricts the role of $\vee$ gates only to the representation of Shannon expansion. That is each $\vee$ gate is associated with a variable and represents a function $(x\wedge f)\vee(\neg x\wedge g)$ where $f$ and $g$ are some Boolean functions not depending on $x$. This restriction enables an easy guess of the variables of variables but it is not clear how to handle the guessing of satisfaction/non-satisfaction of clauses.
The main objective of the research presented in this paper is to make a progress towards resolution of Open Question 1. Before presenting our results, we state reasons why the question is important to resolve.
1.2 Motivation
It has been observed in [8] that the trace of several well known dpll based model counters (##\##-sat solvers) using component analysis can be represented as a Decision dnnf. It is well known that the number of satisfying assignments of a cnf of bounded incidence treewidth can be efficiently computed [17]. On the other hand, it is not clear whether the efficient computation can be carried out by the dppl solvers. Resolving Open Question 1 will provide an important insight in this direction.
Another aspect of the motivation is the relationship between $\wedge_{d}$-fbdd and proof complexity. In particular, there is a close similarity between $\wedge_{d}$-fbdd and regular resolution as a well known fbdd-based representation of regular resolution implicitly uses decomposable conjunction. Indeed, to prove unsatisfiability of two variable disjoint cnfs, it is enough to prove unsatisfiability of just one of them and to completely ignore the other. Importantly, it is an open question whether unsatisfiable cnfs of bounded incidence treewidth have fpt-sized regular resolution refutation. In the context of the similarity as specified above, it is reasonable to expect that understanding the power of $\wedge_{d}$-fbdd as related to cnfs of bounded incidence treewidth would provide an insight for such a power of regular resolution. This connection is especially intriguing as the question for regular resolution is a restriction of the same question for the full-fledged resolution, a well known open problem in the area of proof complexity [4]
1.3 Our results and structure of the paper
In this paper we tackle the above question by restricting it to special cases of $\wedge_{d}$-fbdd. We consider two restrictions: obliviosness and structuredness.
Arguably, the best known restricted class of fbdd is Ordered Binary Decision Diagram (obdd). Applying this restriction to $\wedge_{d}$-fbdds, we obtain $\wedge_{d}$-obdds or, to put it differently, obdds equipped with $\wedge_{d}$ gates. Arguing as in [8], we conclude that $\wedge_{d}$-obdd represents the trace of a dblp model counter with component analysis and fixed variable ordering
To the best of our knowledge the $\wedge_{d}$-obdd was first introduced in [3]. In [10] the model appears under the name $OBDD[\wedge]_{\prec}$ where $\prec$ is an order over a chain.
While $\wedge_{d}$-obdd are known to have fpt representation [12] (as noted in [3] the constructiom in [12] is, in fact, structured $\wedge_{d}$-obdd), it is shown in [3] that representation of cnfs of bounded incidence treewidth requires an xp-sized representation as $\wedge_{d}$-obdds.
In this paper, we restate the proof of [3] by explicitly stating a generic approach to proving lower bounds of $\wedge_{d}$-obdd as a generalization of such an approach for obdd. Using this generic approach we demonstrate an exponential separation between fbdd and $\wedge_{d}$-obdd. In particular, we introduce a class of cnfs representable by poly-sized fbdd and prove an exponential lower bound for representation of this class of cnfs as $\wedge_{d}$-obdds.
Our last result for $\wedge_{d}$-obdds is an exponential separation between $\wedge_{d}$=obdds and obdds. In the context of sat solvers this separation means that the use of component analysis in presence of a fixed variable ordering can lead to an exponential runtime improvement. This situation contrasts with the dynamic variable ordering because of a well known quasipolynomial simulation of $\wedge_{d}$-textscfbdd by fbdd [1].
As stated above, the second restriction on $\wedge_{d}$-fbdds considered in this paper is structuredness. In the context of cnfs of bounded incidence treewidth structured Decision dnff is a rather restrictive model. Indeed, it is shown in [3] that taking a monotone $2$-cnf of bounded primal treewidth and adding one clause with negative literals of all the variables requires structured Decision dnnfs representing such cnfs to be xp-sized in terms of their incidence treewidth. By weakening the structurality restriction, we introduce a more powerful model that we name structured $\wedge_{d}$-fbdds.
The starting point of the construction is that transformation from Decision dnnf into $\wedge_{d}$-fbbd leads to an encapsulation effect: a $\vee$ node and two $\wedge_{d}$ nodes implementing Shannon decomposition are replaced by a single decision node. The structured $\wedge_{d}$-fbdd requires that only the $\wedge_{d}$ nodes explicitly present in the model respect a vtree or to put it differently, no constraint is impossed on the $\wedge_{d}$ nodes ’hidden’ inside the decision nodes.
We demonstrate that structured $\wedge_{d}$-fbdd is a rather strong model. In particular, unlike $\wedge_{d}$-obdd and Structured Decision dnnf, it has fpt-sized representation for cnfs that can be turned into cnfs of bounded primal treewidth by removal of a constant number of clauses. We leave it open as to whether Structured $\wedge_{d}$-fbdds admit an fpt representation of cnfs of bounded incidence treewidth.
In the rest of the paper, Section 2 introduces the necessary technical background, Sections 3 and 4 present the results related to $\wedge_{d}$-obdd and the structured restrictions of Decision dnnf. Proofs of selected statements are postponed to the Appendix.
2 Preliminaries
2.1 Assignments, rectangles, Boolean functions
Let $X$ be a finite set of variables. A (truth) assignment on $X$ is a mapping from $X$ to $\{0,1\}$. The set of truth assignments on $X$ is denoted by $\{0,1\}^{X}$. We reserve lowercase boldface English letters to denote the assignments.
In this paper we will consider several objects for which the set of variables is naturally defined: assignments, Boolean functions, cnfs, and several knowledge compilation models. We use the unified notation $\mathsf{var}(\mathcal{O})$ where $\mathcal{O}$ is the considered object. In particular, the set of variables (that is, the domain) of an assignment ${\bf a}$ is denoted by $\mathsf{var}({\bf a})$.
We observe that if ${\bf a}$ and ${\bf b}$ are assignments and for each $x\in\mathsf{var}({\bf a})\cap\mathsf{var}({\bf b})$, ${\bf a}(x)={\bf b}(x)$ then ${\bf a}\cup{\bf b}$ is an assignment to $\mathsf{var}({\bf a})\cup\mathsf{var}({\bf b})$.
Let $\mathcal{H}$ be a set of assignments, not necessarily over the same set of variables. Let $\mathsf{var}(\mathcal{H})=\bigcup_{{\bf a}\in\mathcal{H}}\mathsf{var}({\bf a})$. We say that $\mathcal{H}$ is uniform if for each ${\bf a}\in\mathcal{H}$, $\mathsf{var}({\bf a})=\mathsf{var}(\mathcal{H})$. For example the set $\mathcal{H}_{1}=\{\{(x_{1},0),(x_{2},1)\}\{(x_{2},0),(x_{3},1)\}\}$ is not uniform and $\mathsf{var}(\mathcal{H}_{1})=\{x_{1},x_{2},x_{3}\}$. On the other hand, $\mathcal{H}_{2}=\{\{(x_{0},1),(x_{2},0),(x_{3},1)\},\\ \{(x_{1},1),(x_{2},1),(x_{3},0)\}\}$ is uniform but, again, $\mathcal{H}_{2}=\{x_{1},x_{2},x_{3}\}$.
We are now going to introduce an operation that is, conceptually, very similar to the Cartesian product. We thus slightly abuse the notation by referring to this operation as the Cartesian product. So, let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be sets of assignments with $\mathsf{var}(\mathcal{H}_{1})\cap\mathsf{var}\mathcal{H}_{2})=\emptyset$. We denote $\{{\bf a}\cup{\bf b}|{\bf a}\in\mathcal{H}_{1},{\bf b}\in\mathcal{H}_{2}\}$ by $\mathcal{H}_{1}\times\mathcal{H}_{2}$. For example, suppose that $\mathcal{H}_{1}=\{\{(x_{1},1),(x_{2},0)\},\{(x_{2},1)\}\}$ and $\mathcal{H}_{2}=\{\{(x_{3},1)\},\{(x_{3},0),(x_{4},0)\}\}$. Then $\mathcal{H}_{1}\times\mathcal{H}_{2}=\{\{(x_{1},1),(x_{2},0),(x_{3},1)\},\\ \{(x_{1},1),(x_{2},0),(x_{3},0),(x_{4},0)\},\{(x_{2},1),(x_{3},1)\},\{(x_{2},1),(x_{3},0),(x_{4},0)\}\}$. We note that for any set $\mathcal{H}$ of assignments, $\mathcal{H}\times\{\emptyset\}=\mathcal{H}$.
If both $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are uniform and both $\mathsf{var}(\mathcal{H}_{1})$ and $\mathsf{var}(\mathcal{H}_{2})$ are non-empty then $\mathcal{H}_{1}\times\mathcal{H}_{2}$ is called a rectangle. More generally, we say that a uniform set $\mathcal{H}$ of assignments is a rectangle if there are uniform sets $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ of assignments such that $\mathsf{var}(\mathcal{H}_{1})$ and $\mathsf{var}(\mathcal{H}_{2})$ partition $\mathsf{var}(\mathcal{H})$ such that $\mathcal{H}=\mathcal{H}_{1}\times\mathcal{H}_{2}$. We say that $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are witnesses for $\mathcal{H}$. Note that a rectangle can have many possible pairs of witnesses, consider, for example $\{0,1\}^{X}$.
The following notion is very important for proving our results.
Definition 1.
Let $\mathcal{H}$ be a rectangle and let $Y\subseteq\mathsf{var}(\mathcal{H})$. We say that $\mathcal{H}$ breaks $Y$ if there is a pair $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ of witnesses of $\mathcal{H}$ such that both $Y\cap\mathsf{var}(\mathcal{H}_{1})$ and $Y\cap var(\mathcal{H}_{2})$ are nonempty.
For example $\{0,1\}^{X}$ breaks any disjoint non-empty subset $Y$ of $X$ of size at least $2$. On the other hand, let $\mathcal{H}_{1}=\{\{(x_{1},1),(x_{2},0)\}\}$ and let $\mathcal{H}_{2}=\{\{(x_{3},0),(x_{4},0)\},\{(x_{3},0),(x_{4},1)\},\{(x_{3},1),(x_{4},0)\}\}$. Then $\mathcal{H}=\mathcal{H}_{1}\times\mathcal{H}_{2}$ does not break $\{x_{3},x_{4}\}$.
We can naturally extend $\mathcal{H}_{1}\times\mathcal{H}_{2}$ to the case of more than two sets of assignments, that is $\mathcal{H}_{1}\times\dots\times\mathcal{H}_{m}$ as e.g. $(\mathcal{H}_{1}\times\dots\times\mathcal{H}_{m-1})\times\mathcal{H}_{m}$. The operation is clearly commutative that is we can permute the elements in any order. We also use the notation $\prod_{i\in[m]}\mathcal{H}_{i}$ (recall that $[m]=\{1,\dots,m\}$) We notice that if $I_{1},I_{2}$ is a partition of $[m]$ then $\prod_{i\in[m]}\mathcal{H}_{i}=\prod_{i\in I_{1}}\mathcal{H}_{i}\times\prod_{i\in I_{2}}\mathcal{H}_{i}$. Therefore, we observe the following.
Proposition 1.
Suppose that $\mathcal{H}=\prod_{i\in[m]}\mathcal{H}_{i}$, $m\geq 2$ and $\mathcal{H}_{i}$ is uniform for each $i\in[m]$. Further on, assume that $\mathcal{H}$ is not a rectangle breaking a set $Y\subseteq\mathsf{var}(\mathcal{H})$ Then there is $i\in[m]$ such that $Y\subseteq\mathsf{var}(\mathcal{H}_{i})$.
Next, we define the notions of projection and restriction.
Definition 2.
  1. 1.
[…]
For example, if =$\mathcal{H}=\{\{(x_{1},0),(x_{2},0),(x_{3},1)\},\{(x_{1},1),(x_{2},0),(x_{3},1)\},\{(x_{1},1),(x_{2},1),(x_{3},0\}\}$ and ${\bf a}=\{(x_{1},1)\}$ then $\mathcal{H}|_{\bf a}=\{\{(x_{2},0),(x_{3},1)\},\{(x_{2},1),(x_{3},0)\}\}$.
Observe that if $\mathcal{H}$ is uniform then $\mathcal{H}_{\bf a}$ is also uniform.
A literal is a variable $x$ or its negation $\neg x$. For a literal $\ell$, we denote its variable by $\mathsf{var}(\ell)$. Naturally for a set $L$ of literals $\mathsf{var}(L)=\{var(\ell)|\ell\in L\}$. $L$ is well formed if for every two distinct literals $\ell_{1},\ell_{2}\in L$, $\mathsf{var}(\ell_{1})\neq\mathsf{var}(\ell_{2})$.
If $x\in\mathsf{var}(L)$, we say that $x$ occurs in $L$. Moreover, $x$ occurs positively in $L$ if $x\in L$ and negatively if $\neg x\in L$.
There is a natural correspondence between assignments and well-formed sets of literals. In particular, an assignment ${\bf a}$ corresponds to a set $L$ of literals over $\mathsf{var}(a)$ such that, for each $x\in\mathsf{var}({\bf a})$, $x\in L$ if ${\bf a}(x)=1$, otherwise $\neg x\in L$. We will use this correspondence to define satisfying assignments for cnfs.
A Boolean function $f$ on variables $X$ is a mapping from $\{0,1\}^{X}$ to $\{0,1\}$. As mentioned above, we denote $X$ by $\mathsf{var}(f)$. Let ${\bf a}$ be an assignment on $\mathsf{var}(f)$ such that $f({\bf a})=1$. We say that ${\bf a}$ is a satisfying assignment for $f$.
We denote by $\mathcal{S}(f)$ the set of satisfying assignments for $f$. Conversely, a uniform set $\mathcal{S}$ naturally gives rise to a function $f$ over $\mathsf{var}(f)$ whose set of satisfying assignments is precisely $\mathcal{S}$. Finally, the set of satisfying assignments naturally leads to the definition of a restriction of a function. In particular, for a function $f$ and an assignment ${\bf a}$, $f|_{\bf a}$ is a function on $\mathsf{var}(f)\setminus\mathsf{var}({\bf a})$ whose set of satisfying assignments is precisely $\mathcal{S}(f)|_{\bf a}$.
2.2 Models of Boolean functions
We assume the reader to be familiar with the notion of Boolean formulas and in particular, with the notion of conjunctive normal forms (cnfs). For a cnf $\varphi$, $\mathsf{var}(\varphi)$ is the set of variables whose literals occur in the clauses of $\varphi$. For example, let $\varphi=(x_{1}\vee x_{2})\wedge(x_{2}\vee\neg x_{3}\vee\neg x_{4})$. $\mathsf{var}(\varphi)=\{x_{1},x_{2},x_{3},x_{4}\}$.
In order to define the function represented by a cnf, it will be convenient to consider a cnf as a set of its clauses and a clause as a set of literals. Let ${\bf a}$ be an assignment and let $L_{\bf a}$ be the well-formed set of literals over $\mathsf{var}({\bf a})$ corresponding to ${\bf a}$. Let $\varphi=\{C_{1},\dots,C_{q}\}$ be a cnf. Then ${\bf a}$ satisfies a clause $C_{i}$ if $L_{\bf a}\cap C_{i}\neq\emptyset$ and ${\bf a}$ satisfies $\varphi$ if ${\bf a}$ satisfies each clause of $\varphi$. The function $f(\varphi)$ represented by $\varphi$ has the set of variables $\mathsf{var}(\varphi)$ and $\mathcal{S}(f(\varphi))$ consists of exactly those assignments that satisfy $\varphi$. In what follows, we slightly abuse the notation and refer to the function represented by $\varphi$ also by $\varphi$. The correct use will always be clear from the context.
We are next going to define the model Decision dnnf, which throughout this paper is referred to as $\wedge_{d}$-fbdd (the reasons will become clear later). The model is based on a directed acyclic graph (dag) with a single source with vertices and some edges possibly being labelled. We will often use the following notation. Let $D$ be a dag and let $u$ be a vertex of $D$. Then we denote by $D_{u}$ the subgraph of $D$ induced by $u$ and all the vertices reachable from $u$ in $D$. The labels on vertices and edges of $D_{u}$ are the same as the labels of the respective vertices and edges of $D$.
We define $\wedge_{d}$-fbdd in the following two definitions. The first one will define $\wedge_{d}$-fbdd as a combinatorial object. In the second definition, we will define the semantics of the object that is the function represented by it.
Definition 3.
A $\wedge_{d}$-fbdd is a dag $B$ with a single source and two sinks. One sink is labelled by ${\bf 0}$, the other by ${\bf 1}$. Each non-sink node has two out-neighbours (children). A non-sink node may be either a conjunction node, labelled with $\wedge$ or a decision node labelled with a variable name. If $u$ is a decision node then one outgoing edge is labelled with 00, the other with $1$.
Let $u$ be a node of $B$. Then $\mathsf{var}(B_{u})$ is the set of variables labelling the decision nodes of $B_{u}$. $B$ obeys two constraints: decomposability and read-onceness. The decomposability constraint means that all the conjunction nodes $u$ are decomposable. In particular, let $u_{0}$ and $u_{1}$ be the children of $u$. Then $\mathsf{var}(B_{u_{0}})\cap\mathsf{var}(B_{u_{1}})=\emptyset$. The read-onceness means that two decision nodes labelled by the same variable cannot be connected by a directed path.
We define the size of $B$ as the number of nodes and refer to it as $|B|$.
We observe that if $B$ is a $\wedge_{d}$-fbdd and $u\in V(B)$ then $B_{u}$ is also a $\wedge_{d}$-fbdd. This allows us to introduce recursive definition from the sinks up to the source of $B$.
Definition 4.
Let $B$ be a $\wedge_{d}$-fbdd. We define the set $\mathcal{A}(B)$ of assignments accepted by $B$ in the following recursive way. If $B$ consists of a sink node only then $\mathcal{A}(B)=\{\emptyset\}$ if the sink is labelled by ${\bf 1}$ and $\emptyset$ otherwise.
Otherwise, let $u$ be the source of $B$, let $u_{0}$ and $u_{1}$ be the children of $u$, and let $\mathcal{A}_{i}=\mathcal{A}(B_{u_{i}})$ for each $i\in\{1,2\}$.
Assume first that $u$ is a decision node and let $x$ be the variable labelling $u$. We assume w.l.o.g that $(u,u_{0})$ is labelled by 00 and $(u,u_{1})$ is labelled by $1$. Then $\mathcal{A}(B)=\mathcal{A}_{0}\times\{(x,0)\}\cup\mathcal{A}_{1}\times\{(x,1)\}$. If $u$ is a conjunction node then $\mathcal{A}(B)=\mathcal{A}_{0}\times\mathcal{A}_{1}$.
The function $f(B)$ represented by $B$ is the function over $var(B)$ whose satisfying assignments are all ${\bf b}$ such that there is ${\bf a}\in\mathcal{A}(B)$ such that ${\bf a}\subseteq{\bf b}$. We denote $\mathcal{S}(f(B))$ by $\mathcal{S}(B)$ and refer to this set as the set of satisfying assignments of $B$.
In read-once models, directed paths are naturally associated with assignments. A formal definition is provided below.
Definition 5.
Let $P$ be a directed path of a $\wedge_{d}$-fbdd. We denote by $\mathsf{var}(P)$ the set of variables labelling the nodes of $P$ but the last one. We denote by ${\bf a}(P)$ the assignment over $\mathsf{var}(P)$ where for each $x\in\mathsf{var}(P)$, ${\bf a}(x)$ is determined as follows. Let $u$ be the node of $P$ labelled by $x$. Then ${\bf a}(x)$ is the label on the outgoing edge of $u$ in $P$. (Note that even if the last node of $P$ is labelled by a variable, the variable is not included into $\mathsf{var}(P)$ exactly because the outgoing edge of the last node is not in $P$ and hence the assignment of the variable cannot be defined by the above procedure.)
We next define subclasses of $\wedge_{d}$-fbdds. Let $B$ be a $\wedge_{d}$-fbdd that does not have $\wedge_{d}$ nodes. Then $B$ is a fbdd. Next, let $B$ be a $\wedge_{d}$-fbdd such that there is a linear order $\pi$ over a superset of $\mathsf{var}(B)$ so that the following holds. Let $P$ be a directed path of $B$ and suppose that $P$ has two decision nodes $u_{1}$ and $u_{2}$ labelled by variables $x_{1}$ and $x_{2}$ and assume that $u_{1}$ occurs before $u_{2}$ in $P$. Then $x_{1}<_{\pi}x_{2}$ (that is $x_{1}$ is smaller than $x_{2}$ according to $\pi$). Then $B$ is called an $\wedge_{d}$-obdd (obeying $\pi$ if not clear from the context). Finally, an $\wedge_{d}$-obdd without $\wedge_{d}$ nodes is just an obdd. Allowing $\pi$ to be defined over a superset of $\mathsf{var}(B)$ will be convenient for our reasoning when we consider subgraphs of a $\wedge_{d}$-obdd.
We will need one more restricted class of $\wedge_{d}$-fbdd, a structured one. We will define it in Section 4 where it will actually be used.
Throughout the paper, when we refer to a smallest model representing a function $f$, we will use the notation $modelname(f)$. For instance $\wedge_{d}$-obdd$(f)$, $(f)$, obdd$(f)$ respectively refer to the smallest $\wedge_{d}$-obdd, fbdd($f$), and obdd$(f)$ representing $f$.
2.3 Graphs, width parameters, and correspondence with cnf
We use the standard terminology related to graph as specified in [7]. Let us recall that a matching $M$ of a graph $G$ is a subset of its edges such that for any distinct $e_{1},e_{2}\in M$, $e_{1}\cap e_{2}=\emptyset$. The treewidth of $G$, denoted by $tw(G)$ is arguably the best known graph parameter measuring closeness to a tree. To define $tw(G)$, we need the notion of a tree decomposition of $G$ which is a pair $(T,{\bf B})$ where $T$ is a tree and ${\bf B}:V(T)\rightarrow 2^{V(G)}$. The set ${\bf B}(t)$ for $t\in V(T)$ is called the bag of $T$. The bags obey the rules of containement meaning that for each $e\in V(G)$ there is $t\in V(T)$ such that $e\subseteq{\bf B}(t)$ and connectivity. For the latter, let $T[v]$ be the subgraph of $T$ induced by $t$ such that $v\in{\bf B}(t)$. Then $T[v]$ is connected. For $t\in V(T)$, the width of ${\bf B}(t)$ is $|{\bf B}(t)|-1$ The width of $(T,{\bf B})$ is the largest width of a bag. Finally, $tw(G)$ is the smallest width of a tree decomposition. The pathwidth of $G$, denoted by $pw(G)$, is the smallest width of a tree decomposition $(T,{\bf B})$ of $G$ such that $T$ is a path.
A central notion used throughout this paper is a matching crossing a sequence of vertices. This notion is formally defined below.
Definition 6.
Let $G$ be a graph. Let $M$ be a matching of $G$ and let $U,V$ be a partition of $V(M)$ such that each edge of $M$ has one end in $U$, the other in $V$. We then say that $M$ is a matching between $U$ and $V$. Let $\pi$ be a linear order of $V(G)$. A matching $M$ of $G$ crosses $\pi$ if there is a prefix $\pi_{0}$ of $\pi$ with the corresponding suffix $\pi_{1}$ so that $M$ s a matching between $\pi_{0}\cap V(M),\pi_{1}\cap V(M)$ We call $\pi_{0}$ a witnessing prefix for $M$.
Remark 1.
In Definition 6, we overloaded the notation by using a prefix of $\pi$ both as a linear order and as a set. The correct use will always be clear from the context.
Example 1.
Let $G=P_{8}$ (a path of $8$ vertices) with $V(G)=\{v_{1},\dots,v_{8}\}$ appearing along the path in the order listed. Consider an order $\pi=(v_{1},\dots,v_{8})$. Then the only matchings crossing $\pi$ are sing leton ones. On the other hand, consider the order $\pi^{*}=(v_{1},v_{3},v_{5},v_{7},v_{2},v_{4},v_{6},v_{8})$. Then the matching $M=\{\{v_{1},v_{2}\},\{v_{3},v_{4}\},\{v_{5},v_{6}\},\{v_{7},v_{8}\}\}$ crosses $\pi^{*}$ as witnessed by $(v_{1},v_{3},v_{5},v_{7})$.
Taking to the maximum the size of a matching crossing the given linear order leads us to the following width parameters.
Definition 7.
Let $G$ be a graph and $\pi$ be linear order of $V(G)$. The linear maximum matching width lmm-width of $\pi$ is the largest size of a matching of $G$ crossing $\pi$. The lmm-width of $G$, denoted by $lmmw(G)$ is the smallest linear matching width over all linear orders of $\pi$.
The linear special induced matching width (lsim-width) of $\pi$ is the largest size of an induced matching of $G$ crossing $\pi$. Accordingly, lsim-width of $G$, denoted by $lsimw(G)$, is the smallest lsim-width of a linear order of $G$.
The parameters lmm-width and lsim-width are linear versions of well known graph parameters maximum matching width [19] and special induced matching width [9].
Lemma 1.
There is a constant $c$ such that for each sufficiently large $k$ there is an infinite class ${\bf G}_{k}$ of graphs such that for each $G\in{\bf G}_{k}$, $tw(G)\leq k$ and $lsimw(G)\geq c\log(|V(G)|)\cdot k$.
Proof.
It is known [15] that the statement as in the lemma holds if $lsimw(G)$ is replaced by $pw(G)$ and, moreover, the considered class of graphs has max-degree $5$. Next, it is known [16] that the $pw(G)$ and $lmmw(G)$are linearly related. This means that the lemma is true if $lsimw(G)$ is replaced by $lmmw(G)$. Finally, for any matching $M$,there is an induced subset $M^{\prime}$ of size at least $|M|/(2d+1)$ where $d$ is the max-degree of $G$. Indeed, greedily choose edges $e\in M$ into $M^{\prime}$ by removing the edges of $M$ connecting $N(e)$ (at most $2d$ edges are removed per chosen edge). Since $d=5$ for the considered class, $lsimw(G)\geq lmmw(G)/11$. ∎
For a cnf $\varphi$, we use two graphs, both regard $\varphi$ as a hypergraph over $\mathsf{var}(\varphi)$ ignoring the literals. The primal graph of $\varphi$ has $\mathsf{var}(\varphi)$ as the set of vertices and the two vertices are adjacent if they occur in the same clause. The incidence graph of $\varphi$ has $\mathsf{var}(\varphi)$ and the clauses as its vertices. A variable $x$ is adjacent to a clause $C$ if and only if $x$ occurs in $C$. We denote by $ptw(\varphi)$ and $itv(\varphi)$ the trewidth of the primal and incidence graphs of $\varphi$ respectively.
Proposition 2.
$itw(\varphi)\leq ptw(\varphi)$ [18]
3 $\wedge_{d}$-obdd: incidence treewidth inefficiency and separations
3.1 An approach to proving $\wedge_{d}$-obbd lower bounds
We start from overviewing an approach for proving lower bounds for obdds and then generalize it to $\wedge_{d}$-obdds.
Let $B$ be an obdd obeying an order $\pi$. Let ${\bf g}$ be an assigment over a prefix of $\pi$. Then we can identify the maximal path $P=P({\bf g})$ starting from the source and such that ${\bf a}(P)\subseteq{\bf g}$. Indeed, existence of such a path and its uniqueness is easy to verify algorithmically. Let the current vertex $u$ of $B$ be the source and the current variable $x$ be the label of $u$. Furthermore, let the current path $P$ be consisting of $u$ only. If $x\notin\mathsf{var}({\bf g})$ then stop. Otherwise, choose the out-neighbour $v$ of $u$ such that $(u,v)$ is labelled with ${\bf g}(x)$. Let the current vertex be $v$, the current variable be the label of $v$ and the current path be $P$ plus $(u,v)$. Repeat the process until the stopping condition reached. We let $u({\bf g})$ be the final vertex of $P({\bf g})$.
Now, let $\mathcal{F}$ be a set of fooling assignments, for each ${\bf g}\in\mathcal{F}$, $Var({\bf g})$ is a prefix of $P$. If we prove that for every two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$, $u({\bf g}_{1})\neq u({\bf g}_{2})$ then this implies that $|B|\geq|\mathcal{F}|$. In particular, if $|\mathcal{F}|$ is large in terms of the number of variables, then the respective large lower bound follows.
If we try to adapt the above methodology to a $\wedge_{d}$-obdd, $B$ obeying an order $\pi$, we will notice that an assignment ${\bf g}$ as above is associated not with a path but rather with a rooted tree. The role is $u({\bf g})$ is played by the leaves of that tree. Consequently, there is no straightforward way to identify a single node associated with an assignment. We propose a generic methodology that helps to identify such a single node. In the rest of this subsection, we provide a formal description of this methodology.
Definition 8.
Let ${\bf g}$ be an assignment over a subset of $\mathsf{var}(B)$. The alignment of $B$ by ${\bf g}$ denoted by $B[{\bf g}]$ is obtained from $B$ by the following two operations.
  1. 1.
    For each decision node $u$ labelled by $x\in\mathsf{var}({\bf g})$, remove the outgoing edge of $u$ labelled with $1-{\bf g}(x)$. Let $B_{0}[{\bf g}]$ be the resulting subgraph of $B$.
  2. 2.
    Remove all the nodes of $B_{0}[{\bf g}]$ that are not reachable from the source of $B[{\bf g}]$.
Example 2.
Figure 1 demonstrates a $\wedge_{d}$-obdd $B$ for $(x_{1}\vee x_{2})\wedge(x_{2}\vee x_{3})\wedge(x_{4}\vee x_{5})\wedge(x_{5}\vee x_{6})$ (on the left) obeying the order $(x_{2},x_{1},x_{3},x_{5},x_{4},x_{6})$ and $B[{\bf g}]$ (on the right) where ${\bf g}=\{(x_{2},1)\}$.
The $true$ and $false$ sinks are respectively labelled with $T$ and $F$. Note that, for the sake of a better visualization, we deviated from the definition requiring a $\wedge_{d}$-obdd to have a single $true$ and a single $false$ sinks. To make the models on the picture consistent with the definition, all $true$ sinks of each model need to be contracted together and the same contraction needs to be done for all the $false$ sinks.
[…]
Figure 1: A $\wedge_{d}$-obdd and its alignment
We note that $B[{\bf g}]$ is not necessarily a $\wedge_{d}$-obdd since some decision nodes may have only one outgoing edge. This situation is addressed in the following definition.
Definition 9.
A decision node of $B[{\bf g}]$ that has only one outgoing neighbour is called an incomplete decision node. A maximal path all intermediate decision nodes of which are incomplete is called an incomplete path. We note that a $\wedge_{d}$ node cannot be a final node of an incoplete path due to its maximality, hence an incomplete path may end only with a sink of $B$ or with a decision node of $B$. We denote by $L({\bf g})$ the set of all decision nodes that are final nodes of the incomplete paths of $B[{\bf g}]$.
Back to Example 2, $L({\bf g})$ is a singleton consisting of the only node labelled by $x_{5}$. We note that $L({\bf g})$ are exactly minimal complete decision nodes of $B[{\bf g}]$.
Lemma 2.
Let $\pi$ be the order obeyed by $B$ containing $\mathsf{var}(B)$ and let ${\bf g}$ be an assignment over a prefix of $\pi$. Then the following statements hold regarding $B[{\bf g}]$.
  1. 1.
    Let $u$ be a complete decision node and let $v$ be a decision node being a successor of $u$ (that is, $B[{\bf g}]$ has a path from $u$ to $v$). Then $v$ is also a complete decision node. In other words, $B[{\bf g}]_{u}$ is a $\wedge_{d}$-obdd.
  2. 2.
    Let $u_{1},u_{2}\in L({\bf g})$. Then $\mathsf{var}(B[{\bf g}]_{u_{1}})\cap\mathsf{var}(B[{\bf g}]_{u_{2}})=\emptyset$.
  3. 3.
    Let $T({\bf g})$ be the union of all incomplete paths ending with nodes of $L({\bf g})$. Then $T({\bf g})$ is a rooted tree.
Proof.
Let $x$ and $y$ be the variables labelling $u$ and $v$ respectively. By assumption $x<_{\pi}y$ and $x\notin Var({\bf g})$, the latter constitutes a prefix of $\pi$. It follows that $y\notin Var({\bf g})$ and hence $v$ is complete simply by construction.
For the second statement, note that, by assumption, none of $u_{1}$ and $u_{2}$ is a successor of the other. Therefore, in $B[{\bf g}]$ there is a node $v$ with two children $v_{1}$ and $v_{2}$ such that $u_{i}$ is a (possibly not proper) successor of $v_{i}$ for each $i\in\{1,2\}$. If $v$ is a decision node then it is complete, in contradiction to the minimality of $u_{1}$ and $u_{2}$. We conclude that $v$ is a conjunction node that hence $\mathsf{var}(B[{\bf g}]_{v_{1}})\cap\mathsf{var}(B[{\bf g}]_{v_{2}})=\emptyset$. As $\mathsf{var}(B[{\bf g}]_{u_{1}})\cap\mathsf{var}(B[{\bf g}]_{u_{2}})\subseteq\mathsf{var}(B[{\bf g}]_{v_{1}})\cap\mathsf{var}(B[{\bf g}]_{v_{2}})$, the second statement follows.
For the third statement, it is sufficient to prove that any two distinct incomplete paths $P_{1}$ and $P_{2}$ do not have nodes in common besides their comon prefixes. So, let $P$ be the longest common prefix of $P_{1}$ and $P_{2}$ and let $u$ be the last vertex of $P_{1}$ and $P_{2}$. Arguing as in the previous paragraph, we conclude that $u$ is a $\wedge_{d}$ node. Let $u_{1}$ and $u_{2}$ be the children of $u$. Then, by the maximality of $P$, we can assume w.l.o.g. that $u_{1}\in V(P_{1})$ and $u_{2}\in V(P_{2})$.
Suppose that $P_{1}$ and $P_{2}$ have a common vertex $v$ that is a proper successor of $u$. As $v$ is a predecessor of a vertex of $L({\bf g})$, $\mathsf{var}(B_{v})\neq\emptyset$. As $\mathsf{var}(B_{v})\subseteq\mathsf{var}(B_{u_{1}})\cap\mathsf{var}(B_{u_{2}})$, we get a contradiction to the decomposability of $u$. ∎
Following Lemma 2, we denote by $T({\bf g})$ the union of incomplete paths of $B[{\bf g}]$ ending with a node of $L({\bf g})$. The tree $T({\bf g})$ will not be essential for our further reasoning. However, it is important to underscore the direction of upgrade reasoning as we move from obddto $\wedge_{d}$-obdd: instead of the path $P({\bf g})$, we consider the tree $T({\bf g})$.
With the premises of Lemma 2, $L({\bf g})$ provides a neat characterization of $\mathcal{S}(B)|_{\bf g}$ as established in the next lemma.
Lemma 3.
Let the notation and premises be as in Lemma 2. In addition, assume that $\mathcal{S}(B)|_{\bf g}\neq\emptyset$. Let $X=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\bigcup_{u\in L_{\bf g}}\mathsf{var}(B_{u})$. If $L({\bf g})\neq\emptyset$ then $\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}\times\prod_{u\in L({\bf g})}\mathcal{S}(B_{u})$. Otherwise, $\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}$.
Lemma 3 is proved by an elementary though lengthy inductive exploration of the structure of $\wedge_{d}$-obdd. A complete proof is postponed to the Appendix.
We are now ready to outline a methodology for obtaining a lower bound for a $\wedge_{d}$-obdd $B$ obeying an order $\pi$. Like in the case for obdds, we introduce a large set $\mathcal{F}$ set of fooling assignments and seek to prove that $|B|\geq|\mathcal{F}|$. Since, we are dealing with a tree $T({\bf g})$ rather than $P({\bf g})$, we make an extra effort of identifying a leaf $u({\bf g})$ of $T({\bf g})$. Having done that, we prove, like in case of obdd, that for two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$, $u({\bf g}_{1})\neq u({\bf g}_{2})$ thus implying the desired lower bound $|B|\geq|\mathcal{F}|$.
In order to identify $u({\bf g})$, we introduce a set $UB({\bf g})\subseteq\mathsf{var}(B)\setminus\mathsf{var}({\bf g})$ which we refer to as the unbreakable set for ${\bf g}$. This is a set of at least two variables and we prove that $\mathcal{S}(B)|_{\bf g}$ does not break $UB({\bf g})$ (hence the name).
Next, we employ Lemma 3 to claim that either $UB({\bf g})\subseteq\mathsf{var}(B_{u})$ for some $u\in L({\bf g})$ or $UB({\bf g})\subseteq X$ where $X$ is as in the statement of Lemma 3. The latter possibility is ruled outr since $UB({\bf g})$ is of size at least $2$ and $\mathsf{var}(X)$ can be further broken into singletons. We thus set $u({\bf g})$ to be the $u\in L({\bf g})$ such tha $UB({\bf g})\subseteq\mathsf{var}(B_{u})$.
For the last stage of the proof, if it happens that $u({\bf g}_{1})=u({\bf g}_{2})=u$ then $UB({\bf g}_{1})\cup UB({\bf g}_{2})\subseteq\mathsf{var}(B_{u})$ simply by definition. It thus follows from Lemma 3 that $\mathcal{S}|_{{\bf g}_{1}}$ and $\mathcal{S}|_{{\bf g}_{2}}$ have the same projections to $UB({\bf g}_{1})\cup UB({\bf g}_{2})$. We prove that these projections are distinct and hence establish that $u({\bf g}_{1})\neq u({\bf g}_{2})$.
3.2 $\wedge_{d}$-obdd lower bounds and separations
In [3] we established an xp lower bound for $\wedge_{d}$-obdds representing cnfs of bounded incidence treewidth. We prove this result here using the approach outlined in Section 3.1.
Theorem 4.
There is a constant $c$ and an infinite set of values $k$ such that for each $k$ there is an infinite set $\Phi$ of cnfs such that for each $\varphi\in\Phi$ $itw(\varphi)\leq k$ and $\wedge_{d}$-obdd$(\varphi)\geq|\mathsf{var}(\varphi)|^{c\cdot k}$.
Thus a natural restriction on $\wedge_{d}$-fbdd separates primal [11] and incidence treewidth. This makes $\wedge_{d}$-obdd an interesting object to study in greater detail. We establish the relationship of $\wedge_{d}$-obdd with two closely related formalisms obdd and fbdd by proving exponential separations as specified below (for both statements, the lower bounds are proved using the approach as in Section 3.1).
Theorem 5.
There is an infinite class of cnfs polynomially representable as fbdds but requiring exponential-size representation as $\wedge_{d}$-obdds.
Theorem 6.
There is an infinite class of cnfs polynomially representable as $\wedge_{d}$-obdds but requiring exponential-size representation as obdds.
Remark 2.
As $\wedge_{d}$-obdd is not a special case of fbdd it is natural to ask whether an opposite of Theorem 5 holds. It is known from [1] that $\wedge_{d}$-fbdd can be quasipolynomially simulated by fbdd. As $\wedge_{d}$-obdd is a special case of $\wedge_{d}$-fbdd, the simulation clearly applies to $\wedge_{d}$-obdd.
For the corresponding quasipolynomial separation, take a class of cnfs of $ptw$ at most $k$ that require fbdd of size $n^{\Omega(k)}$ (see [15] for such a class of cnfs). Let $k=\log n$ where $n$ is the number of variables. Then the resulting class of cnfs will require $n^{\Omega(\log(n)}$ size fbdds for their representation but can be represented by a polynomial size $\wedge_{d}$-obdd according to [11].
Remark 3.
Theorem 6 looks somewhat surprising because equipping fbdds with the $\wedge$ gates leads only to quasipolynomial gains in size [1]
In the two definitions below, we introduce classes of cnfs needed for the stated lower bounds and separations.
Definition 10.
Let $G$ be a graph without isolated vertices. Then $\varphi(G)$ is a cnf whose set of variables is $V(G)$ and the set of clauses is $\{(u\vee v)|\{u,v\}\in E(G)\}$.
Let $V=V(G)$ and let $V[1]=\{v[1]|v\in V\}$ and $V[2]=\{v[2]|v\in V\}$ be two disjoint copies of $V$. Let $G^{*}$ be the graph with $V(G^{*})=V[1]\cup V[2]$ and $E(G^{*})=\bigcup_{\{u,v\}\in E(G)}\{\{u[1],v[2]\},\{v[1],u[2]\}\}$.
The cnf $\psi(G)$ is obtained from $\varphi(G^{*})$ by introducing two additional clauses $\neg V[1]=\{\neg v[1]|v\in V\}$ and $\neg V[2]=\{\neg v[2]|v\in V\}$.
Extending the notation $V[i]$, for $U\subseteq V(G)$, and $i\in\{1,2\}$, we let $U[i]=\{u[i]|u\in U\}$.
Example 3.
Let $G$ be a graph with $V(G)=\{u_{1},\dots,u_{5}\}$ and
$E(G)=\{\{u_{1},u_{2}\},\{u_{2},u_{3}\},\{u_{3},u_{4}\},\{u_{4},u_{5}\},\{u_{5},u_{1}\},\{u_{1},u_{3}\}\}$. In other words, $G$ is a cycle of $5$ vertices with an extra chord. Then $V(G^{*})=\{u_{1}[1],\dots,u_{5}[1],u_{1}[2],\dots,u_{5}[2]\}$ and $E(G^{*})=\{\{u_{1}[1],u_{2}[2]\},\{u_{2}[1],u_{3}[2]\},\{u_{3}[1],u_{4}[2]\},\{u_{4}[1],u_{5}[2]\},\{u_{5}[1],u_{1}[2]\},\{u_{1}[1],u_{3}[2]\},\\ \{u_{1}[2],u_{2}[1]\},\{u_{2}[2],u_{3}[1]\},\{u_{3}[2],u_{4}[1]\},\{u_{4}[2],u_{5}[1]\},\{u_{5}[2],u_{1}[1]\},\{u_{1}[2],u_{3}[1]\}\}$. $\varphi(G)$ and $\varphi(G^{*})$ simply turn the edges $\{u,v\}$ of the respective graphs into clauses $(u\vee v)$. $\psi(G)$ is obtained from $\varphi(G^{*})$ by adding clauses ($(\neg u_{1}[1]\vee\dots\vee\neg u_{5}[1])$ and ($(\neg u_{1}[2]\vee\dots\vee\neg u_{5}[2])$
We introduce two more classes of formulas that are obtained from $\varphi(G)$ and $\psi(G)$ by introducing junction variables.
Definition 11.
Let $G$ be a graph without isolated vertices and let $E_{1},E_{2}$ be a partition of $E(G)$ so that $V(G[E_{1}])=V(G[E_{2}])=V(G)$.
We denote by $\varphi(G,E_{1},E_{2})$ the function $(\textsc{jn}\rightarrow\varphi(G[E_{1}]))\wedge(\neg\textsc{jn}\rightarrow\varphi(G[E_{2}])$ where jn is a junction variable that does not belong to $V(G)$. It is not hard to see that, $\varphi(G,E_{1},E_{2})$ can be expressed as a cnf $\bigwedge_{\{u,v\}\in E_{1}}(\neg\textsc{jn}\vee u\vee v)\wedge\bigwedge_{\{u,v\}\in E_{2}}(\textsc{jn}\vee u\vee v)$.
Similarly, we denote by $\psi(G,E_{1},E_{2})$ the function $(\textsc{jn}\rightarrow\psi(G[E_{1}]))\wedge(\neg\textsc{jn}\rightarrow\psi(G[E_{2}])$
Remark 4.
We will use cnfs with junction variables for Theorems 5 and 6. This approach is used in [20] for separation of obdds and fbdds. The rough idea is that a ’hard’ function is ’split’ into two ’easy’ ones using the junction variable. The fbdd, due to its flexibility, can adjust and represent each ’easy’ function with the suitable variable ordering. On the other hand, obdd is rigid in the sense that it needs to use the same ordering for both ’easy’ functions and the ordering leads to an exponential size representation for one of them.
In order to proceed, we need to delve deeper into the structure of matchings of graph $G^{*}$. This is done in the next definition and two statements that follow.
Definition 12.
Let $M$ be a matching of $G^{*}$ between $V^{\prime}$ and $V^{\prime\prime}$. We say that $V^{\prime},V^{\prime\prime}$ is the neat partition of $V(M)$ if there are $U,W\in V(G)$ such that,say $V^{\prime}=U[1]$ and $V^{\prime\prime}=W[2]$.
Let $\pi^{*}$ be a linear order of $V(G^{*})$. We say that $M$ neatly crosses $\pi^{*}$ if there a prefix $\pi^{*}_{0}$ of $\pi^{*}$ such that $\pi^{*}_{0}\cap V(M)$ and $(\pi^{*}\setminus\pi^{*}_{0})\cap V(M)$ form a neat partition of $M$.
In other words $M$ neatly crossing $\pi^{*}$ means that $M$ crosses $\pi^{*}$ with a witnessing prefix satisfying the constraint as in the definition. In the next statement, we demonstrate that if each linear order of $G$ is crossed by a large induced matching then every linear order of $G^{*}$ is neatly crossed by a large induced matching. The statement is formulated in terms of parameter lsim-width defined earlier.
Lemma 7.
Every linear order of $G^{*}$ is neatly crossed by an induced matching of size at least $lsimw(G)$
Proof.
Let $\pi^{*}$ be a linear order of $V(G^{*})$. Let $\pi$ be a linear order of $V(G)$ obtained from $\pi^{*}$ as follows. For each $v\in V(G)$, let $ind(v)=1$ if $v[1]$ appears in $\pi^{*}$ before $v[2]$ and $2$ otherwise. In $\pi$, $u$ is ordered before $v$ if and only if $u[ind(u)]$ is ordered before $v[ind(v)]$.
Let $\pi_{0}$ be a prefix of $\pi$ witnessing a matching $M$ of size at least $lsimw(G)$ crossing $\pi$. Let $U=V(M)\cap\pi_{0}$. Then $U$ is can be weakly partitioned into $U_{1}$ and $U_{2}$ such that $U_{i}$ is the set of $u\in U$ such that $ind(u)=i$. By the pigeonhole principle, the size of either $U_{1}$ or $U_{2}$ is at least $|U|/2$. We assume the former w.l.o.g. Let $M_{1}$ be the subset of $M$ consisting of all $e\in M$ such that $U_{1}\cap e\neq\emptyset$. Let $W_{1}$ be $V(M_{1})\setminus U_{1}$. Clearly, $M_{1}$ is a matching between $U_{1}$ and $W_{1}$. Let $M^{*}_{1}=\{\{u[1],w[2]\}|\{u,v\}\in M_{1}$, $u\in U_{1},w\in W_{1}\}$. It is not hard to see that $M^{*}_{1}$ is an induced matching of $G^{*}$ between $U_{1}[1]$ and $W_{1}[2]$.
Let $\pi^{*}_{0}$ be the largest prefix of $\pi^{*}$ finishing with a vertex of $U_{1}[1]$ and let $\pi^{*}_{1}$ be the matching suffix. We claim that $\pi^{*}_{0}$ is a witnessing prefix for $\pi^{*}$ being neatly crossed by $M^{*}_{1}$ thus establishing the lemma. It is immediate from the definition that $U_{1}[1]\subseteq\pi^{*}_{0}$. It thus remains to observe that $W_{1}[2]\subseteq\pi^{*}_{1}$.
Let $u[1]$ be the last vertex of $\pi^{*}_{0}$. Let $w[2]\in W_{1}[2]$. Note that by definition of $\pi_{0}$, $u\in\pi_{0}$ while $w$ is not. It follows from $u[ind(u)]<w[ind(w)]$ according to $\pi^{*}$. Now $ind(u)=1$ by definition of $U_{1}$. That is $u[1]$ occurs before $w[ind(w)]$ and surely before $w[3-ind(w)]$ in $\pi^{*}$. As $u[1]$ is the last element of $\pi^{*}_{0}$, both occurrences of $w$ are outside of $\pi^{*}_{0}$. ∎
Lemma 8.
Let $G$ be a graph without isolated vertices. Let $E_{1},E_{2}$ be a partition of $E(G)$ so that $V(G[E_{1}])=V(G[E_{2}])=V(G)$. (Note that this means that $V(G^{*})=V(G[E_{1}]^{*})=V(G[E_{2}]^{*})$). Then the following statements hold.
  1. 1.
    Let $\pi$ be a linear order of $V(G)$. Then either $G[E_{1}]$ or $G[E_{2}]$ has an induced matching of size at least $lsimw(G)/2$ crossing $\pi$.
  2. 2.
    Let $\pi^{*}$ be a linear order of $V(G^{*})$, Then either $G[E_{1}]^{*}$ or $G[E_{2}]^{*}$ has an induced matching of size at least $lsimw(G)/4$ neatly crossing $\pi^{*}$.
Proof.
Let $M$ be a matching of $G$ of size at least $lsimw(G)$ crossing $\pi$. Clearly either $M\cap E_{1}$ or $M\cap E_{2}$ is of size at least $|M|/2$. Assume the former w.l.o.g. As $M\cap E_{1}\subseteq M$, clearly, $M\cap E_{1}$ crosses $\pi$. This proves the first statement.
For $i\in\{1,2\}$, Let $E^{*}_{i}=E(G[E_{i}]^{*})$. It is not hard to see that $E^{*}_{1},E^{*}_{2}$ partition $E(G^{*})$ and $G^{*}[E^{*}_{i}]=G[E_{i}]^{*}$ for each $i\in\{1,2\}$. Let $M^{*}$ be an induced matching of size at least $lsimw(G)/2$ neatly crossing $\pi^{*}$ existing by Lemma 7. For each $i\in\{1,2\}$, let $M^{*}_{i}=M^{*}\cap E^{*}_{i}$. Clearly, both $M^{*}_{1}$ and $M^{*}_{2}$ cross $\pi^{*}$ ans least one of them is of size at least half the size of $M^{*}$ that is at least $lsimw(G)/4$. ∎
Remark 5.
Note that Lemma 8 does not make any statements regarding $lsimw(G[E_{1}])$ and $lsimw(G[E_{2}])$. Indeed, if $G$ is an $n\times n$ grid, $E_{1}$ is the set of ’horizontal’ edges and $E_{2}$ is the set of ’vertical’ edges then $lsimw((G[E_{1}])=lsimw(G[E_{2}])=1$, while $lsimw(G)=\Omega(n)$. Nevertheless, Lemma 8 states that for each linear order $\pi$ of $V(G)$ must be $\Omega(n)$ either for $G[E_{1}]$ or for $G[E_{2}]$. Continuing on the discussion in Remark 4, the above phenomenon is precisely the reason of rigidity of obdd-based models representing cnfs with junction variables. We will see in the proof of Theorems 5 and 6 below how this leads to exponential lower bounds.
Next, we state two theorems, whose proof are provided in Section 3.3, that serve as engines for proving lower bounds for Theorems 4, 5, and 6. We will then use Lemmas 7 and Lemma 8 for ’harnessing’ these engines in order to prove the theorems.
Theorem 9.
Let $B$ be a $\wedge_{d}$-obdd representing $\psi(G)$. Let $\pi^{*}$ be a linear order of $\mathsf{var}(\psi(G))$ obeyed by $B$. Let $M$ be a matching of $G^{*}$ neatly crossing $\pi^{*}$. Then $|B|\geq 2^{\Omega(|M|)}$.
Theorem 10.
Let $B$ be an obdd representing $\varphi(G)$. Let $\pi$ be a linear order of $V(G)$ obeyed by $B$. Let $M$ be a matching of $G$ crossing $\pi$. Then $|B|\geq 2^{\Omega(|M|)}$.
Proof.
(Theorem 4) For a sufficiently large $k$, let ${\bf G}_{k}$ be a class of graphs $G_{k}$ of treewidth at most and lsim-width $\Omega(k\log n)$. Existence of such a class of graphs follows Lemma 1. In order to prove the theorem it is sufficient to demonstrate that $\wedge_{d}$-obdd representing $\psi(G_{k})$ is of size $n^{\Omega(k)}$ and that $itw(\psi(G))=O(k)$.
Let $B$ be a $\wedge_{d}$-obdd representing $\psi(G_{k})$. Let $\pi^{*}$ be a linear order of $V(G^{*}_{k})$ obeyed by $B$. By Lemma 7, there is a matching $M$ of $G^{*}_{k}$ of size $\Omega(lsimw(G_{k}))$ neatly crossing $\pi^{*}$. Hence, by Theorem 9, the size of $2^{\Omega(lsimw(G))}$ which is $n^{\Omega(k)}$ according to the previous paragraph.
For the proof that $itw(\psi(G))=O(k)$, let $(T,{\bf B})$ be a tree decomposition of $G$ of width at most $k$. For each $t\in V(T)$, replace each $v\in{\bf B}(t)$ with $v[1]$ and $v[2]$. It is not hard to see that we obtain a tree decomposition for the primal graph $\varphi(G^{*})$. As the largest bag size of $(T,{\bf B})$ is at most $k+1$, the largest bag size of the new tree decomposition is at most $22k+$ 2 and hence $ptw(\varphi(G^{*}))$ is at most $12k+$ 1. By Proposition 2 $itw(\varphi(G^{*}))$ is at most $12k+$ 1. Take a tree decomposition witnessing $itw(\varphi(G^{*}))$ and add to each bag two vertices corresponding to the long clauses. As a result we obtain a tree decomposition of the incidence graph $\psi(G)$ of width at most $32k+$ 3. ∎
We are now turning to the proof of Theorems 5 and 6. For both theorems, the target class of graphs are $n\times n$ grids that we denote by $G_{n}$. A formal definition is provided below.
Definition 13.
Let $n>0$ be an integer. A $n\times n$ grid denoted by $G_{n}$ is a graph with $V(G_{n})=\{(i,j,)|1\leq i\leq n,1\leq j\leq n\}$ and $E(G_{n})=E_{hor}\cup E_{vert}$ where $E_{hor}=\{\{(i,j),(i,j+1)\}|1\leq i\leq n,1\leq j\leq n-1\}$ and $E_{vert}=\{\{(i,j),(i+1,j\}|1\leq i\leq n-1,1\leq j\leq n\}$. We also refer to $E_{hor}$ and $E_{vert}$ as, respectively, horizontal and vertical edges of $G_{n}$.
As the max-degree of $G_{n}$ is $4$, $lsimw(G_{n})$ and $lmmw(G)$ are linearly related. Next, it is known [16] that $pw(G_{n})$ and $lmmw(G_{n})$ are, in turn, linearly related. Finally, it is well know that $pw(G_{n})\geq tw(G_{n})\geq\Omega(n)$. Therefore, we conclude the following.
Proposition 3.
$lsimw(G_{n})=\Omega(n)$.
We also need an auxuliary lemma whose proof is provided in the Appendix.
Lemma 11.
Let $B$ be a $\wedge_{d}$-obdd obeying an order $\pi$. Let ${\bf g}=\{(x,i)\}$ where $x\in var(B)$ and $i\in\{0,1\}$. Suppose that all the variables of $f(B)|_{\bf g}$ are essential. Let $B^{\prime}$ be obtained from $B$ by the following operations.
  1. 1.
    Removal of all the edges $(u,v)$ such that $u$ is a decision node associated with $x$ and $(u,v)$ is labelled with $2-i$.
  2. 2.
    Contraction of all the edges $(u,v)$ such $u$ is a decision node associated with $x$ and $(u,v)$ is labelled with $i$ (meaning that $(u,v)$ is removed and $u$ is identified with $v$).
  3. 3.
    Removal of all the nodes not reachable from the source of the resulting graph.
Then $B^{\prime}$ is a $\wedge_{d}$-obdd obeying $\pi$ and $f(B^{\prime})=f(B)|_{\bf g}$.
Proof.
(Theorem 5) We demonstrate that the statement of the theorem holds for the family of cnfs $\psi(G_{n},E_{hor},E_{vert})$ with $n>0$. For the lower bound, let $B$ be a $\wedge_{d}$-obdd representing $\psi(G_{n},E_{hor},E_{vert})$. Let $\pi_{0}$ be a linear order obeyed by $B$. Let $\pi^{*}$ be the linear order on $V(G^{*})$ induced by $\pi_{0}$ (that is, two elements are ordered by $\pi^{*}$ in exactly the same way as they are ordered by $\pi_{0}$).
By Lemma 8, we can assume w.l.o.g, that $\pi^{*}$ is neatly crossed by a matching $M$ of $G_{n}[E_{hor}]^{*}$ of size $\Omega(lsimw(G_{n}))$. By Proposition 3, $|M|=\Omega(n)$. Then, by Theorem 9, we conclude that
Claim 1.
A $\wedge_{d}$-obdd representing $\psi(G_{n}[E_{hor}])$ and obeying $\pi^{*}$ is of size $2^{\Omega(n)}$.
Let ${\bf g}=\{(\textsc{jn},1)\}$ It is not hard to see that $\psi(G_{n},E_{hor},E_{vert})|_{\bf g}=\psi(G_{n}[E_{hor}])$ and the latter does not have non-essential variables. By Lemma 11, a $\wedge_{d}$-obdd representing $\psi(G_{n}[E_{hor}])$ and obeying $\pi^{*}$ can be obtained from $B$ by a transformation that does not increase its size By Claim 1, the size of $B$ is $2^{\Omega(n)}$ as required.
For the upper bound, we consider an fbdd whose source $rt$ is associated with jn. Let $u_{0}$ and $u_{1}$ be the children of $rt$ with respective edges $(rt,u_{0})$ and $(rt,u_{1})$ labelled with 00 and $1$. Then $u_{0}$ is the source of an fbdd representing $\psi(G_{n}[E_{vert}])$ and $u_{1}$ is the source of an fbdd representing $\psi(G_{n}[E_{hor}])$. It is not hard to see that the resulting fbdd indeed represents $\psi(G_{n},E_{hor},E_{vert})$. It remains to demonstrate that both $\psi(G_{n}[E_{hor}])$ and $\psi(G_{n}[E_{vert}])$ can be represented by poly-size fbdds. We demonstrate that the incidence pathwidth of both $\psi(G[E_{hor}])$ and $\psi(G_{n}[E_{hor}])$ and $\psi(G_{n}[E_{vert}])$ is at most $7$. It then follows [13] that both these cnfs can be represented by linear-size fbdds.
Let $\pi$ be the linear order over $V(G)=V(G[E_{hor}])$ where the vertices occur in the ’dictionary’ order: $(i_{1},j_{1})$ precedes $(i_{2},j_{2})$ if $i_{1}<i_{2}$ or $i_{1}=i_{2}$ and $j_{1}<j_{2}$ and let $\pi_{i}$ be the vertex in the position $i$ of $\pi$. That is, $\pi=(\pi_{1},\dots,\pi_{n^{2}})$. We consider a path decomposition $(P,{\bf B})$ where $P=\pi_{1},\dots,\pi_{n^{2}}$ and for each $u\in V(P)$, ${\bf B}(u)$ is defined as follows.
If $u=(i,1)$ for $1\leq i\leq n$ then ${\bf B}(u)=\{u[1],u[2],\neg V(G)[1],\neg V(G)[2]\}$ (recall that the last two elements correspond to the long clauses). If $u=(i,j)$ for $1\leq i\leq n$ and $2\leq j\leq n$, let $v=(i,j-1)$. Then ${\bf B}(u)=\{u[1],u[2],v[1],v[2],(u[1]\vee v[2]),(u[2]\vee v[1]),\neg V(G)[1],\neg V(G)[2]\}$. It follows from a direct inspection that $(P,{\bf B})$ is a tree decomposition of the incidence graph of $\psi(G_{n}[E_{hor}])$ of width $7$. For $\psi(G_{n}[E_{vert}])$, the construction is analogous with rows and columns changing their roles. ∎
Proof.
(Theorem 6) We consider the class of cnfs $\varphi(G_{n},E_{hor},E_{vert})$. Let $B$ be an obdd representing $\varphi(G_{n},E_{hor},E_{vert})$. Let $\pi_{0}$ be an order obeyed by $B$ and let $\pi$ be a linear order of $V(G_{n})$ induced by $\pi_{0}$. By Lemma 8, we can assume w.l.o.g. the existence of a matching $M$ of $\varphi(G_{n}[E_{1}])$ of size at least $\Omega(n)$ crossing $\pi$. By Theorem 10, an obdd obeying $\pi$ and representing $\varphi(G_{n}[E_{1}])$ is of size $2^{\Omega(n)}$. As all the variables of $\varphi(G_{n}[E_{1}])$ are essential, It follows from Lemma 11 that an obdd representing $\varphi(G_{n}[E_{1}])$ can be obtained from $B$ by a transformation that does not increase its size (note that the transformation does not introduce conjunction nodes hence, being applied to an obdd produces another obdd). We thus conclude that the size of $B$ is at least $2^{\Omega(n)}$.
For the upper bound, we demonstrate existence of a poly-size $\wedge_{d}$-obdd representing $\varphi(G_{n}E_{hor},E_{vert})$ and obeying the order $\pi_{d}$ where the first element is jn followed by $V(G_{n})$ in the dictionary order, that is $(i_{1},j_{1})$ precedes $(i_{2},j_{2})$ if $i_{1}<i_{2}$ or $i_{1}=i_{2}$ and $j_{1}<j_{2}$.
Let $Hor_{i}$ be the subgraph of $G_{n}$ induced by row $i$, that is the path $(i,1),\dots(i,n)$ and $Vert_{i}$ be the subgraph of $G_{n}$ induced by the column $i$, that is the path $(1,i),\dots,(n,i)$. By direct layerwise inspection, we conclude that
Claim 2.
There are linear sized obdds obeying $\pi_{d}$ and representing $\varphi(Hor_{i})$ and $\varphi(Vert_{i})$ for each $1\leq i\leq n$.
Proof.
In order to represent $\mathsf{var}(Hor_{i})$, let us query the variables by the ascending order of their second coordinate, that is $(i_{1},),\dots,(i,n)$. For each $11\leq j\leq n-$ 1, let ${\bf H}_{j}$ be the set of all functions $\varphi(Hr_{i})|_{\bf g}$ where ${\bf g}$ is an assignment over $(i,1),\dots,(i,j)$. It is not hard to see that that there are at most $3$ such functions: constant zero, and two functions on $(i,j+1),\dots,(i,n)$ completely determined by the assignment of $(i,j)$. It is well known [20] that the resulting obdd can be seen as a dag with layers $11,\dots,n+$ 1, the last layer is reserved for the sink nodes, and each layer $1\leq j\leq n$ containing nodes labelled with $(i,j)$ and having $O(1)$ size.
The argument for $\varphi(Vert_{i})$ is similar with the first coordinate of its variables being used instead of the second one. ∎
Let $B^{h}_{i}$ and $B^{v}_{i}$ be obdds obeying $\pi_{d}$ and representing $Hor_{i}$ and $Vert_{i}$ as per Claim 2. By using $O(n)$ conjunction nodes, it is easy to create $\wedge_{d}$-obdds $B_{1}$ and $B_{0}$ that are, respectively conjunctions of all $B^{v}_{i}$ and all $B^{h}_{i}$ and hence, respectively represent $\wedge_{1\leq i\leq n}\varphi(Hor_{i})$ and $\wedge_{1\leq i\leq n}\varphi(Vert_{i})$. It is not hard to see that the former conjunction is $\varphi(G[E_{hor}])$ and the latter one is $\varphi(G[E_{vert}])$, hence they are, respectively represented by $B_{1}$ and $B_{0}$. The resulting $\wedge_{d}$-obdd is then constructed as follows. The source $rt$ is a decision node labelled with jn. Let $u_{1}$ and $u_{0}$ be the children of $rt$ labelled by $1$ and 00 respectively. Then $u_{1}$ is the source of $B_{1}$ and $u_{0}$ is the source of $B_{0}$. ∎
Remark 6.
Note that the variables of a single column of $G_{n}$ explored along the column satisfy the dictionary order. However, if we try to explore first one column and then another one then the dictionary order will be violated. Thus, the upper bound in the proof of Theorem 6 demonstrates how the conjunction nodes overcome the rigidity of obdd by splitting the set of all variables in chunks that do obey a fixed order.
3.3 Proofs of Theorem 9 and 10
We first prove Theorem 9 using the approach as outlined in Section 3.1 . The proof of Theorem 10 will follow the same approach but with a simplified reasoning.
The first step of the approach is to define a fooling set of assignments. W.l.o.g. we assume existence of sets $U,W\subseteq V(G)$. such that $M$ is a matching between $U[1]$ and $W[2]$ and that each element of $U[1]$ is ordered by $\pi^{*}$ before each element of $W[2]$. We let $\pi_{0}$ be the prefix of $\pi^{*}$ whose last element is the last element of $U[1]$ in $\pi^{*}$.
Definition 14.
An assignment ${\bf g}$ over $\pi_{0}$ is fooling if it satisfies the following two conditions.
  1. 1.
    At least one variable of $U[1]$ is assigned with 00 and at least two variables of $U[1]$ are assigned with $1$.
  2. 2.
    All the variables of $\pi_{0}\setminus U[1]$ are assigned with $1$.
We denote by $\mathcal{F}$ the set of all fooling assignments.
Example 4.
Let $G$ be a matching consisting of three edges $\{u_{1},w_{1}\},\{u_{2},w_{2}\},\{u_{3},w_{3}\}$ Let $\pi^{*}=(u_{1}[1],u_{2}[1],u_{1}[2],u_{2}[2],w_{1}[1],w_{2}[1],u_{3}[1],w_{1}[2],w_{2}[2],w_{3}[2],u_{3}[2],w_{3}[1])$ Let $M=\{\{u_{1}[1],w_{1}[2]\},\{u_{2}[1],w_{2}[2]\},\{u_{3}[1],w_{3}[2]\}$. In this specific case, $U=\{u_{1}[1],u_{2}[1],u_{3}[1]\}$ and $W=\{w_{1}[2],w_{2}[2],w_{3}[2]\}$. Further on $\pi_{0}=(u_{1}[1],u_{2}[1],u_{1}[2],u_{2}[2],w_{1}[1],w_{2}[1],u_{3}[1])$. Now, $\mathcal{F}$ consists of all the assignments that map all of $u_{1}[2],u_{2}[2],w_{1}[1],w_{1}[2]$ to $1$ and assign $u_{1}[1],u_{2}[1],u_{3}[1]$ so that exactly two of them are assigned with $1$.
Before we proceed, we extend the notation by letting $U=\{u_{1},\dots,u_{q}\}$, $W=\{w_{1},\dots,w_{q}\}$ and $M=\{\{u_{1}[1],w_{1}[2]\},\dots,\{u_{q}[1],w_{q}[2]\}\}$. Further on, for $I\subseteq\{1,\dots,q\}$, we let $U_{I}=\{u_{i}|i\in I\}$ and $U[1]_{I}=\{u_{i}[1]|i\in I\}$. The sets $W_{I}$ and $W[2]_{I}$ are defined accordingly.
We now proceed to the next stage, that is, defining unbreakable sets.
Definition 15.
Let ${\bf g}\in\mathcal{F}$. We denote by $I({\bf g})$ the set of all $1\leq i\leq q$ such that $u_{i}[1]$ is assigned with $1$ by ${\bf g}$. We call $W[2]_{I({\bf g})}$ the unbreakable set of ${\bf g}$. Further on for $J\subseteq I({\bf g})$, we let ${\bf h}_{J}[{\bf g}]$ to be the assignment over $V(G^{*})$ such that ${\bf g}\subseteq{\bf h}_{J}[{\bf g}]$, for each $u\in V(G^{*})$, ${\bf h}_{J}[{\bf g}]=0$ if $u\in W[2]_{J}$ and $1$ otherwise.
Note that we introduce a notation for unbrekable sets specifically tailored to demonstrate the role of $I({\bf g})$.
Example 5.
Continuing on Example 4, let ${\bf g}=\{(u_{1}[1],0),(u_{2}[1],1),(u_{3}[1],1),\\ (u_{1}[2],1),(u_{2}[2],1),(w_{1}[1],1),(w_{2}[1],1)\}$. Then $I({\bf g})=\{2,3\}$ and the unbreakable set for ${\bf g}$ is $\{w_{2}[2],w_{3}[2]\}$. For $J=I({\bf g})$, ${\bf h}_{J}({\bf g})$ is the extension of ${\bf g}$ assigning $w_{1}[2]$ with $1$, $w_{2}[2]$ with 00 and $w_{3}[2]$ with 00.
The next stage of the proof is to show that the unbreakable sets are indeed unbrekable by $\mathcal{S}(B)|_{\bf g}$. We need first an auxiliary statement.
Lemma 12.
${\bf h}_{J}[{\bf g}]$ satisfies $\psi(G)$ if and only if $J\neq\emptyset$.
Proof.
First of all, if $J=\emptyset$ then $U[2]$ is assigned with $1$ for each $u\in V(G)$ thus falsifying the clause $\neg V(G)[2]$. Otherwise, we observe that $\neg V(G)[1]$ is satisfied because one of $U[1]$ is assigned negatively and $\neg V(G)[2]$ is satisfied because a variable of $W[2]_{J}$ is assigned negatively. Hence, $\psi(G)$ may be falsified only if both variables of a cause $(u[1]\vee w[2])$ are assigned with zeroes. By assumption $w[2]=w_{j}[2]$ for some $j\in J$. Further on, $u_{j}[1]$ assigned with $1$ by ${\bf g}$, but on the other hand, the only variables assigned by zeroes belong to $U[1]$. Hence $u[1]=u_{i}[1]$ for some $i\neq j$. By definition of $\psi(G)$, $G^{*}$ has an edge $\{u_{i}[1],w_{j}[2]\}$ in contradiction to $M$ being an induced matching. ∎
Now, we are ready to prove the unbreakbility of the unbreakable sets.
Lemma 13.
For each ${\bf g}\in\mathcal{F}$, $\mathcal{S}(B)|_{\bf g}$ does not break $W[2]_{I({\bf g})}$.
Proof.
Assume the opposite. Then $\mathcal{S}(B)|_{\bf g}=\mathcal{S}_{1}\times\mathcal{S}_{2}$ such that there is a partition $I_{1},I_{2}$ of $I({\bf g})$ so that $\mathsf{var}(\mathcal{S}_{j})\cap W[2]_{I({\bf g})}=W[2]_{I_{j}}$ for each $j\in\{1,2\}$.
By Lemma 12, ${\bf h}_{2}=Proj({\bf h}_{I_{1}}[{\bf g}],\mathsf{var}(\mathcal{S}_{2}))\in\mathcal{S}_{2}$ and ${\bf h}_{1}=Proj({\bf h}_{I_{2}}[{\bf g}],\mathsf{var}(\mathcal{S}_{1}))\in\mathcal{S}_{1}$. Then ${\bf g}\cup{\bf h}_{1}\cup{\bf h}_{2}$ is a satisfying assignment of $\psi(G)$. However, this is a contradiction since ${\bf g}\cup{\bf h}_{1}\cup{\bf h}_{2}$ sets all the variables of $V(G)[2]$ to $1$ thus falsifying $\neg V(G)[2]$. ∎
For the existence $u({\bf g})$, we now need to rule out the possibility that $W[2]_{I({\bf g})}\subseteq X$ where $X$ is as in Lemma 3.
Lemma 14.
Let ${\bf g}\in\mathcal{F}$. Then there is $u\in L({\bf g})$ such that $W[2]_{I({\bf g})}\subseteq\mathsf{var}(B_{u})$.
Proof.
It follows from the combination of Lemma 3 that either the statement of the present lemma is true or $W[2]_{I({\bf g})}\subseteq X$, where $X=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\bigcup_{u\in L_{\bf g}}\mathsf{var}B_{u}$.
First all, we observe that by Lemma 12 $\mathcal{S}(B)|_{\bf g}$ includes at least one satisfying assignment, namely ${\bf h}_{J}[{\bf g}]$ where $J$ is an arbitrary non-empty subset of $I({\bf g})$. Next, we observe that $X$ cannot include all the variables of $\mathcal{S}(B)|_{\bf g}$ as otherwise, assignment all of them with $1$ will falsify $\neg V(G)[2]$. It then follows from Lemma 3 that $\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}\times\mathcal{S}$ where $\mathsf{var}(\mathcal{S})\neq\emptyset$.
By definition, $|W[2]_{I({\bf g})}|\geq 2$. Therefore, if $W[2]_{I({\bf g})}\subseteq X$, then we can partition $X$ into $X_{1}$ and $X_{2}$ each $X_{i}$ has a non-empty intersection with $|W[2]_{I({\bf g})}|$. But then $\mathcal{S}(B)|_{\bf g}=X_{1}^{\{0,1\}}\times X_{2}^{\{0,1\}}\times\mathcal{S}$ thus breaking $W[2]_{I({\bf g})}$ in contradiction to Lemma 13. ∎
In light of Lemma 14, we now define $u({\bf g})$ as the $u\in L({\bf g})$ such that $W[2]_{I({\bf g})}\subseteq\mathsf{var}(B_{u})$.
We proceed demonstrating that for distinct elements ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$, $u({\bf g}_{1})\neq u({\bf g}_{2})$.
Lemma 15.
Let ${\bf g}_{1},{\bf g}_{2}$ be two distinct elements of $\mathcal{F}$. Then $u({\bf g}_{1})\neq u({\bf g}_{2})$.
Proof.
Assume the opposite, that is, $u({\bf g}_{1})=u({\bf g}_{2})$. Let $u_{i}[1]$ be a variable assigned differently by ${\bf g}_{1}$ and ${\bf g}_{2}$. Assume w.l.o.g. that ${\bf g}_{1}(u_{i}[1])=0$ while ${\bf g}_{2}(u_{i}[1])=1$. That is, $i\in I({\bf g}_{2})$.
Let $I=\{i\}$. By Lemma 12, ${\bf h}_{I}[{\bf g}_{2}]$ satisfies $\psi(G)$. Let ${\bf h}=Proj({\bf h}_{I}[{\bf g}_{2}],\mathsf{var}(B_{u({\bf g}_{2})}))$. In particular, note that $w_{i}[2]\in\mathsf{var}({\bf h})$ and is mapped to 00. By Lemma 3, ${\bf h}\in\mathcal{S}(B_{u({\bf g}_{2})})$. Hence, by assumption, ${\bf h}\in\mathcal{S}(B_{u({\bf g}_{1})})$. By another application of Lemma 3, ${\bf g}_{1}\cup{\bf h}$ can be extended to a satisfying assignment of $\psi(G)$. However, this is a contradiction because ${\bf g}_{1}\cup{\bf h}$ maps both $u_{i}[1]$ and $w_{i}[2]$ to zeroes thus falsifying the clause $(u_{i}[1]\vee w_{i}[2])$. ∎
Theorem 9 is now an easy corollary of Lemma 15.
Proof.
(of Theorem 9) It is immediate from Lemma 15 that $|B|\geq\mathcal{F}$. It remains to notice that $\{Proj({\bf g},U[1])|{\bf g}\in\mathcal{F}\}$ includes all possible assignments but all zeroes and all ones. It thus follows that $|\mathcal{F}|\geq 2^{q}-q-2$. ∎
Proof.
(of Theorem10)
Claim 3.
Let $B$ be an obdd obeying an order $\pi$ and let ${\bf g}$ be an assignment to a prefix of $\pi$. Assume that $\emptyset\subset\mathcal{S}(B)|_{\bf g}\subset(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))^{\{0,1\}}$. Then $L({\bf g})$ is a singleton. Let $u({\bf g})$ be the only element of $L({\bf g})$. Then $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B_{u({\bf g})})\times X^{\{0,1\}}$, where $X=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\mathsf{var}(\mathcal{S}(B_{u({\bf g})})$.
Proof.
We observe that $|L({\bf g})|\leq 1$. Indeed, in case of two distinct nodes of $L({\bf g})$, we employ the argument as in the proof of the second statement of Lemma 2 to demonstrate that these nodes must have a $\wedge_{d}$ node as their common ancestor. As an obdd does not have such nodes, this possibility is ruled out. If we assume that $L({\bf g})=\emptyset$ then, taking into account that $\mathcal{S}(B)|_{\bf g}\neq\emptyset$, we conclude that the only node of $B[{\bf g}]$ besides the incomplete decision node is the $true$ sink. However, this means that all possible extensions of ${\bf g}$ are satisfying assignments of $B$ again contradicting our assumptions. We thus conclude that $L({\bf g})$ is indeed a singleton. The remaining part of the claim is immediate from Lemma 3. ∎
Note that Claim 3 formalizes the approach to proving obdd lower bounds that we outlined in Section 3.1. In particular, $u({\bf g})$ can be identified straight away without resorting to unbreakable sets. We thus proceed to define a fooling set $\mathcal{F}_{0}$ and to demonstrate that for every two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}_{0}$, $u({\bf g}_{1})\neq u({\bf g}_{2})$.
Let $\pi$ be an order obeyed by an obdd $B$ representing $\varphi(G)$ such that $M$ crosses $B$. Let $U,W\subseteq V(G)$ be two subsets of $V(G)$ such that $M$ is an induced matching of $G$ between $U$ and $W$ and each vertex of $U$ is ordered by $\pi$ before each vertex of $W$. Let $\pi_{0}$ be the prefix of $\pi$ whose last vertex is the last vertex of $U$ in $\pi$. Let $\mathcal{F}_{0}$ be the set of all assignments over $\pi_{0}$ that assign all variables of $\pi_{0}\setminus U$ by $1$ and assign at least one variable of $U$ by 00. It is not hard to observe that each ${\bf g}\in\mathcal{F}_{0}$ satisfies the premises of Claim 3.
Claim 4.
Let ${\bf g}_{1},{\bf g}_{2}$ be two distinct elements of $\mathcal{F}_{0}$ Then $u({\bf g}_{1})\neq u({\bf g}_{2})$.
Proof.
Assume the opposite. Let $U=\{u_{1},\dots,u_{q}\}$, $W=\{w_{1},\dots,w_{q}\}$ so that $M=\{\{u_{1},w_{1}\},\dots,\{u_{q},w_{q}\}\}$. Assume w.l.o.g. that there is $1\leq i\leq q$ such that ${\bf g}_{1}(u_{i})=0$ while ${\bf g}_{2}(u_{i})=1$.
Let ${\bf h}_{2}$ be an assignment over $V(G)$ such that ${\bf g}_{2}\subseteq{\bf h}_{2}$, ${\bf h}_{2}(w_{i})=0$ and ${\bf h}_{2}(w)=1$ for each other $w\in\pi\setminus\pi_{0}$. We observe that ${\bf h}_{2}$ satisfies $\varphi(G)$. Indeed, assume that ${\bf h}_{2}$ falsifies some $(u\vee v)$. By definition, both $u$ and $v$ belong to $V(M)$ but $\{u,v\}\in E(G)\setminus M$. But this is a contradiction to $M$ being an induced matching.
Let ${\bf h}=Proj({\bf h}_{2},\mathsf{var}(B_{u({\bf g}_{2})}))$. We note that $w_{i}\in\mathsf{var}({\bf h})$ and ${\bf h}(w_{i})=0$ simply by definition of ${\bf h}_{2}$. By Claim 3, ${\bf h}\in\mathcal{S}(B_{u({\bf g}_{2})})$ and hence, by our assumption, ${\bf h}\in\mathcal{S}(B_{u({\bf g}_{1})})$ implying that ${\bf g}_{1}\cup{\bf h}$ is a satisfying assignment of $\varphi(G)$. However, this is a contradiction since ${\bf g}_{1}\cup{\bf h}$ falsifies $(u_{i}\vee v_{i})$
It follows from Claim 4 that $|B|\geq|\mathcal{F}_{0}|$. It remains to notice that the projections of the elements of $\mathcal{F}_{0}|$ to $U$ include all possible assignments to $U$ but one. Hence, $|\mathcal{F}_{0}|\geq 2^{q}-1$.
4 The effect of structuredness on Decision dnnf
An important property of obdd is efficiency of conjunction of two obdds obeying the same order. In particular, if $B_{1}$ and $B_{2}$ are two obdds over the same set of variables and obeying the same order $\pi$ then the function $f(B_{1})\wedge f(B_{2})$ can be represented by an obdd obeying $\pi$ of size $|B_{1}|\cdot|B_{2}|$ and can be computed at time $O(|B_{1}|\cdot|B_{2}|)$. This nice property is lost by fbdd. In fact, it is not hard to demonstrate that if two obdd $B_{1}$ and $B_{2}$ obey different orders then the size an fbdd representing $f(B_{1})\wedge f(B_{2})$ can be exponentially lower-bounded by $|B_{1}|+|B_{2}|$.
dnnf, being a generalization of fbdd, does not enable efficient conjunction. A restriction on dnnf that enables efficient conjunction is structuredness. The purpose of this section is to define two different forms of structuredness for Decision dnnfs and discuss the power of these models.
Recall that a dnnf is just a deMorgan circuit with all the $\wedge$ gates being decomposable. Assume that both $\wedge$ and $\vee$ gates have fan-in $2$. To define the structuredness, we need to define first a variable tree or a vtree.
Definition 16.
A vtree $VT$ over a set $V$ of variables is a rooted binary tree with every non-leaf node having two children and whose leaves are in a bijective correspondence with $V$. The function $\mathsf{var}$ naturally extends to vtrees. In particular, for each $x\in V(VT)$, $\mathsf{var}(VT_{x})$ is the set of variables labelling the leaves of $VT_{x}$.
Definition 17.
Let $D$ be a dnnf A $\wedge_{d}$ node $u$ of $D$ with inputs $u_{1}$ and $u_{2}$ respects a vtree $VT$ if there is a node $x$ of $VT$ with children $x_{1}$ and $x_{2}$ such that $\mathsf{var}(D_{u_{1}})\subseteq\mathsf{var}(VT_{x_{1}})$ and $\mathsf{var}(D_{u_{2}})\subseteq\mathsf{var}(VT_{x_{2}})$. A dnnf $D$ over $V$ respects $VT$ if each node of $D$ respects $VT$ A dnnf $D$ is structured if it respects a vtree.
To apply the structuredness to Decision dnnf, we need to be aware of two equivalent definitions of the model: as a special case of dnnf and as a generalization of fbdd. So far, we have used the latter definition which we referred to as $\wedge_{d}$-fbdd. The former definition states that a Decision dnnf is a dnnf where each $\vee$ node $u$ obeys the following restriction. Let $u_{0}$ and $u_{1}$ be the inputs of $u$. Then they both are $\wedge_{d}$ nodes and there is a variable $x\in\mathsf{var}(D)$ such that one of inputs of $u_{0}$ is an input gate labelled with $\neg x$ and one of the inputs of $u_{1}$ is an input gate labelled with $x$. Then a structured decision dnnf is just a $\wedge_{d}$-fbdd whose representation as a Decision dnnf respects some vtree. As the fpt upper bound of [11], in fact, applies to structured Decision dnnfs, the following holds.
Theorem 16.
A cnf $\varphi$ of primal treewidth at most $k$ can be represented as a structured decision dnnf of size at most $2^{k}\cdot|\varphi|$.
Yet the model is quite restrictive and can be easily ’fooled’ into being xp-sized even with one long clause. In particular, the following has been established in [3].
Theorem 17.
Let $G$ be a graph without isolated vertices. Let $\varphi^{*}(G)$ be a cnf obtained from $\varphi(G)$ by adding a single clause $\neg V$. Then the size of structured Decision dnnf representing $\varphi^{*}(G)$ is at least $\Omega(\textsc{fbdd}(G))$.
In particular, there is an infinite set of positive integers $k$ for each of them there is a class ${\bf G}_{k}$ of graphs of treewidth at most $k$ (and hence the $itw(\varphi^{*}(G))\leq k+1$) and the structured Decision dnnf representation being of size at least. $n^{\Omega(k)}$ where $n=|V(G)|$.
In order to get an additional insight into the ability of $\wedge_{d}$-dnnf to efficiently represent cnfs of bounded incidence treewidth, we introduce an alternative definition of structuredness, this time based on $\wedge_{d}$-fbdd.
Definition 18.
A $\wedge_{d}$-fbdd $B$ respects a vtree $VT$ if each $\wedge_{d}$ node of $B$ respects $VT$. That is, unlike in the case of the ’proper’ structuredness, we do not impose any restrictions on the decision nodes of $B$. We say that $B$ is structured if $B$ respects a vtree $VT$.
A structured $\wedge_{d}$ fbdd is a generalization of fbdd as it does not impose any restriction on a $\wedge_{d}$-fbdd in the absence of $\wedge_{d}$ nodes. As a result the property of efficient conjunction held for structured dnnf does not hold structured $\wedge_{d}$-fbdds.
Remark 7.
Note that, even though Decision dnnf and $\wedge_{d}$-fbdd, their structured versions are not.
We consider structured $\wedge_{d}$-fbdd a very interesting model for further study because it is a powerful special case of $\wedge_{d}$-fbdd which may lead to insight into size lower bounds of $\wedge_{d}$-fbdds for cnf of bounded incidence treewidth. In particular, we do not know whether such cnfs can be represented by fpt-sized structured $\wedge_{d}$-fbdds. We know, however, that the instances that we used so far for deriving xp lower bounds admit an fpt-sized representation as structured $\wedge_{d}$-fbdds. In fact, a much more general result can be established.
Theorem 18.
Let $\varphi$ be a cnf that can be turned into a cnf of primal treewidth at most $k$ by removal of at most $p$ clauses. Then $\varphi$ can be represented as a structured $\wedge_{d}$-fbdd of size at most $O(n^{p}\cdot 2^{k})$ where $n=|\mathsf{var}(\varphi)|$.
In the rest of this section, we provide a proof of Theorem 18.
Let ${\bf g}$ be an assignment to a subset of variables of a cnf $\psi$. Let $\psi[{\bf g}]$ be the cnf obtained from $\psi$ by removal of all the clauses satisfied by ${\bf g}$ and removal of the occurrences of $\mathsf{var}({\bf g})$ from the remaining clauses.
We note that $\psi[{\bf g}]$ is not necessarily the same function as $\psi|_{\bf g}$ as the set set of the variables of the former may be a strict subset of the set of variables of the latter. Yet, it is not hard to demonstrate the following.
Proposition 4.
If $\mathsf{var}(\psi[{\bf g}])=\mathsf{var}(\psi|_{\bf g})$ then $\mathcal{S}(\psi[{\bf g}])=\mathcal{S}(\psi|_{\bf g})$. Otherwise, $\mathcal{S}(\psi|_{\bf g})=\mathcal{S}(\psi[{\bf g}])\times Y^{\{0,1\}}$ where $Y=\mathsf{var}(\psi|_{\bf g})\setminus\mathsf{var}(\psi[{\bf g}])$.
In order to prove Theorem 18, we identify a set ${\bf C}$ of at most $p$ ’long’ clauses whose removal from $\varphi$ results in a cnf of primal treewidth at most $k$. An important part of the resulting structured $\wedge_{d}$-fbdd is representation of ${\bf C}$ by an $O(n^{p})$-sized decision tree. We are now going to define a decision tree and to demonstrate the possibility of such a representation.
Definition 19.
A decision tree is a rooted binary tree where each leaf node is labeled by either $true$ or $false$ and each non-leaf node is labelled with a variable. Moreover, for each non-leaf node one outgoing edge is labelled with $1$, the other with 00. A decision tree must be read-once, that is the same variable does not label two nodes along a directed path. Then each directed path $P$ corresponds to the assignment ${\bf a}(P)$ in the same way as we defined it for $\wedge_{d}$-fbdd.
For a decision tree $DT$, $\mathsf{var}(DT)$ is the set of variables labelinbg the non-leaf nodes of $DT$. We say that $DT$ represents a function $f$ is $\mathsf{var}(DT)\subseteq\mathsf{var}(f)$ and the following two statements hold.
  1. 1.
    Let $P$ be a root-leaf path of $DT$ ending with a node labeled by $true$. Then $f({\bf g})=1$ for each ${\bf g}$ such that $\mathsf{var}({\bf g})=\mathsf{var}(f)$ and ${\bf a}(P)\subseteq{\bf g}$.
  2. 2.
    Let $P$ be a root-leaf path of $DT$ ending with a node labeled by $false$. Then $f({\bf g})=0$ for each ${\bf g}$ such that $\mathsf{var}({\bf g})=\mathsf{var}(f)$ and ${\bf a}(P)\subseteq{\bf g}$.
Lemma 19.
A cnf of $p$ clauses can be represented as a decision tree $DT(\varphi)$ of size $O(n^{p})$ where $n=|\mathsf{var}(\varphi)|$.
Proof.
If $\varphi$ has no clauses then the resulting decision tree consists of a single $true$ node. If $\varphi$ has an empty clause then the decision tree consists of a single $false$ node.
Otherwise, pick a clause $C$ with variables $x_{1},\dots,x_{r}$. We define $DT(\varphi)$ as having a root-leaf path $P$ consisting of vertices $u_{1},\dots,u_{r+1}$ where $u_{1},\dots,u_{r}$ are respectively labelled with $x_{1},\dots,x_{r}$. Furthermore for each $1\leq i\leq r$, the outgoing edge of $u_{i}$ is labelled with 00 if $x_{i}\in C$ and with $1$ if $\neg x_{i}\in C$. Clearly, ${\bf a}(P)$ falsifies $C$, hence we label $u_{r+1}$ with $false$.
Next, for each $1\leq i\leq r$, we introduce a new neighbour $w_{i}$ of $v_{i}$ and, of course, label the edge $(u_{i},w_{i})$ with the assignment of $x_{i}$ as in $C$. Let $P_{i}$ be the path from the root to $w_{i}$ and let ${\bf g}_{i}={\bf g}(P_{i})$ Clearly, ${\bf g}_{i}$ satisfies $C$. Therefore $\varphi[{{\bf g}_{i}}]$ can be expressed as a cnf of at most $p-1$ clauses. Make each $w_{i}$ the root of $DT(\varphi[{{\bf g}_{i}}])$. Clearly, the resulting tree represents $\varphi$. By the induction assumption, the trees $DT(\varphi[{{\bf g}_{i}}])$ are of size at most $n^{p-1}$ and there are at most $|C|\leq n$ such trees. Hence the lemma follows. ∎
Proof.
(of Theorem 18) Let ${\bf C}$ be at most $p$ clauses of $\varphi$ whose removal makes the remaining cnf to be of primal treewidth at most $k$.
Let $DT({\bf C})$ be the decision tree of size $O(n^{p})$ representing ${\bf C}$ guaranteed to exist by Lemma 19. Let $\psi=\varphi\setminus{\bf C}$. Since structured Decision dnnf is a special case of structured $\wedge_{d}$-fbdd, Theorem 16 holds for structured $\wedge_{d}$-dnnf. In fact, it is not hard to establish the following strengthening.
Claim 5.
There is a vtree $VT$ such that the following holds. Let $P$ be a root-leaf path of $DT({\bf C})$. Then there is a structured $\wedge_{d}$-fbdd $B_{P}$ respecting $VT$ and representing $\psi[{\bf a}(P)]$ and of size $O(2^{k}\cdot|var(\psi)|)$.
We construct a structured $\wedge_{d}$-fbdd $B$ representing $\varphi$ of size at most $O(n^{p}2^{k}\cdot|\mathsf{var}(\varphi))$ in the following three stages.
  1. 1.
    For each root-leaf path $P$ of $DT({\bf C})$ that ends at a $true$ sink, remove the $true$ label off the sink and identify the sink with the source of $B_{P}$ as per Claim 5.
  2. 2.
    Contract all the $true$ sinks of the resulting dag into a single $true$ sink and all the $false$ sinks into a single $false$ sink.
  3. 3.
    Let $B^{\prime}$ be the dag obtained as a result of the above two stages. Let $X=\mathsf{var}(\varphi)\setminus\mathsf{var}(B^{\prime})$. If $X=\emptyset$ we set $B=B^{\prime}$. Otherwise, let $x_{1},\dots,x_{a}$ be the variables of $X$. Then $B$ is obtained from $B^{\prime}$ by introducing new nodes $u_{1},\dots,u_{a}$ and setting $u_{a+1}$ to be the source of $B^{\prime}$. For each $1\leq i\leq a$, label $u_{i}$ with $x_{i}$ and introduce two edges from $u_{i}$ to $u_{i+1}$, one labelled with $1$, the oth er labelled with 00.
It is not hard to see that $B$ is indeed a $\wedge_{d}$-fbdd representing $\varphi$ and with the required upper bound on its size. We further note that each $\wedge_{d}$ node of $B$ is a $\wedge_{d}$ node of some $B_{P}$ and hence respects $VT$ by construction.
5 Conclusion
The main purpose of this paper is obtaining better understanding of the power of $\wedge_{d}$-fbdd in terms of the representation of cnfs of bounded incidence treewidth.
For that purpose, we considered several restrictions on $\wedge_{d}$. The first considered model is $\wedge_{d}$-obdd. The xp lower bound for $\wedge_{d}$-obdd representing cnfs of bounded incidence treewidth leads us to the first open question. One motivation for considering $\wedge_{d}$-fbdd for cnf of bounded incidence treewidth is that this would provide an insight into the understanding of complexity of regular resolution for such cnfs. The xp lower bound for $\wedge_{d}$-obbd thus naturally raises a question whether such a lower bound would hold for the corresponding restriction over the regular resolution.
It is well known that a regular resolution refutation of a cnf $\varphi$ can be represented as a read-once branching program whose sinks are labelled with clauses of $\varphi$. The constraint is that the assignment carried by each source-sink path must falsify the clause associated with the sink. Let us say that the a regular resolution refutation is oblivious if the corresponding read-once branching program obeys a linear order of its variables.
Open question 2.
What is the complexity of oblivious regular resolution parameterized by the incidence treewidth?
One important property of obdds is a possibility of efficient conjunction of two obdds obeying the same order: such obdds can be seen as structured dnnfs respecting the same vtree. One important consequence is a polynomial time algorithm for satisfiability testing of conjunction of two obdds obeying the same order. To put it differently, the conjunction of two obdds obeying the same order meets the minimal requirement for being considered a knowledge compilation model. It is thus interesting whether the same holds for conjunction of two $\wedge_{d}$-obdds obeying the same order.
Open question 3.
Let $B_{1}$ and $B_{2}$ be two $\wedge_{d}$ obdds obeying the same order $\pi$. Is there a polynomial time algorithm for testing whether $f(B_{1})\wedge f(B_{2})$ has at least one satisfying assignment?
Finally, we believe that understanding the complexity of structured $\wedge_{d}$-fbdd representing cnfs of bounded incidence treewidth is an important step towards understanding such complexity for $\wedge_{d}$-fbdds. In this context, we will pose a more specific open question related to Theorem 18.
Open question 4.
The upper bound presented in Theorem 18 is fpt in $k$ but xp in $p$. Does an fpt upper bound parameterized by $k+p$ hold for structured $\wedge_{d}$-fbdd representing the class of cnfs as specified in Theorem 18?
Acknowledgement
We would like to thank Florent Capelli for useful feedback.
Appendix A Proof of Lemma 11
Claim 6.
Let $X=(\mathsf{var}(B)\setminus\{x\})\setminus\mathsf{var}(B^{\prime})$. Then $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B^{\prime})\times X^{\{0,1\}}$.
It is immediate from the essentiality assumption that $X=\emptyset$. Therefore, the statement of the lemma is immediate from the claim. It thus remains to prove the claim. The proof is by induction on $|B|$. For $|B|=1$, $\mathsf{var}(B)=\emptyset$, hence the claim is true in a vacuous way.
For $|B|>1$, we assume first that the source $u$ is a decision node labelled by a variable $y$. Further on, assume first that $y=x$. Let $(u,v)$ be the outgoing edge of $u$ labelled with $i$. It is clear by construction that $B^{\prime}=B_{v}$. But then the claim is immediate from Lemma 20. We thus assume that $y\neq x$.
Let ${\bf h}\in\mathcal{S}(B)|_{\bf g}$. Let ${\bf h}_{0}=Proj({\bf h},\mathsf{var}(B^{\prime}))$. We need to demonstrate that ${\bf h}_{0}\in\mathcal{S}(B^{\prime})$.
Let ${\bf a}={\bf h}\cup\{(x,i)\}$. Let $j$ be the assignment of $y$ by ${\bf a}$. Let ${\bf b}={\bf a}\setminus\{(y,j)\}$.
Let $(u,v)$ be the outgoing edge of $u$ labelled by $j$. Let ${\bf b}_{0}=Proj({\bf b},\mathsf{var}(B_{v}))$. By Lemma 20, ${\bf b}_{0}\in\mathcal{S}(B_{v})$. Let ${\bf c}_{0}={\bf b}_{0}\setminus\{(x,i)\}$. Let ${\bf c}_{1}=Proj({\bf c}_{0},\mathsf{var}(B^{\prime}_{v}))$. Let ${\bf h}_{1}={\bf c}_{1}\cup\{(y,j)\}$.
On the one hand ${\bf h}_{1}\subseteq{\bf h}$. On the other hand $\mathsf{var}(B^{\prime}_{v})\subseteq var(B^{\prime})$ simply by construction. We conclude that ${\bf h}_{1}\subseteq{\bf h}_{0}$. By Lemma 20 applied to $B^{\prime}$ and $B^{\prime}_{v}$, we conclude that an arbitrary extension of ${\bf h}_{1}$ to $\mathsf{var}(B^{\prime})$ is an element of $\mathcal{B^{\prime}}$. It follows that ${\bf h}_{0}\in\mathcal{S}(B^{\prime})$.
For the opposite direction, pick $j\in\{0,1\}$ Let $v$ be the outgoing edge of $u$ in $B^{\prime}$ such that $(u,v)$ is labelled with $j$. Let ${\bf q}=\{(x,i),(y,j)\}$. Let $Y=(\mathsf{var}(B)\setminus\{x,y\})\setminus\mathsf{var}(B^{\prime}_{v})$. Let ${\bf c}\in\mathcal{S}(B^{\prime}_{v})$ and let ${\bf y}\in Y^{\{0,1\}}$. It is sufficient to prove that ${\bf c}\cup{\bf y}\in\mathcal{S}(B)|_{\bf q}$.
The set $Y$ can be weakly partitioned (meaning that some of partition classes may be empty) into $Y_{0}=(\mathsf{var}(B_{v})\setminus\mathsf{var}({\bf g}))\setminus\mathsf{var}(B^{\prime}_{v})$ and $Y_{1}=(\mathsf{var}(B)\setminus\{y\})\setminus var(B_{v})$. Let ${\bf y}_{0}=Proj({\bf y},Y_{0})$ and ${\bf y}_{1}=Proj({\bf y},Y_{1})$. By the induction assumption, ${\bf c}\cup{\bf y}_{0}\in\mathcal{S}(B_{v})|_{\bf g}$. Let ${\bf c}_{0}={\bf c}\cup{\bf y}_{0}\cup{\bf g}$. By Lemma 20, ${\bf c_{0}}\cup{\bf y}_{1}\in\mathcal{S}(B)|_{\{(y,j)\}}$. We note that ${\bf c}_{0}\cup{\bf y}_{1}={\bf c}\cup{\bf g}$. Hence ${\bf c}\in\mathcal{S}(B)|_{\bf q}$ as required.
It remains to assume that $u$ is a conjunction node. Let $u_{1}$ and $u_{2}$ be the children of $u$. For the sake of brevity, we denote $B_{u_{1}}$ and $B_{u_{2}}$ by $B_{1}$ and $B_{2}$ respectively. Assume w.l.o.g. that $x\in\mathsf{var}(B_{1})$. By Lemma 21
$\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B_{1})|_{\bf g}\times\mathcal{S}(B_{2})$ (1)
By applying the induction assumption to $B_{1}$, (1) is transformed into
$\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B^{\prime}_{1})\times\mathcal{S}(B_{2})$ (2)
By construction, $B^{\prime}$ is obtained by transforming $B_{1}$ int $B^{\prime}_{1}$. Therefore, by Lemma 21 used with the empty assignment
$\mathcal{S}(B^{\prime})=\mathcal{S}(B^{\prime}_{1})\times\mathcal{S}(B_{2})$ (3)
The right-hand parts of (2) and (3) are the same so the left-hand parts are implying the statement for the considered case.
The proof of the lemma is now complete.
Appendix B Proof of Lemma 3
Lemma 20.
Let $B$ be a $\wedge_{d}$-obdd. Let ${\bf g}$ be an assignment. Suppose that the source node $u$ of $B$ is a decision node associated with a variable $x$. Let $i\in\{1,2\}$ be such that $(x,i)\in{\bf g}$. Let $(u,v)$ be the outgoing edge of $u$ labelled with $i$. Let ${\bf g}_{0}={\bf g}\setminus\{(x,i)\}$. Let $X=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\mathsf{var}(B_{v})$. Then $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B_{v})|_{{\bf g}_{0}}\times X^{\{0,1\}}$
Proof.
Let ${\bf h}\in\mathcal{S}(B)|_{\bf g}$. Let ${\bf a}={\bf g}\cup{\bf h}$. Then ${\bf a}\in\mathcal{S}(B)$. This means that there is ${\bf b}\in\mathcal{A}(B)$ such that ${\bf b}\subseteq{\bf a}$.
By construction,$(x,i)\in{\bf b}$ and ${\bf b}_{0}={\bf b}\setminus\{(x,i)\}$ is an element of $\mathcal{A}(B_{v})$. Hence, $(x,i)\in{\bf a}$. Let ${\bf a}_{0}={\bf a}\setminus\{(x,i)\}$. Clearly, $\mathsf{var}({\bf b}_{0})\subseteq\mathsf{var}(B_{v})\subseteq\mathsf{var}({\bf a}_{0})$. Let ${\bf c}_{0}=Proj({\bf a},var(B_{v}))$. We conclude that ${\bf c}_{0}\in\mathcal{S}(B_{v})$. Let ${\bf d}={\bf c}_{0}\setminus{\bf g}_{0}$. It follows that ${\bf d}\in\mathcal{S}(B_{v})|_{{\bf g}_{0}}$.
We note that ${\bf h}={\bf a}_{0}\setminus{\bf g}_{0}$ and hence ${\bf d}\subseteq{\bf h}$. Let ${\bf x}={\bf h}\setminus{\bf d}$. Then $\mathsf{var}({\bf x})=(\mathsf{var}(B)\setminus\mathsf{var}(g))\setminus(\mathsf{var}(B_{v})\setminus\mathsf{var}({\bf g}_{0})=(\mathsf{var}(B)\setminus\mathsf{var}(g))\setminus\mathsf{var}(B_{v})=X$. We conclude that ${\bf h}\in\mathcal{S}(B_{v})|_{{\bf g}_{0}}\times X^{\{0,1\}}$.
Conversely, let ${\bf d}\in\mathcal{S}(B_{v})|_{{\bf g}_{0}}$ and let ${\bf x}\in X^{\{0,1\}}$. Let ${\bf h}={\bf d}\cup{\bf x}$ and let ${\bf a}={\bf g}\cup{\bf h}$. We need to prove that ${\bf a}\in\mathcal{S}(B)$. Let ${\bf g}^{\prime}_{0}=Proj({\bf g}_{0},\mathsf{var}(B_{v}))$. Let ${\bf c}_{0}={\bf g}^{\prime}_{0}\cup{\bf d}$. Then ${\bf c}_{0}\in\mathcal{S}(B_{v})$. Consequently, there is ${\bf b}_{0}\subseteq{\bf c}_{0}$ such that ${\bf b}_{0}\in\mathcal{A}(B_{v})$. Let ${\bf b}={\bf b}_{0}\cup\{(x,i)\}$. By construction, ${\bf b}\in\mathcal{A}(B)$. As ${\bf b}\subseteq{\bf c}_{0}\cup\{(x,i)\}\subseteq{\bf a}$, we conclude that ${\bf a}\in\mathcal{S}(B)$. ∎
Lemma 21.
Let $B$ be a $\wedge_{d}$-obdd. Let ${\bf g}$ be an assignment. Assume that the source $u$ of $B$ is a conjunction node. Let $u_{1}$ and $u_{2}$ be children of $u$. Let $V=\mathsf{var}(B)$, $V_{1}=\mathsf{var}(B_{u_{1}})$ and $V_{2}=\mathsf{var}(B_{u_{2}})$. Further on, let ${\bf g}_{1}=Proj({\bf g},V_{1})$, ${\bf g}_{2}=Proj({\bf g},V_{2})$. Then $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B_{u_{1}})|_{\bf g_{1}}\times\mathcal{S}(B_{u_{2}})|_{\bf g_{2}}$.
Proof.
Let ${\bf h}\in\mathcal{S}(B)|_{\bf g}$. Let ${\bf a}={\bf g}\cup{\bf h}$. It follows that there is ${\bf b}\subseteq{\bf a}$ such that ${\bf b}\in\mathcal{A}(B)$.
For each $i\in\{1,2\}$, let ${\bf a}_{i}=Proj({\bf a},V_{i})$ and let ${\bf b}_{i}=Proj({\bf b},V_{i})$. By construction, for each $i\in\{1,2\}$, ${\bf b}_{i}\in\mathcal{A}(B_{u_{i}})$ and hence ${\bf a}_{i}\in\mathcal{S}(B_{u_{i}})$. For each ${\bf a}_{i}$, let ${\bf h}_{i}={\bf a}_{i}\setminus{\bf g}_{i}$. Clearly, ${\bf h}={\bf h}_{1}\cup{\bf h}_{2}$ and ${\bf h}_{i}\in\mathcal{S}(B_{u_{i}})|_{{\bf g}_{i}}$ for each $i\in\{1,2\}$. Hence, ${\bf h}\in\mathcal{S}(B_{u_{1}})|_{\bf g_{1}}\times\mathcal{S}(B_{u_{2}})|_{\bf g_{2}}$.
Conversely, let ${\bf h}_{i}\in\mathcal{S}(B_{u_{i}})|_{{\bf g}_{i}}$ for each $i\in\{1,2\}$. Let ${\bf h}={\bf h}_{1}\cup{\bf h}_{2}$ and let ${\bf a}={\bf g}\cup{\bf h}$. We need to demonstrate that ${\bf a}\in\mathcal{S}$. For each $i\in\{1,2\}$, let ${\bf a_{i}}={\bf g}_{i}\cup{\bf h}_{i}$. By assumption, ${\bf a}_{i}\in\mathcal{S}(B_{u_{i}})$ for each $i\in\{1,2\}$ implying existence of ${\bf b}_{1}\subseteq{\bf a}_{1}$ and ${\bf b}_{2}\subseteq{\bf a}_{2}$ such that ${\bf b}_{1}\in\mathcal{A}(B_{u_{1}})$ and ${\bf b}_{2}\in\mathcal{A}(B_{u_{2}})$. Let ${\bf b}={\bf b}_{1}\cup{\bf b}_{2}$. By construction, ${\bf b}\in\mathcal{A}(B)$. As clearly, ${\bf b}\subseteq{\bf a}$, we conclude that ${\bf a}\in\mathcal{S}(B)$. ∎
Lemma 22.
Let $B$ be a $\wedge_{d}$-obdd obeying an order $\pi$. Let ${\bf g}$ be an assignment to a prefix of $\pi$. Suppose that the source node $u$ of $B$ is a decision node associated with a variable $x$. Let $i\in\{0,1\}$ be such that $(x,i)\in{\bf g}$ Let $(u,v)$ be the outgoing edge of $u$ labelled with $i$. Let ${\bf g}_{0}={\bf g}\setminus\{(x,i)\}$. Then $B_{v}$ obeys $\pi_{0}=\pi\setminus\{x\}$, ${\bf g}_{0}$ is an assignment to a prefix of $\pi_{0}$ and $B[{\bf g}]$ is obtained from $B_{v}[{\bf g_{0}}]$ by adding the edge $(u,v)$ labelled with $i$
Proof.
$B_{v}$ obeys $\pi$ but does not contain $x$ hence clearly $B_{v}$ obeys $\pi_{0}$. Clearly, removal of $x$ from a prefix of $\pi$ results in a prefix of $\pi_{0}$. Therefore, ${\bf g}_{0}$ assigns a prefix of $\pi_{0}$.
It is not hard to see that $B[{\bf g}]$ can be obtained as follows.
  1. 1.
    Remove the outgoing edge of $u$ labelled with $2-i$.
  2. 2.
    Remove all the nodes besides $u$ and those in $B_{v}$.
  3. 3.
    Inside $B_{v}$ remove all the edges whose tails are labelled with variables of ${\bf g}$ but the labels on the edges are opposite to their respective assignments in ${\bf g}$. Then remove all the nodes all the nodes of $B_{v}$ that are not reachable from $v$ in the resulting graph.
It is not hard to see that the graph resulting from the above algorithm is the same as when we first compute $B_{v}[{\bf g}_{0}]$ and then add $(u,v)$. ∎
Corollary 1.
With the notation and premises as of Lemma 22, $L_{B}({\bf g})=L_{B_{v}}({\bf g}_{0})$
Proof.
By construction of $B[{\bf g}]$ as specified in Lemma 22, it is immediate that $B[{\bf g}]$ and $B_{v}[{\bf g}_{0}]$ have the same set of complete decision nodes. Moreover, as the construction does not introduce new paths between the complete decision nodes nor removes an existing one, minimal complete decision node of $B_{v}[{\bf g_{0}}]$ remains minimal in $B[{\bf g}]$ and no new minimal complete decision node appears in $B[{\bf g}]$. ∎
We now introduce special notation for the variable $X$ as in Lemma 3. In particular, we let $X_{B}({\bf g})=(\mathsf{var}(B)\setminus\mathsf{var}(g))\setminus\bigcup_{w\in L_{B}({\bf g})}\mathsf{var}(B_{w})$.
Corollary 2.
With the notation and premises as of Lemma 22, $X_{B}({\bf g})=((\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\mathsf{var}(B_{v}))\cup X_{B_{v}}({\bf g}_{0})$.
Proof.
By Corollary 1
$X_{B}({\bf g})=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\bigcup_{w\in L_{B_{v}}({\bf g}_{0})}\mathsf{var}(B_{w})$ (4)
As $var(B_{v})\setminus\mathsf{var}({\bf g}_{0})\subseteq\mathsf{var}(B)\setminus\mathsf{var}({\bf g})$, we conclude that $X_{B_{v}}({\bf g}_{0})\subseteq X_{B}({\bf g})$. But what are $X_{B}({\bf g})\setminus X_{B_{v}}({\bf g}_{0})$? They are all the variables of $\mathsf{var}(B)\setminus\mathsf{var}(B_{v})$ but $\mathsf{var}({\bf g})$. Hence the corollary follows.
Lemma 23.
Let $B$ be a $\wedge_{d}$-obdd obeying an order $\pi$. Let ${\bf g}$ be an assignment to a prefix of $\pi$. Suppose that the source $u$ of $B$ is a conjunction node with children $u_{1}$ and $u_{2}$. Let $V=\mathsf{var}(B)$, $V_{1}=\mathsf{var}(B_{u_{1}})$ and $V_{2}=\mathsf{var}(B_{u_{2}})$. Further on, let $\pi_{1}=\pi[V_{1}]$, $\pi_{2}=\pi[V_{2}]$ (where $\pi[V_{i}]$ is the suborder of $\pi$ induced by $V_{i}$), ${\bf g}_{1}=Proj({\bf g},V_{1})$, ${\bf g}_{2}=Proj({\bf g},V_{2})$. Then, for each $i\in\{1.2\}$, $B_{u_{i}}$ is a $\wedge_{d}$-obdd obeying $\pi_{i}$ and ${\bf g}_{i}$ s an assignment over a prefix of $\pi_{i}$. Furthermore, $B[{\bf g}]$ is obtained from $B_{u_{1}}[{\bf g}_{1}]$ and $B_{u_{2}}[{\bf g}_{2}]$ by adding a conjunction node $u$ with children $u_{1}$ and $u_{2}$.
Proof.
For each $i\in\{1,2\}$, $B_{u_{i}}$ obeys $\pi$. Hence, $B_{u_{i}}$ obeys the suborder of $\pi$ induced by $\mathsf{var}(B_{u_{i}})$. Assume that say $\mathsf{var}({\bf g}_{1})$ do not form a prefix of $\pi_{1}$. Then there is $x\in\mathsf{var}({\bf g}_{1})$ and $y\in\pi_{1}\setminus\mathsf{var}({\bf g}_{1})$ such that $y<_{\pi_{1}}x$. Then $y<_{\pi}x$ and $y\in\pi\setminus\mathsf{var}({\bf g})$ in contradiction to $\mathsf{var}({\bf g})$ being a prefix of $\pi$.
It is not hard to see that, in the considered case, the second stage of obtaining the alignment can be reformulated as removal of nodes reachable from both $u_{1}$ and $u_{2}$. Now, let us obtain $B_{u_{1}}[{\bf g}_{1}]\cup B_{u_{2}}[{\bf g}_{2}]$ as follows.
  1. 1.
    Carry out the edge removal in both $B_{u_{1}}$ and $B_{u_{1}}$. Clearly, this is equivalent to edges removal from $B$ to obtain $B[{\bf g}]$.
  2. 2.
    Carry out node removal separately for $B_{u_{1}}[{\bf g_{1}}]$ and $B_{u_{2}}[{\bf g_{2}}]$ and then perform the graph union operation. Clearly, the nodes removed (apart from the source) will be precisely those that are not reachable from $u_{1}$ in $B_{u_{1}}$ and in $u_{2}$ in $B_{u_{2}}$ after the removal of edges.
Then, adding the source as specified will result in $B[{\bf g}]$ as required. ∎
Corollary 3.
With the notation and premises as in Lemma 23, $L_{B}({\bf g})=L_{B_{1}}({\bf g}_{1})\cup L_{B_{2}}({\bf g}_{2})$.
Proof.
It is clear from the construction of $B[{\bf g}]$ by Lemma 23 that the set of complete decision nodes of $B[{\bf g}]$ is the union of the sets of such nodes of $B_{1}[{\bf g}_{1}]$ and $B_{2}[{\bf g}_{2}]$. Let $w\in L_{B}({\bf g}]$. Assume w.l.o.g. that $w$ is a node of $B_{1}[{\bf g}_{1}]$ as the minimality cannot be destroyed by moving from a graph to a subgraph, we conclude that $w\in L_{B_{1}}({\bf g}_{1})$.
Conversely, let $w\in L_{B_{1}}({\bf g}_{1})\cup L_{B_{2}}({\bf g}_{2})$. We assume w.l.o.g. that $w\in L_{B_{1}}({\bf g}_{1})$ We note that the construction of $B[{\bf g}]$ by Lemma 23 does not add new edges between vertices of $B_{1}[{\bf g}_{1}]$ and a path from a decision node of $B_{2}[{\bf g}_{2}]$ to a decision node of $B_{1}[{\bf g}_{1}]$ is impossible due to the decomposability of $u$. Therefore, we conclude that $w\in L_{B}({\bf g}]$. ∎
Corollary 4.
With the notation and premises as in Lemma 23, $X_{B}({\bf g})=X_{B_{1}}({\bf g}_{1})\cup X_{B_{2}}({\bf g}_{2})$
Proof.
Due to the decomposability of $u$, $X_{B}({\bf g})$ is the disjoint union of $X_{1}=X_{B}({\bf g})\cap\mathsf{var}(B_{1})$ and $X_{2}=X_{B}({\bf g})\cap\mathsf{var}(B_{2})$. We demonstrate that for each $i\in\{1,2\}$ $X_{i}=X_{B_{i}}({\bf g}_{i})$. The argument is symmetric for both possible values of $i$, so we assume w.l.o.g that $i=1$.
By definition $X_{1}=Y_{0}\setminus Y_{1}$ where $Y_{0}=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\cap\mathsf{var}(B_{1})=\mathsf{var}(B_{1})\setminus\mathsf{var}({\bf g}_{1})$ and $Y_{1}=(\bigcup_{w\in L_{B}({\bf g})}\mathsf{var}(B_{w}))\cap\mathsf{var}(B_{1})$ By decomposability of $u$ and Corollary 3, $Y_{1}=\bigcup_{w\in L_{B_{1}}({\bf g}_{1})}\mathsf{var}(B_{w})$. Substituting the obtained expressions fr $Y_{0}$ and $Y_{1}$ into $Y_{0}\setminus Y_{1}$, we obtain the definition of $X_{B_{1}}({\bf g}_{1})$. ∎
Proof.
(of Lemma 3) By induction of $|\mathsf{var}(B)|$. Suppose that $|\mathsf{var}(B)|=0$. Then $B$ consists of a singe sink node. The sink cannot be a $false$ node because otherwise, we get a contradiction with $|\mathcal{S}(B)|_{\bf g}\neq\emptyset$. Thus the sink is a $true$ node and hence $\mathcal{S}(B)=\{\emptyset\}$, any alignment of $B$ is $B$ itself and for any assignment ${\bf g}$, $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B)$. Clearly, $X_{B}({\bf g})=\emptyset$ and $\mathcal{S}(B)|_{\bf g}=\emptyset^{0,1}$ as required by the lemma.
Assume now that $\mathsf{var}(B)|>0$ and $u$ be the source node of $B$.
Assume that $u$ is a decision node. Let $x$ be the variable associated with $u$. If $\mathsf{var}({\bf g})\cap\mathsf{var}(B)=\emptyset$ then $B[{\bf g}]=B=B_{u}$, that is $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B_{u})$. As $u$ is the only minimal decision node of $B$, the lemma holds in this case.
Otherwise, we observe that $x\in\mathsf{var}({\bf g})$. Indeed, let $y\in\mathsf{var}({\bf g})\cap\mathsf{var}(B)$. This means that $B$ has a decision node $w$ associated with a variable $y$. If $w=u$ then $y=x$ and we are done. Otherwise, as $u$ is the source node $B$ has a path from $u$ to $w$. It follows that $x$ precedes $y$ in $\pi$. As $\pi_{0}$ is a prefix including $y$, it must also include $x$.
Let $(x,i)$ be an element of ${\bf g}$. Let $(u,v)$ be the outgoing edge of $u$ labelled with $i$. Let ${\bf g}_{0}={\bf g}\setminus\{x,i\}$. Let $Y=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\mathsf{var}(B_{v})$. We note that by Lemma 20, $\mathcal{S}(B)|_{\bf g}\neq\emptyset$ implies that $\mathcal{S}(B_{v})({\bf g}_{0})\neq\emptyset$. Therefore, we may apply the induction assumption to $B_{v}[{\bf g}_{0}]$.
Assume that $L_{B_{v}}({\bf g}_{0})\neq\emptyset$. Combining Lemma 20 and the induction assumption, we observe that
$\mathcal{S}(B)|_{\bf g}=\prod_{w\in L_{B_{v}}({\bf g}_{0})}\mathcal{S}(B_{w})\times(X_{B_{v}}({\bf g}_{0})\cup Y)^{0,1}$ (5)
The lemma is immediate from application of Corollaries 1 and 2 to (5).
If $L_{B_{v}}({\bf g}_{0})=\emptyset$ then the same reasoning applies with $\mathcal{S}(B)|_{\bf g}=(X_{B_{v}}({\bf g}_{0})\cup Y)^{0,1}$ used instead (5).
It remains to assume that $u$ is a conjunction node. Let $u_{1}$ and $u_{2}$ be the children of $u$. For $i\in\{1,2\}$, let $B_{i}=B_{u_{i}}$, $V_{i}=\mathsf{var}(B)\cap\mathsf{var}(B_{i})$, ${\bf g}_{i}=Proj({\bf g},V_{i})$, $\pi_{i}=\pi[V_{i}]$. By Lemma 21, for each $i\in\{1,2\}$, $B_{i}$ obeys $\pi_{i}$, ${\bf g}_{i}$ is an assignment over a prefix of $\pi_{i}$, and $\mathcal{S}(B_{i})|_{{\bf g}_{i}}\neq\emptyset$. Therefore, we can apply the induction assumption to each $B_{i}[{\bf g_{i}}]$.
Assume that both $L_{B_{1}}({\bf g}_{1})$ and $L_{B_{2}}({\bf g}_{2})$ are non-empty. Therefore, by Lemma 21 combined with the induction assumption, we conclude that
$\mathcal{S}(B)=\prod_{i\in\{1,2\}}(\prod_{w\in L_{B_{i}}({\bf g}_{i})}(\mathcal{S}(B_{w}))\times X_{B_{i}}({\bf g}_{i})^{0,1})$ (6)
The lemma is immediate by applying Corollary 3 and Corollary 4 to (6).
Assume that, say $L_{B_{1}}({\bf g}_{1})\neq\emptyset$ while $L_{B_{2}}({\bf g}_{2})=\emptyset$. Then the same reasoning applies with $\mathcal{S}(B)=\prod_{w\in L_{B_{1}}({\bf g}_{1})}(\mathcal{S}(B_{w}))\times X_{B_{1}}({\bf g}_{1})^{0,1}\times X_{B_{2}}({\bf g}_{2})$ being used instead (6). Finally, if $L_{B_{1}}({\bf g}_{1})=L_{B_{2}}({\bf g}_{2})=\emptyset$ then we use the same argumentation with $\mathcal{S}(B)=X_{B_{1}}({\bf g}_{1})\times X_{B_{2}}({\bf g}_{2})$ instead (6). ∎