### v2

# On complexity of restricted fragments of Decision DNNF

Abstract
Decision dnnf (a.k.a. $\wedge_{d}$-fbdd) is an important special case of Decomposable Negation Normal Form (dnnf), a landmark knowledge compilation model. Like other known dnnf restrictions, Decision dnnf admits fpt sized representation of cnfs of bounded primal treewidth. However, unlike other restrictions, the complexity of representation for cnfs of bounded incidence treewidth is wide open.
In [5], we resolved this question for two restricted classes of Decision dnnf that we name $\wedge_{d}$-obdd and Structured Decision dnnf. In particular, we demonstrated that, while both these classes have fpt-sized representations for cnfs of bounded primal treewidth, they need xp-size for representation of cnfs of bounded incidence treewidth.
In the main part of this paper we carry out an in-depth study of the $\wedge_{d}$-obdd model. We formulate a generic methodology for proving lower bounds for the model. Using this methodology, we reestablish the xp lower bound provided in [5]. We also provide exponential separations between fbdd and $\wedge_{d}$-obdd and between $\wedge_{d}$-obdd and an ordinary obdd. The last separation is somewhat surprising since $\wedge_{d}$-fbdd can be quasipolynomially simulated by fbdd.
We also study the complexity of Apply operation for $\wedge_{d}$-obdd. We demonstrate that, in general, the Apply operation leads to exponential blow up of the resulting model. We identify a special restricted case where the Apply operation can be carried out efficiently. In order to define the special case, we consider a subclass of $\wedge_{d}$-obdd embeddable into a linear order and introduce a novel concept of irregularity index measuring how far the model from an ordinary obdd. We demonstrate that the efficiency of Apply for the cases where two models are embeddable into the same order with low irregularity indices.
In the remaining part of the paper, we introduce a relaxed version of Structured Decision dnnf that we name Structured $\wedge_{d}$-fbdd. In particular, we explain why Decision dnnf is equivalent to $\wedge_{d}$-fbdd but their structured versions are distinct. We demonstrate that this model is quite powerful for cnfs of bounded incidence treewidth: it has an fpt representation for cnfs that can be turned into ones of bounded primal treewidth by removal of a constant number of clauses (while for both $\wedge_{d}$-obdd and Structured Decision dnnf an xp lower bound is triggered by just two long clauses).
We conclude the paper with an extensive list of open questions.
1 Introduction
1.1 The main objective
Decomposable negation normal forms (dnnfs) [7] is a model for representation of Boolean functions that is of a landmark importance in the area of (propositional) Knowledge compilation (kc). The difference of the model from deMorgan circuits is that all the $\wedge$-gates are decomposable meaning, informally, that two distinct inputs of the gate cannot be reached by the same variable. We will refer to such gates as $\wedge_{d}$-gates. Another way to define dnnfs is as a Non-deterministic read-once branching program ($1$-nbp) equipped with $\wedge_{d}$-nodes (in the context of branching programs, we refer to gates as nodes). This latter view is useful for understanding that (from the univariate perspective, that is when the size bounds depend on the number $n$ of input variables) dnnfs are not very strong compared to $1$-nbps since the latter can quasipolynomially simulate the former [18]. In particular, a class of Boolean functions requiring for its presentation $1$-nbps exponentially large in the number of variables, also requires exponentially large dnnfs.
The strength of $\wedge_{d}$-gates becomes distinctive when we consider dnnfs from the multivariate perspective. In particular, cnfs of primal treewidth at most $k$ have fpt presentation as a dnnf [7] but, in general require an xp-sized representation as a $1$-nbp. [19]. The fpt upper bound as above turned out to be robust in the sense that it is preserved by all known restrictions of dnnfs that retain $\wedge_{d}$-gates. In light of this, it is natural to investigate if fpt upper bounds hold for the incidence treewidth.
Among all known dnnf restrictions using $\wedge_{d}$-gates, there are two that are minimal in the sense that every other restriction is a generalization of one of them. One such minimal restriction is a Sentential Decision Diagram [8] that admits fpt presentation for a much more general class than cnfs of bounded incidence treewidth [4]. The other restriction is Decision dnnf [10, 14] where the parameterized complexity of representation of cnfs of bounded incidence treewidth is an open question formally stated below.
Open question 1.
Are there a function $f$ and a constant $a$ so that a cnf of incidence treewidth at most $k$ can be represented as a Decision dnnf of size at most $f(k)\cdot n^{a}$?
To understand the reason of difficulty of Open Question 1, it is convenient to view Decision dnnfs as Free Binary Decision Diagrams (fbdds) equipped with $\wedge_{d}$-nodes [3]. Throughout this paper, we will refer to this equivalent representation as $\wedge_{d}$-fbdd. This representation clearly illustrates that the $\vee$-gates of Decision dnnf are only capable to ‘guess’ the value of variables associated with them. However, it is not clear how to handle the guessing of satisfaction/non-satisfaction of clauses.
The main objective of the research presented in this paper is to make a progress towards resolution of Open Question 1. In the next subsection of the Introduction, we state our results, the motivation is discussed in Subsection 1.3, the structure of the rest of the paper is discussed in Subsection 1.4.
1.2 Our results
In this paper we tackle the Open Question 1 by restricting it to special cases of $\wedge_{d}$-fbdd. We consider two restrictions: obliviosness and structuredness.
Arguably, the best known restricted class of fbdd is Ordered Binary Decision Diagram (obdd). Applying this restriction to $\wedge_{d}$-fbdds, we obtain $\wedge_{d}$-obdds or, to put it differently, obdds equipped with $\wedge_{d}$-nodes. Arguing as in [10], we conclude that $\wedge_{d}$-obdd represents the trace of a dpll model counter with component analysis and fixed variable ordering.
To the best of our knowledge the $\wedge_{d}$-obdd was first introduced in [5]. In [13] the model appears under the name $OBDD[\wedge]_{\prec}$ where $\prec$ is an order over a chain.
While $\wedge_{d}$-obdd are known to have fpt representation [15] parameterized by the primal treewidth of the input cnf, (as noted in [5] the construction in [15] is, in fact, structured $\wedge_{d}$-obdd), it is shown in [5] that representation of cnfs parameterized by the bounded incidence treewidth requires an xp-sized representation as $\wedge_{d}$-obdds.
In this paper we introduce a generic methodology for proving lower bounds for $\wedge_{d}$-obdds. The key part of this methodology is a novel concept of an ($\wedge_{d}$-obdd-) fooling set $\mathcal{F}$ of assignments defined for a Boolean function $f$ and a permutation $\pi$ of $\mathsf{var}(f)$, the set of variables of $f$. We demonstrate that if $B$ is a $\wedge_{d}$-obdd representing function $f$ and respecting the order $\pi$ then $|B|\geq|\mathcal{F}|$. Thus proving large lower bounds reduces to establishing existence of a large fooling set.
We use the above methodology to obtain two lower bounds for $\wedge_{d}$-obdds. First, we restate the xp lower bound proof of [5]. Second, we demonstrate an exponential separation between fbdd and $\wedge_{d}$-obdd. In particular, we introduce a class of cnfs representable by poly-sized fbdd and prove an exponential lower bound for representation of this class of cnfs as $\wedge_{d}$-obdds.
We also demonstrate an exponential separation between $\wedge_{d}$-obdds and obdds.
An important property of obdd is an efficient implementation of Apply (conjunction) operation. In particular, let $B_{1}$ and $B_{2}$ be two obdds over the same set of variables and respecting the same order that represent functions $f(B_{1})$ and $f(B_{2})$ respectively. Then the function $f(B_{1})\wedge f(B_{2})$ can be represented by an obdd of size $|B_{1}|\times|B_{2}|$. It is thus natural to study whether $\wedge_{d}$-obdd allows efficient Apply operation. We demonstrate that in general this is not the case. In particular, we show that vertical and horizontal lines of an $n\times n$ grid can be represented as linear sized $\wedge_{d}$-obdd respecting the same ‘dictionary’ order. On the other hand, their conjunction is a ‘fully-fledged’ grid and it requires an exponential-sied representation even as a dnnf, let alone an $\wedge_{d}$-obdd.
In light of the above, it is natural to study a restriction of $\wedge_{d}$-obdd that enables efficient implimentation of Apply We introduce a notion of embeddability of a model into a linear order that can be seen as a weak form of structuredness. For Boolean circuits, this notion is known under the name of compatible orders [1], for probabilistic circuits, a similar idea of contiguity was used in [26] for a similar purpose of circuits multiplication.
We introduce a novel notion of an irregularity index that, intutively speaking, shows how far an $\wedge_{d}$-obdd respecting an order $\pi$ is from an obdd embeddable in the same order. We demonstrate that if $B_{1}$ and $B_{2}$ are $\wedge_{d}$-obdds respecting the same order with the respective irregularity indices $k_{1}$ and $k_{2}$ then $f(B_{1})\wedge f(B_{2})$ can be represented as a $\wedge_{d}$-obdd of size $(|B_{1}|\times|B_{2}|)^{O(k_{1}+k_{2})}$. We believe that the notion of irregularity idex gives rise to a new research direction of general interest in the area of kc. We provide the related discussion in the Conclusion section posing a number of conjectures and open questions.
As we stated above, the second restriction on $\wedge_{d}$-fbdds considered in this paper is structuredness. In the context of cnfs of bounded incidence treewidth, structured Decision dnnf is a rather restrictive model. Indeed, it is shown in [5] that taking a monotone $2$-cnf of bounded primal treewidth and adding one clause with negative literals of all the variables requires structured Decision dnnfs representing such cnfs to be xp-sized in terms of their incidence treewidth. By weakening the structurality restriction, we introduce a more powerful model that we name structured $\wedge_{d}$-fbdds.
The starting point of the construction is that transformation from Decision dnnf into $\wedge_{d}$-fbbd leads to an encapsulation effect: a $\vee$-gate and two $\wedge_{d}$-gates implementing Shannon decomposition are replaced by a single decision node. The structured $\wedge_{d}$-fbdd requires that only the $\wedge_{d}$-nodes explicitly present in the model respect a vtree or to put it differently, no constraint is imposed on the $\wedge_{d}$-gates ‘hidden’ inside the decision nodes.
We demonstrate that structured $\wedge_{d}$-fbdd is a rather strong model. In particular, unlike $\wedge_{d}$-obdd and Structured Decision dnnf, it has fpt-sized representation for cnfs that can be turned into cnfs of bounded primal treewidth by removal of a constant number of clauses. We leave it open as to whether Structured $\wedge_{d}$-fbdds admit an fpt representation of cnfs of bounded incidence treewidth.
1.3 Motivation
The motivation of the proposed results is double-fold. First, we make a progress towards resolution of Open Question 1. In this context, we demonstrate that $\wedge_{d}$-fbdd respecting a specific order cannot efficiently represent cnfs of bounded incidence treewidth. To put it differently, if $\wedge_{d}$-fbdd admits an fpt representation of cnfs of bounded incidence treewidth then this representation must heavily depend on absence of any static order of querying of variables.
We also introduce a novel concept of structured $\wedge_{d}$-fbdd, demonstrate its power as compared to the known models and propose understanding the complexity of its representation of cnfs of bounded incidence treewidth as the next step towards resolution of Open Question 1.
The second aspect of motivation is that our results for $\wedge_{d}$-obdd are interesting in their own right. Indeed, in the context of tracing of model counters, the model can be seen as tracing model counters with component analysis and a static ordering of variables. Our results lead to two conclusions about such solvers. First, the xp lower bound for cnfs of bounded incidence treewidth implies that the static ordering does not allow efficient handling of CNFs of bounded incidence treewidth.
Second, it is interesting to see the exponential separation between $\wedge_{d}$-obdd in contrast with quasi-polynomial simulation of $\wedge_{d}$-fbdd by fbdd [3]. This discrepancy demonstrates that, in presence of static ordering, the component analysis is way more important for efficiency of the solver than in case of a dynamical ordering. From the practical perspective, this conclusion means that, in presence of a static ordering, the time overhead spent to component analysis is highly worthwhile.
Furthermore, our results related to Apply operation for $\wedge_{d}$-obdd may be of interest for the area of kc in general beyond the formalisms considered in this paper. For example, the novel notion of irregularity index provides an efficient way of turning a restricted dnnf into a model of the corresponding class of read-once branching programs. Similar transformations are known for Decision dnffs and unrestricted dnnfs and result in a quasi-polynomial blow-up in the size of the resulting read-once branching programs [2, 12]. Moreover, as we show in this paper, the transformation from $\wedge_{d}$-obdd to obdd may even result in an exponential blow-up. However when the transformation is controlled by the irregularity index parameter, it becomes efficient if the value of the parameter is small.
1.4 Structure of the paper
The rest of the paper is organized as follows. The necessary terminology and the related statements are provided in Section 2. The xp lower bound and the separations for $\wedge_{d}$-obdd are provided in Section 3. The complexity of Apply operation for $\wedge_{d}$-obdd is discussed in Section 4. The structuredness applied to $\wedge_{d}$-fbdd is discussed in Section 5. The concluding discussion and the list of related open questions are provided in Section 6.
2 Preliminaries
2.1 Assignments, rectangles, Boolean functions
Let $X$ be a finite set of variables. A (truth) assignment on $X$ is a mapping from $X$ to $\{0,1\}$. The set of truth assignments on $X$ is denoted by $\{0,1\}^{X}$. We reserve lowercase boldface English letters to denote the assignments.
In this paper we will consider several objects for which the set of variables is naturally defined: assignments, Boolean functions, cnfs, and several knowledge compilation models. We use the unified notation $\mathsf{var}(\mathcal{O})$ where $\mathcal{O}$ is the considered object. In particular, the set of variables (that is, the domain) of an assignment $\bf a$} is denoted by $\mathsf{var}({\bf a})$.
We observe that if $\bf a$} and $\bf b$} are assignments and for each $x\in\mathsf{var}({\bf a})\cap\mathsf{var}({\bf b})$, ${\bf a}(x)={\bf b}(x)$ then ${\bf a}\cup{\bf b}$ is an assignment to $\mathsf{var}({\bf a})\cup\mathsf{var}({\bf b})$.
Let $\mathcal{H}$ be a set of assignments, not necessarily over the same set of variables. Let $\mathsf{var}(\mathcal{H})=\bigcup_{{\bf a}\in\mathcal{H}}\mathsf{var}({\bf a})$. We say that $\mathcal{H}$ is uniform if for each $\bf a$$\in\mathcal{H}$, $\mathsf{var}({\bf a})=\mathsf{var}(\mathcal{H})$. For example the set $\mathcal{H}_{1}=\{\{(x_{1},0),(x_{2},1)\}\{(x_{2},0),(x_{3},1)\}\}$ is not uniform and $\mathsf{var}(\mathcal{H}_{1})=\{x_{1},x_{2},x_{3}\}$. On the other hand, $\mathcal{H}_{2}=\{\{(x_{0},1),(x_{2},0),(x_{3},1)\},\\ \{(x_{1},1),(x_{2},1),(x_{3},0)\}\}$ is uniform but, again, $\mathcal{H}_{2}=\{x_{1},x_{2},x_{3}\}$.
We are now going to introduce an operation that is, conceptually, very similar to the Cartesian product. We thus slightly abuse the notation by referring to this operation as the Cartesian product. So, let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be sets of assignments with $\mathsf{var}(\mathcal{H}_{1})\cap\mathsf{var}(\mathcal{H}_{2})=\emptyset$. We denote $\{{\bf a}\cup{\bf b}|{\bf a}\in\mathcal{H}_{1},{\bf b}\in\mathcal{H}_{2}\}$ by $\mathcal{H}_{1}\times\mathcal{H}_{2}$. For example, suppose that $\mathcal{H}_{1}=\{\{(x_{1},1),(x_{2},0)\},\{(x_{2},1)\}\}$ and $\mathcal{H}_{2}=\{\{(x_{3},1)\},\{(x_{3},0),(x_{4},0)\}\}$. Then $\mathcal{H}_{1}\times\mathcal{H}_{2}=\{\{(x_{1},1),(x_{2},0),(x_{3},1)\},\\ \{(x_{1},1),(x_{2},0),(x_{3},0),(x_{4},0)\},\{(x_{2},1),(x_{3},1)\},\{(x_{2},1),(x_{3},0),(x_{4},0)\}\}$. We note that for any set $\mathcal{H}$ of assignments, $\mathcal{H}\times\{\emptyset\}=\mathcal{H}$.
If both $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are uniform and both $\mathsf{var}(\mathcal{H}_{1})$ and $\mathsf{var}(\mathcal{H}_{2})$ are non-empty then $\mathcal{H}_{1}\times\mathcal{H}_{2}$ is called a rectangle. More generally, we say that a uniform set $\mathcal{H}$ of assignments is a rectangle if there are uniform sets $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ of assignments such that $\mathsf{var}(\mathcal{H}_{1})$ and $\mathsf{var}(\mathcal{H}_{2})$ partition $\mathsf{var}(\mathcal{H})$ such that $\mathcal{H}=\mathcal{H}_{1}\times\mathcal{H}_{2}$. We say that $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are witnesses for $\mathcal{H}$. Note that a rectangle can have many possible pairs of witnesses, consider, for example $\{0,1\}^{X}$.
The following notion is very important for proving our results.
Definition 1.
Let $\mathcal{H}$ be a rectangle and let $Y\subseteq\mathsf{var}(\mathcal{H})$. We say that $\mathcal{H}$ breaks $Y$ if there is a pair $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ of witnesses of $\mathcal{H}$ such that both $Y\cap\mathsf{var}(\mathcal{H}_{1})$ and $Y\cap\mathsf{var}(\mathcal{H}_{2})$ are nonempty.
For example $\{0,1\}^{X}$ breaks any disjoint non-empty subset $Y$ of $X$ of size at least $2$. On the other hand, let $\mathcal{H}_{1}=\{\{(x_{1},1),(x_{2},0)\}\}$ and let $\mathcal{H}_{2}=\{\{(x_{3},0),(x_{4},0)\},\{(x_{3},0),(x_{4},1)\},\{(x_{3},1),(x_{4},0)\}\}$. Then $\mathcal{H}=\mathcal{H}_{1}\times\mathcal{H}_{2}$ does not break $\{x_{3},x_{4}\}$.
We can naturally extend $\mathcal{H}_{1}\times\mathcal{H}_{2}$ to the case of more than two sets of assignments, that is $\mathcal{H}_{1}\times\dots\times\mathcal{H}_{m}$ as e.g. $(\mathcal{H}_{1}\times\dots\times\mathcal{H}_{m-1})\times\mathcal{H}_{m}$. The operation is clearly commutative that is we can permute the elements in any order. We also use the notation $\prod_{i\in[m]}\mathcal{H}_{i}$ (recall that $m]=\{1,\dots,m\}$) We notice that if $I_{1},I_{2}$ is a partition of $[m]$ then $\prod_{i\in[m]}\mathcal{H}_{i}=\prod_{i\in I_{1}}\mathcal{H}_{i}\times\prod_{i\in I_{2}}\mathcal{H}_{i}$. Therefore, we observe the following.
Proposition 1.
Suppose that $\mathcal{H}=\prod_{i\in[m]}\mathcal{H}_{i}$, $m\geq 2$ and $\mathcal{H}_{i}$ is uniform for each $i\in[m]$. Further on, assume that $\mathcal{H}$ is not a rectangle breaking a set $Y\subseteq\mathsf{var}(\mathcal{H})$ Then there is $i\in[m]$ such that $Y\subseteq\mathsf{var}(\mathcal{H}_{i})$.
Next, we define the notions of projection and restriction.
Definition 2.
  1. 1.
    Let $\bf g$} be an assignment and let $U$ be a set of variables. The projection $Proj({\bf g},U)$ is the assignment $\bf h$} such that $\mathsf{var}({\bf h})=\mathsf{var}({\bf g})\cap U$ and for each $v\in\mathsf{var}({\bf h})$, ${\bf h}(v)={\bf g}(v)$. Let $\mathcal{H}$ be a set of assignments and let $\bf a$} be an assignment such that $\mathsf{var}({\bf a})\subseteq\mathsf{var}(\mathcal{H})$. Then $\mathcal{H}|_{\bf a}=\{{\bf b}\setminus{\bf a}|{\bf b}\in\mathcal{H},{\bf a}\subseteq{\bf b}\}$. We call $\mathcal{H}|_{\bf a}$ the restriction of $\mathcal{H}$ by $\bf a$}. To extend the notion of restriction to the case where $\mathsf{var}({\bf a})$ is not necessarily a subset of $\mathsf{var}(\mathcal{H})$, we set $\mathcal{H}|_{\bf a}=\mathcal{H}_{Proj({\bf a},\mathsf{var}(\mathcal{H}))}$.
For example, if $\mathcal{H}=\{\{(x_{1},0),(x_{2},0),(x_{3},1)\},\{(x_{1},1),(x_{2},0),(x_{3},1)\},\{(x_{1},1),(x_{2},1),(x_{3},0\}\}$ and ${\bf a}=\{(x_{1},1)\}$ then $\mathcal{H}|_{\bf a}=\{\{(x_{2},0),(x_{3},1)\},\{(x_{2},1),(x_{3},0)\}\}$.
Observe that if $\mathcal{H}$ is uniform then $\mathcal{H}_{\bf a}$ is also uniform.
A literal is a variable $x$ or its negation $\neg x$. For a literal $\ell$, we denote its variable by $\mathsf{var}(\ell)$. Naturally for a set $L$ of literals $\mathsf{var}(L)=\{\mathsf{var}(\ell)|\ell\in L\}$. $L$ is well formed if for every two distinct literals $\ell_{1},\ell_{2}\in L$, $\mathsf{var}(\ell_{1})\neq\mathsf{var}(\ell_{2})$.
If $x\in\mathsf{var}(L)$, we say that $x$ occurs in $L$. Moreover, $x$ occurs positively in $L$ if $x\in L$ and negatively if $\neg x\in L$.
There is a natural correspondence between assignments and well-formed sets of literals. In particular, an assignment $\bf a$} corresponds to a set $L$ of literals over $\mathsf{var}(a)$ such that, for each $x\in\mathsf{var}({\bf a})$, $x\in L$ if ${\bf a}(x)=1$, otherwise $\neg x\in L$. We will use this correspondence to define satisfying assignments for cnfs.
A Boolean function $f$ on variables $X$ is a mapping from $\{0,1\}^{X}$ to $\{0,1\}$. As mentioned above, we denote $X$ by $\mathsf{var}(f)$. Let $\bf a$} be an assignment on $\mathsf{var}(f)$ such that $f({\bf a})=1$. We say that $\bf a$} is a satisfying assignment for $f$.
We denote by $\mathcal{S}(f)$ the set of satisfying assignments for $f$. Conversely, a uniform set $\mathcal{S}$ naturally gives rise to a function $f$ over $\mathsf{var}(f)$ whose set of satisfying assignments is precisely $\mathcal{S}$. Finally, the set of satisfying assignments naturally leads to the definition of a restriction of a function. In particular, for a function $f$ and an assignment $\bf a$}, $f|_{\bf a}$ is a function on $\mathsf{var}(f)\setminus\mathsf{var}({\bf a})$ whose set of satisfying assignments is precisely $\mathcal{S}(f)|_{\bf a}$.
2.2 Models of Boolean functions
We assume the reader to be familiar with the notion of Boolean formulas and in particular, with the notion of conjunctive normal forms (cnfs). For a cnf $\varphi$, $\mathsf{var}(\varphi)$ is the set of variables whose literals occur in the clauses of $\varphi$. For example, let $\varphi=(x_{1}\vee x_{2})\wedge(x_{2}\vee\neg x_{3}\vee\neg x_{4})$. Then $\mathsf{var}(\varphi)=\{x_{1},x_{2},x_{3},x_{4}\}$.
In order to define the function represented by a cnf, it will be convenient to consider a cnf as a set of its clauses and a clause as a set of literals. Let $\bf a$} be an assignment and let $L_{\bf a}$ be the well-formed set of literals over $\mathsf{var}({\bf a})$ corresponding to $\bf a$}. Let $\varphi=\{C_{1},\dots,C_{q}\}$ be a cnf. Then $\bf a$} satisfies a clause $C_{i}$ if $L_{\bf a}\cap C_{i}\neq\emptyset$ and $\bf a$} satisfies $\varphi$ if $\bf a$} satisfies each clause of $\varphi$. The function $f(\varphi)$ represented by $\varphi$ has the set of variables $\mathsf{var}(\varphi)$ and $\mathcal{S}(f(\varphi))$ consists of exactly those assignments (over $\mathsf{var}(\varphi)$) that satisfy $\varphi$. In what follows, we slightly abuse the notation and refer to the function represented by $\varphi$ also by $\varphi$. The correct use will always be clear from the context.
Remark 1.
Let us remark that, for a cnf $\varphi$, $\mathcal{S}(\varphi)$ consists of the satisfying assignment of $\varphi$ over all the variables of $\varphi$ (simplyt because, in this context we consider $\varphi$ a function). On the other if $\varphi$ is viewed as a cnf, it can have much shorter satisfying assignments. For example, $\varphi=(x_{1}\vee x_{2})\wedge(x_{2}\vee\neg x_{3}\vee\neg x_{4})$ is satisfied with $\{(x_{2},1)\}$. It is well known that $\mathcal{S}(\varphi)$ is the set of all extensions (to the whole of $\mathsf{var}(\varphi)$) of all the minimal satisfying assignments of $\varphi$.
We are next going to define the model Decision dnnf, which throughout this paper is referred to as $\wedge_{d}$-fbdd (the reasons will become clear later). The model is based on a directed acyclic graph (dag) with a single source with vertices and some edges possibly being labelled. We will often use the following notation. Let $D$ be a dag. Then $V(D)$ is the set of vertices (nodes) of $D$. Further on, let $u\in V(D)$. Then we denote by $D_{u}$ the subgraph of $D$ induced by $u$ and all the vertices reachable from $u$ in $D$. The labels on vertices and edges of $D_{u}$ are the same as the labels of the respective vertices and edges of $D$.
We define $\wedge_{d}$-fbdd in the following two definitions. The first one will define $\wedge_{d}$-fbdd as a combinatorial object. In the second definition, we will define the semantics of the object that is the function represented by it.
Definition 3.
A $\wedge_{d}$-fbdd is a dag $B$ with a single source and two sinks. One sink is labelled by $True$, the other by $False$ (on the illustrations we denote the sinks by $T$ and $F$ respectively). Each non-sink node has two out-neighbours (children). A non-sink node may be either a decomposable conjunction node, labelled with $\wedge_{d}$ (we often refer to such nodes as $\wedge_{d}$-nodes) or a decision node labelled with a variable name. If $u$ is a decision node then one outgoing edge is labelled with 0, the other with $1$.
Let $u$ be a node of $B$. Then $\mathsf{var}(B_{u})$ is the set of variables labelling the decision nodes of $_{u}. BB$ obeys two constraints: decomposability and read-onceness. The decomposability constraint means that if $u$ is a $\wedge_{d}$-node and $u_{0}$ and $u_{1}$ are the children of $u$ then $\mathsf{var}(B_{u_{0}})\cap\mathsf{var}(B_{u_{1}})=\emptyset$. The read-onceness means that two decision nodes labelled by the same variable cannot be connected by a directed path.
We define the size of $B$ as the number of nodes and refer to it as $|B|$.
We observe that if $B$ is a $\wedge_{d}$-fbdd and $u\in V(B)$ then $B_{u}$ is also a $\wedge_{d}$-fbdd. This allows us to introduce recursive definition from the sinks up to the source of $B$.
Definition 4.
Let $B$ be a $\wedge_{d}$-fbdd. We define the set $\mathcal{A}(B)$ of assignments accepted by $B$ in the following recursive way. If $B$ consists of a sink node only then $\mathcal{A}(B)=\{\emptyset\}$ if the sink is labelled by $True$ and $\emptyset$ otherwise.
Otherwise, let $u$ be the source of $B$, let $u_{0}$ and $u_{1}$ be the children of $u$, and let $\mathcal{A}_{i}=\mathcal{A}(B_{u_{i}})$ for each $i\in\{1,2\}$.
Assume first that $u$ is a decision node and let $x$ be the variable labelling $u$. We assume w.l.o.g that $(u,u_{0})$ is labelled by 0 and $(u,u_{1})$ is labelled by $1$. Then $\mathcal{A}(B)=\mathcal{A}_{0}\times\{(x,0)\}\cup\mathcal{A}_{1}\times\{(x,1)\}$. If $u$ is a conjunction node then $\mathcal{A}(B)=\mathcal{A}_{0}\times\mathcal{A}_{1}$.
The function $f(B)$ represented by $B$ is the function over $\mathsf{var}(B)$ whose satisfying assignments are all $\bf b$} such that there is ${\bf a}\in\mathcal{A}(B)$ such that ${\bf a}\subseteq{\bf b}$. We denote $\mathcal{S}(f(B))$ by $\mathcal{S}(B)$ and refer to this set as the set of satisfying assignments of $B$.
Example 1.
Figure 1 illustrates $\wedge_{d}$-fbdd $B$ representing function $(x\rightarrow(y_{1}\wedge y_{2}))\vee(\neg x\rightarrow(y_{1}\wedge y_{3}))$. In particular, $\mathcal{A}(B)=\{\{(x,1),(y_{1},1),(y_{2},1)\},\{(x,0),(y_{1},1),(y_{3},1)\}\}$. Clearly, $\mathcal{S}(B)$ are all possible supersets of the assignments of $\mathcal{A}(B)$ assigning all of $x,y_{1},y_{2},y_{3}$.
[…]
Figure 1: An example of a $\wedge_{d}$-obdd
In read-once models, directed paths are naturally associated with assignments. A formal definition is provided below.
Definition 5.
Let $P$ be a directed path of a $\wedge_{d}$-fbdd. We denote by $\mathsf{var}(P)$ the set of variables labelling the nodes of $P$ but the last one. We denote by ${\bf a}(P)$ the assignment over $\mathsf{var}(P)$ where for each $x\in\mathsf{var}(P)$, ${\bf a}(x)$ is determined as follows. Let $u$ be the node of $P$ labelled by $x$. Then ${\bf a}(x)$ is the label on the outgoing edge of $u$ in $P$. (Note that even if the last node of $P$ is labelled by a variable, the variable is not included into $\mathsf{var}(P)$ exactly because the outgoing edge of the last node is not in $P$ and hence the assignment of the variable cannot be defined by the above procedure.)
We next define subclasses of $\wedge_{d}$-fbdds. Let $B$ be a $\wedge_{d}$-fbdd that does not have $\wedge_{d}$ nodes. Then $B$ is a fbdd. Next, let $B$ be a $\wedge_{d}$-fbdd such that there is a linear order $\pi$ over a superset of $\mathsf{var}(B)$ so that the following holds. Let $P$ be a directed path of $B$ and suppose that $P$ has two decision nodes $u_{1}$ and $u_{2}$ labelled by variables $x_{1}$ and $x_{2}$ and assume that $u_{1}$ occurs before $u_{2}$ in $P$. Then $x_{1}<_{\pi}x_{2}$ (that is $x_{1}$ is smaller than $x_{2}$ according to $\pi$). Then $B$ is called an $\wedge_{d}$-obdd (respecting $\pi$ if not clear from the context). Note that the $\wedge_{d}$-fbdd $B$ considered in Example 1 is in fact an $\wedge_{d}$-obdd respecting an arbitrary order over $\{x,y_{1},y_{2},y_{3}\}$ provided that $x$ is ordered first. Finally, an $\wedge_{d}$-obdd without $\wedge_{d}$ nodes is just an obdd.
We remark that allowing $\pi$ to be defined over a superset of $\mathsf{var}(B)$ will be convenient for our reasoning when we consider subgraphs of a $\wedge_{d}$-obdd.
We will need one more restricted class of $\wedge_{d}$-fbdd, a structured one. We will define it in Section 5 where it will actually be used.
Throughout the paper, when we refer to the size of a smallest model representing a function $f$, we will use the notation $modelname(f)$. For instance $\wedge_{d}$-obdd(f)(f), fbdd(f)(f), obdd(f)(f) respectively refer to the smallest sizes of $\wedge_{d}$-obdd, fbdd, and obdd representing $f$.
2.3 Graphs, width parameters, and correspondence with CNF
We use the standard terminology related to graph as specified in [9]. In particular for a graph $G$, $U\subseteq V(G)$ and $E^{\prime}\subseteq E(G)$, $G[U]$ and $G[E^{\prime}]$ respectively denote the subgraphs of $G$ induced by $U$ and $E^{\prime}$ Also, let us recall that a matching $M$ of a graph $G$ is a subset of its edges such that for any distinct $e_{1},e_{2}\in M$, $e_{1}\cap e_{2}=\emptyset$. The treewidth of $G$, denoted by $tw(G)$ is arguably the best known graph parameter measuring closeness to a tree. To define $tw(G)$, we need the notion of a tree decomposition of $G$ which is a pair $(T,{\bf B})$ where $T$ is a tree and ${\bf B}:V(T)\rightarrow 2^{V(G)}$. The set ${\bf B}(t)$ for $t\in V(T)$ is called the bag of $T$. The bags obey the rules of containement meaning that for each $e\in V(G)$ there is $t\in V(T)$ such that $e\subseteq{\bf B}(t)$ and connectivity. For the latter, let $T[v]$ be the subgraph of $T$ induced by $t$ such that $v\in{\bf B}(t)$. Then $T[v]$ is connected. For $t\in V(T)$, the width of ${\bf B}(t)$ is $|{\bf B}(t)|-1$ The width of $(T,{\bf B})$ is the largest width of a bag. Finally, $tw(G)$ is the smallest width of a tree decomposition. The pathwidth of $G$, denoted by $pw(G)$, is the smallest width of a tree decomposition $(T,{\bf B})$ of $G$ such that $T$ is a path.
A central notion used throughout this paper is a matching crossing a sequence of vertices. This notion is formally defined below.
Definition 6.
Let $G$ be a graph. Let $M$ be a matching of $G$ and let $U,V$ be a partition of $V(M)$ such that each edge of $M$ has one end in $U$, the other in $V$. We then say that $M$ is a matching between $U$ and $V$. Let $\pi$ be a linear order of $V(G)$. A matching $M$ of $G$ crosses $\pi$ if there is a prefix $\pi_{0}$ of $\pi$ with the corresponding suffix $\pi_{1}$ so that $M$ s a matching between $\pi_{0}\cap V(M),\pi_{1}\cap V(M)$ We call $\pi_{0}$ a witnessing prefix for $M$.
Remark 2.
In Definition 6, we overloaded the notation by using a prefix of $\pi$ both as a linear order and as a set. The correct use will always be clear from the context.
Example 2.
Let $G=P_{8}$ (a path of $8$ vertices) with $V(G)=\{v_{1},\dots,v_{8}\}$ appearing along the path in the order listed. Consider an order $\pi=(v_{1},\dots,v_{8})$. Then the only matchings crossing $\pi$ are singleton ones. On the other hand, consider the order $\pi^{*}=(v_{1},v_{3},v_{5},v_{7},v_{2},v_{4},v_{6},v_{8})$. Then the matching $M=\{\{v_{1},v_{2}\},\{v_{3},v_{4}\},\{v_{5},v_{6}\},\{v_{7},v_{8}\}\}$ crosses $\pi^{*}$ as witnessed by $(v_{1},v_{3},v_{5},v_{7})$.
Taking to the maximum the size of a matching crossing the given linear order leads us to the following width parameters.
Definition 7.
Let $G$ be a graph and $\pi$ be linear order of $V(G)$. The linear maximum matching width lmm-width of $\pi$ is the largest size of a matching of $G$ crossing $\pi$. The lmm-width of $G$, denoted by $lmmw(G)$ is the smallest linear matching width over all linear orders of $\pi$.
The linear special induced matching width (lsim-width) of $\pi$ is the largest size of an induced matching of $G$ crossing $\pi$. Accordingly, lsim-width of $G$, denoted by $lsimw(G)$, is the smallest lsim-width of a linear order of $G$.
The parameters lmm-width and lsim-width are linear versions of well known graph parameters maximum matching width [24] and special induced matching width [11].
Lemma 1.
There is a constant $c$ such that for each sufficiently large $k$ there is an infinite class ${\bf G}_{k}$ of graphs such that for each $G\in{\bf G}_{k}$, $tw(G)\leq k$ and $lsimw(G)\geq c\log(|V(G)|)\cdot k$.
Proof.
It is known [19] that the statement as in the lemma holds if $lsimw(G)$ is replaced by $pw(G)$ and, moreover, the considered class of graphs has max-degree $5$. Next, it is known [20] that the $pw(G)$ and $lmmw(G)$ are linearly related. This means that the lemma is true if $lsimw(G)$ is replaced by $lmmw(G)$. Finally, for any matching $M$, there is an induced subset $M^{\prime}$ of size at least $|M|/(2d+1)$ where $d$ is the max-degree of $G$. Indeed, greedily choose edges $e\in M$ into $M^{\prime}$ by removing the edges of $M$ connecting $N(e)$ (at most $2d$ edges are removed per chosen edge). Since $d=5$ for the considered class, $lsimw(G)\geq lmmw(G)/11$. ∎
For a cnf $\varphi$, we use two graphs, both regard $\varphi$ as a hypergraph over $\mathsf{var}(\varphi)$ ignoring the literals with clauses corresponding to hyperedges. The primal graph of $\varphi$ has $\mathsf{var}(\varphi)$ as the set of vertices and the two vertices are adjacent if they occur in the same clause. The incidence graph of $\varphi$ has $\mathsf{var}(\varphi)$ and the clauses as its vertices. A variable $x$ is adjacent to a clause $C$ if and only if $x$ occurs in $C$. We denote by $ptw(\varphi)$ and $itw(\varphi)$ the trewidth of the primal and incidence graphs of $\varphi$ respectively.
Proposition 2.
$itw(\varphi)\leq ptw(\varphi)$ [23]
3 $\wedge_{d}$-OBDD: inefficiency in terms of incidence treewidth and separations
3.1 An approach to proving $\wedge_{d}$-OBDD lower bounds
We start from overviewing an approach for proving lower bounds for obdds and then generalize it to $\wedge_{d}$-obdds. This is the standard ‘lower bound by the number of subfunctions’ approach as described in [25] but rephrased for the purpose of the further generalization.
Let $B$ be an obdd respecting an order $\pi$. Let $\bf g$} be an assignment over a prefix of $\pi$. Then we can identify the maximal path $P=P({\bf g})$ starting from the source and such that ${\bf a}(P)\subseteq{\bf g}$. Indeed, existence of such a path and its uniqueness is easy to verify algorithmically. Let the current vertex $u$ of $B$ be the source and the current variable $x$ be the label of $u$. Furthermore, let the current path $P$ be consisting of $u$ only. If $x\notin\mathsf{var}({\bf g})$ then stop. Otherwise, choose the out-neighbour $v$ of $u$ such that $(u,v)$ is labelled with ${\bf g}(x)$. Let the current vertex be $v$, the current variable be the label of $v$ and the current path be $P$ plus $(u,v)$. Repeat the process until the stopping condition reached. We let $u({\bf g})$ be the final vertex of $P({\bf g})$.
Next, we define $\mathcal{F}$, a set of fooling assignments, for each $\bf g$$\in\mathcal{F}$, $\mathsf{var}({\bf g})$ is a prefix of $\pi$. The idea is to prove that for every two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$, $u({\bf g}_{1})\neq u({\bf g}_{2})$. This implies that $|B|\geq|\mathcal{F}|$. In particular, if $|\mathcal{F}|$ is large in terms of the number of variables, then the respective large lower bound follows.
For $\wedge_{d}$-obdds, an assignment to a prefix of the respected order is associated with a rooted tree rather than with a path. In the rest of the subsection, we upgrade the above methodology for this case.
Definition 8.
Let $\bf g$} be an assignment over a subset of $\mathsf{var}(B)$. The alignment of $B$ by $\bf g$} denoted by $B[{\bf g}]$ is obtained from $B$ by the following two operations.
  1. 1.
    For each decision node $u$ labelled by $x\in\mathsf{var}({\bf g})$, remove the outgoing edge of $u$ labelled with $1-{\bf g}(x)$. Let $B_{0}[{\bf g}]$ be the resulting subgraph of $B$.
  2. 2.
    Remove all the nodes of $B_{0}[{\bf g}]$ that are not reachable from the source of $B[{\bf g}]$.
Example 3.
Figure 2 demonstrates a $\wedge_{d}$-obdd $B$ for $(x_{1}\vee x_{2})\wedge(x_{2}\vee x_{3})\wedge(x_{4}\vee x_{5})\wedge(x_{5}\vee x_{6})$ (on the left) respecting the order $(x_{2},x_{1},x_{3},x_{5},x_{4},x_{6})$ and $B[{\bf g}]$ (on the right) where ${\bf g}=\{(x_{2},1)\}$.
The $True$ and $False$ sinks are respectively labelled with $T$ and $F$. Note that, for the sake of a better visualization, we deviated from the definition requiring a $\wedge_{d}$-obdd to have a single $True$ and a single $False$ sinks. To make the models on the picture consistent with the definition, all $True$ sinks of each model need to be contracted together and the same contraction needs to be done for all the $False$ sinks.
[…]
Figure 2: A $\wedge_{d}$-obdd and its alignment
We note that $B[{\bf g}]$ is not necessarily a $\wedge_{d}$-obdd since some decision nodes may have only one outgoing edge. This situation is addressed in the following definition.
Definition 9.
A decision node of $B[{\bf g}]$ that has only one outgoing neighbour is called an incomplete decision node. A maximal path with all intermediate decision nodes being incomplete is called an incomplete path. We note that a $\wedge_{d}$ node cannot be a final node of an incomplete path due to its maximality, hence an incomplete path may end only with a sink of $B$ or with a decision node of $B$. We denote by $L({\bf g})$ the set of all decision nodes that are final nodes of the incomplete paths of $B[{\bf g}]$.
Back to Example 3, $L({\bf g})$ is a singleton consisting of the only node labelled by $x_{2}$. We note that $L({\bf g})$ are exactly minimal complete decision nodes of $B[{\bf g}]$ in the sense that each decision node that is a proper descendant of a node of $L({\bf g})$ must be complete.
Lemma 2.
Let $\pi$ be the order respected by $B$ and let $\bf g$} be an assignment over a prefix of $\pi$. Then the following statements hold regarding $B[{\bf g}]$.
  1. 1.
    Let $u$ be a complete decision node and let $v$ be a decision node being a successor of $u$ (that is, $B[{\bf g}]$ has a path from $u$ to $v$). Then $v$ is also a complete decision node. In other words, $B[{\bf g}]_{u}$ is a $\wedge_{d}$-obdd.
  2. 2.
    Let $u_{1},u_{2}\in L({\bf g})$. Then $\mathsf{var}(B[{\bf g}]_{u_{1}})\cap\mathsf{var}(B[{\bf g}]_{u_{2}})=\emptyset$.
  3. 3.
    Let $T({\bf g})$ be the union of all incomplete paths ending with nodes of $L({\bf g})$. Then $T({\bf g})$ is a rooted tree. Moreover, all the intermediate nodes of $T({\bf g})$ with out-degree $2$ are conjunction nodes. In particular, this implies that if $B$ is an obdd then $T({\bf g})$ is just a path.
Proof.
Let $x$ and $y$ be the variables labelling $u$ and $v$ respectively. By assumption $x<_{\pi}y$ and $x\notin\mathsf{var}({\bf g})$, the latter constitutes a prefix of $\pi$. It follows that $y\notin\mathsf{var}({\bf g})$ and hence $v$ is complete simply by construction.
For the second statement, note that, by assumption, none of $u_{1}$ and $u_{2}$ is a successor of the other. Therefore, in $B[{\bf g}]$ there is a node $v$ with two children $v_{1}$ and $v_{2}$ such that $u_{i}$ is a (possibly not proper) successor of $v_{i}$ for each $i\in\{1,2\}$. If $v$ is a decision node then it is complete, in contradiction to the minimality of $u_{1}$ and $u_{2}$. We conclude that $v$ is a conjunction node that hence $\mathsf{var}(B[{\bf g}]_{v_{1}})\cap\mathsf{var}(B[{\bf g}]_{v_{2}})=\emptyset$. As $\mathsf{var}(B[{\bf g}]_{u_{1}})\cap\mathsf{var}(B[{\bf g}]_{u_{2}})\subseteq\mathsf{var}(B[{\bf g}]_{v_{1}})\cap\mathsf{var}(B[{\bf g}]_{v_{2}})$, the second statement follows.
For the third statement, it is sufficient to prove that any two distinct incomplete paths $P_{1}$ and $P_{2}$ do not have nodes in common besides their comon prefix and that the last node of the prefix is $\wedge_{d}$ one. So, let $P$ be the longest common prefix of $P_{1}$ and $P_{2}$ and let $u$ be the last vertex of $P$. Arguing as in the previous paragraph, we conclude that $u$ is a $\wedge_{d}$ node. Let $u_{1}$ and $u_{2}$ be the children of $u$. Then, by the maximality of $P$, we can assume w.l.o.g. that $u_{1}\in V(P_{1})$ and $u_{2}\in V(P_{2})$.
Suppose that $P_{1}$ and $P_{2}$ have a common vertex $v$ that is a proper successor of $u$. As $v$ is a predecessor of a vertex of $L({\bf g})$, $\mathsf{var}(B_{v})\neq\emptyset$. As $\mathsf{var}(B_{v})\subseteq\mathsf{var}(B_{u_{1}})\cap\mathsf{var}(B_{u_{2}})$, we get a contradiction to the decomposability of $u$. ∎
Lemma 2 is illustrated in Figure 3. Suppose that the picture illustrates $B[{\bf g}]$ for some $\wedge_{d}$-obdd $B$ and an assignment $\bf g$} over a prefix of the order respected by $B$. Then the rooted tree within the outer dashed rectangle is $T({\bf g})$, the set of nodes in the inner dashed is $L({\bf g})$. We note that the nodes of $T({\bf g})$ with two children are the $\wedge_{d}$ ones. Each $u\in L({\bf g})$ is contained in an ellipse and the ellipse denotes $B_{u}$, this is demonstrated for the leftmost ellipse. All the $\wedge_{d}$-obdd are assumed to have the same $True$ and $False$ sink, we do not demonstrate this on the picture for the sake of clarity.
$uB_{u}$$L({\bf g})$$T({\bf g})$$\wedge_{d}$$\wedge_{d}$
Figure 3: An intuitive illustration of $T({\bf g})$ and $L({\bf g})$
The notion of $L({\bf g})$ allows a neat characterization of $\mathcal{S}(B)|_{\bf g}$ as specified in the next lemma.
Lemma 3.
Let the notation and premises be as in Lemma 2. In addition, assume that $\mathcal{S}(B)|_{\bf g}\neq\emptyset$. Let $X=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\bigcup_{u\in L({\bf g})}\mathsf{var}(B_{u})$. (In other words, $X$ is the set of variables not used by the model under the restriction posed by $\bf g$}.) If $L({\bf g})\neq\emptyset$ then $\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}\times\prod_{u\in L({\bf g})}\mathcal{S}(B_{u})$. Otherwise, $\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}$.
We believe that Figure 3 makes the statement of Lemma 3 intuitively clear. We formally prove the lemma by induction on $|\mathsf{var}(B)|$. The proof is rather tedious due to the need to inductively account $L({\bf g})$ and $X$ as composed from the corresponding objects for subgraphs of $B$ and the need to prove that the induction assumption holds for these subgraphs. We therefore postpone the full proof to the Appendix and proceed to the use of the lemma for the proposed method of proving lower bounds for $\wedge_{d}$-obdds.
The central part of the method is the notion of a fooling set of a Boolean function.
Definition 10.
Let $f$ be a Boolean function and let $\pi$ be a permutation of $\mathsf{var}(f)$. Let $\mathcal{F}$ be a set of assignments over a prefix $\pi_{0}$ of $\pi$. We say that $\mathcal{F}$ is a ($\wedge_{d}$-obdd-)fooling set for $f$ under $\pi$ if for each $\bf g$$\in\mathcal{F}$ there is an unbreakable set $UB({\bf g})\subseteq\pi\setminus\pi_{0}$ of size at least two having the following properties.
  1. 1.
    Unbreakability: For each $\bf g$$\in\mathcal{F}$, $\mathcal{S}(f)$ does not break $UB({\bf g})$.
  2. 2.
    Distinct projections: For every two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$, $Proj(\mathcal{S}(f)|_{{\bf g}_{1}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))\neq Proj(\mathcal{S}(f)|_{{\bf g}_{2}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))$.
In the next lemma, we demonstrate that if $\mathcal{F}$ is a fooling set of a Boolean function $f$ under a permutation $\pi$ then any $\wedge_{d}$-obdd representing $f$ and respecting $\pi$ must have size at least $|\mathcal{F}|$. The proof consists of two stages. At the first stage, we consider an arbitrary $\bf g$$\in\mathcal{F}$. Combining the unbreakability property with Lemma 3, we identify precisely one $u\in L({\bf g})$ such that $UB({\bf g})\subseteq\mathsf{var}(B_{u})$. We refer to this $u$ as $u({\bf g})$.
In the second part of the proof, we demonstrate that the mapping from $\bf g$} to $u({\bf g})$ is injective thus implying the desired lower bound. For that purpose, we use the distinct projections property. In particular, for the sake of contradiction, we assume that there are two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$ such that $u({\bf g}_{1})$ and $u({\bf g}_{2})$ is the same node and denote this node by $u$. Another application of Lemma 3 demonstrates that, for each $i\in[2]$, the projection of $\mathcal{S}(f)|_{{\bf g}_{i}}$ to $UB({\bf g}_{1})\cup UB({\bf g}_{2})$ is $Proj(\mathcal{S}(B_{u}),UB({\bf g}_{1})\cup UB({\bf g}_{2}))$. That is, both projections are the same in contradiction to the distinct projections property.
Now, let us formally state and prove the lemma.
Lemma 4.
Let $f$ be a Boolean function and let $\mathcal{F}$ be a fooling set for $f$ under a permutation $\pi$ of $\mathsf{var}(f)$. Let $B$ be a $\wedge_{d}$-obdd representing $f$ and respecting $\pi$. Then $|B|\geq|\mathcal{F}|$.
Proof.
First, we note that $\mathcal{S}(B)=\mathcal{S}(f)$. Let $\bf g$$\in\mathcal{F}$. By Proposition 1 combined with Lemma 3, either $UB({\bf g})\subseteq X$ or there is $u\in L({\bf g})$ with $UB({\bf g})\subseteq\mathsf{var}(B_{u})$. The former is ruled out simply by contradiction to unbreakability ($X$ can be arbitrarily broken whereas $UB({\bf g})$ is of size at least $2$ so $UB({\bf g})\subseteq X$ would mean that breaking across $UB({\bf g})$ is possible). We denote the $u$ such that $UB({\bf g})\subseteq\mathsf{var}(B_{u})$ by $u({\bf g})$. We prove that the mapping from $\mathcal{F}$ to $V(B)$ such that $\bf g$$\in\mathcal{F}$ is mapped to $u({\bf g})$ is injective and this implies the desired lower bound.
Indeed, assume the opposite. Then there are distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$ such that $u=u({\bf g}_{1})=u({\bf g}_{2})$. By Lemma 3, for each $i\in[2]$, $Proj(\mathcal{S}(B)|_{{\bf g}_{i}},\mathsf{var}(B_{u}))=\mathcal{S}(B_{u})$. This in particular, implies that for each $i\in[2]$, $Proj(\mathcal{S}(B)|_{{\bf g}_{i}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))=Proj(\mathcal{S}(B_{u}),UB({\bf g}_{1})\cup UB({\bf g}_{2}))$ in contradiction to the distinct projections assumption. ∎
Thus, in order to prove that a $\wedge_{d}$-obdd representing a function $f$ and respecting a specific order $\pi$ is large, it is enough to demonstrate existence of a large fooling set for $f$ under $\pi$. If $\pi$ is chosen arbitrary, this also implies a lower bound for an arbitrary $\wedge_{d}$-obdd representing $f$. The advantage of this approach is that it is completely ‘function-based’ and does not require direct reasoning about the model. Of course, proving existence of a large $\wedge_{d}$-fooling set may also be challenging. However, the difficulty is comparable with the difficulty of establishing a large fooling set for obdds. Indeed, in the latter case, we effectively need to establish the distinct projection property whereas, in the former case, we also need to establish existence of unbreakable sets. We believe that this extra layer of difficulty is reasonable since $\wedge_{d}$-obdd is a much more powerful model than obdd (as demonstrated in the next subsection). Thus, the proposed approach is a useful tool for establishing lower bounds. We demonstrate its work in Section 3.3 where we use the tool for proving Theorem 10.
3.2 $\wedge_{d}$-OBDD lower bounds and separations
In [5] we established an xp lower bound for $\wedge_{d}$-obdds representing cnfs of bounded incidence treewidth. We prove this result here using the approach presented in Section 3.1.
Theorem 5.
[…]
Thus a natural restriction on $\wedge_{d}$-fbdd separates primal [14] and incidence treewidth. This makes $\wedge_{d}$-obdd an interesting object to study in greater detail. We establish the relationship of $\wedge_{d}$-obdd with two closely related formalisms obdd and fbdd by proving exponential separations as specified below (for both statements, the lower bounds are proved using the approach as in Section 3.1). In the next section, we study the Apply operation for $\wedge_{d}$-obdds.
Theorem 6.
There is an infinite class of cnfs polynomially representable as fbdds but requiring exponential-size representation as $\wedge_{d}$-obdds.
Theorem 7.
There is an infinite class of cnfs polynomially representable as $\wedge_{d}$-obdds but requiring exponential-size representation as obdds.
Remark 3.
As $\wedge_{d}$-obdd is not a special case of fbdd it is natural to ask whether an opposite of Theorem 6 holds. It is known from [3] that $\wedge_{d}$-fbdd can be quasipolynomially simulated by fbdd. As $\wedge_{d}$-obdd is a special case of $\wedge_{d}$-fbdd, the simulation clearly applies to $\wedge_{d}$-obdd.
For the corresponding quasipolynomial separation, take a class of cnfs of ptwptw at most $k$ that require fbdd of size $n^{\Omega(k)}$ (see [19] for such a class of cnfs). Let $k=\log n$ where $n$ is the number of variables. Then the resulting class of cnfs will require $n^{\Omega(\log(n)}$ size fbdds for their representation but can be represented by a polynomial size $\wedge_{d}$-obdd according to [14].
Remark 4.
Theorem 7 looks somewhat surprising because equipping fbdds with the $\wedge_{d}$-nodes leads only to quasipolynomial gains in size [3]
In the two definitions below, we introduce classes of cnfs needed for the stated lower bounds and separations.
Definition 11.
Let $G$ be a graph without isolated vertices. Then $\varphi(G)$ is a cnf whose set of variables is $V(G)$ and the set of clauses is $\{(u\vee v)|\{u,v\}\in E(G)\}$.
Let $V=V(G)$ and let $V[1]=\{v[1]|v\in V\}$ and $V[2]=\{v[2]|v\in V\}$ be two disjoint copies of $V$. Let $G^{*}$ be the graph with $V(G^{*})=V[1]\cup V[2]$ and $E(G^{*})=\bigcup_{\{u,v\}\in E(G)}\{\{u[1],v[2]\},\{v[1],u[2]\}\}$.
The cnf $\psi(G)$ is obtained from $\varphi(G^{*})$ by introducing two additional clauses $\neg V[1]=\{\neg v[1]|v\in V\}$ and $\neg V[2]=\{\neg v[2]|v\in V\}$.
Extending the notation $V[i]$, for $U\subseteq V(G)$, and $i\in\{1,2\}$, we let $U[i]=\{u[i]|u\in U\}$.
Example 4.
Let $G$ be a graph with $V(G)=\{u_{1},\dots,u_{5}\}$ and
$E(G)=\{\{u_{1},u_{2}\},\{u_{2},u_{3}\},\{u_{3},u_{4}\},\{u_{4},u_{5}\},\{u_{5},u_{1}\},\{u_{1},u_{3}\}\}$. In other words, $G$ is a cycle of $5$ vertices with an extra chord. Then $V(G^{*})=\{u_{1}[1],\dots,u_{5}[1],u_{1}[2],\dots,u_{5}[2]\}$ and $E(G^{*})=\{\{u_{1}[1],u_{2}[2]\},\{u_{2}[1],u_{3}[2]\},\{u_{3}[1],u_{4}[2]\},\{u_{4}[1],u_{5}[2]\},\{u_{5}[1],u_{1}[2]\},\{u_{1}[1],u_{3}[2]\},\\ \{u_{1}[2],u_{2}[1]\},\{u_{2}[2],u_{3}[1]\},\{u_{3}[2],u_{4}[1]\},\{u_{4}[2],u_{5}[1]\},\{u_{5}[2],u_{1}[1]\},\{u_{1}[2],u_{3}[1]\}\}$. $\varphi(G)$ and $\varphi(G^{*})$ simply turn the edges $\{u,v\}$ of the respective graphs into clauses $(u\vee v)$. $\psi(G)$ is obtained from $\varphi(G^{*})$ by adding clauses $(\neg u_{1}[1]\vee\dots\vee\neg u_{5}[1])$ and $(\neg u_{1}[2]\vee\dots\vee\neg u_{5}[2])$
We introduce two more classes of formulas that are obtained from $\varphi(G)$ and $\psi(G)$ by introducing junction variables.
Definition 12.
Let $G$ be a graph without isolated vertices and let $E_{1},E_{2}$ be a partition of $E(G)$ so that $V(G[E_{1}])=V(G[E_{2}])=V(G)$.
We denote by $\varphi(G,E_{1},E_{2})$ the function $(\textsc{jn}\rightarrow\varphi(G[E_{1}]))\wedge(\neg\textsc{jn}\rightarrow\varphi(G[E_{2}])$ where jn is a junction variable that does not belong to $V(G)$. It is not hard to see that, $\varphi(G,E_{1},E_{2})$ can be expressed as a cnf $\bigwedge_{\{u,v\}\in E_{1}}(\neg\textsc{jn}\vee u\vee v)\wedge\bigwedge_{\{u,v\}\in E_{2}}(\textsc{jn}\vee u\vee v)$.
Similarly, we denote by $\psi(G,E_{1},E_{2})$ the function $(\textsc{jn}\rightarrow\psi(G[E_{1}]))\wedge(\neg\textsc{jn}\rightarrow\psi(G[E_{2}])$
Remark 5.
We will use cnfs with junction variables for Theorems 6 and 7. This approach is used in [25] for separation of obdds and fbdds. The rough idea is that a ’hard’ function is ’split’ into two ’easy’ ones using the junction variable. The fbdd, due to its flexibility, can adjust and represent each ’easy’ function with the suitable variable ordering. On the other hand, obdd is rigid in the sense that it needs to use the same ordering for both ’easy’ functions and the ordering leads to an exponential size representation for one of them.
In order to proceed, we need to delve deeper into the structure of matchings of the graph $G^{*}$. This is done in the next definition and the two statements that follow.
Definition 13.
Let $M$ be a matching of $G^{*}$ between $V^{\prime}$ and $V^{\prime\prime}$. We say that $V^{\prime},V^{\prime\prime}$ is the neat partition of $V(M)$ if there are $U,W\subseteq V(G)$ such that $V^{\prime}=U[1]$ and $V^{\prime\prime}=W[2]$.
Let $\pi^{*}$ be a linear order of $V(G^{*})$. We say that $M$ neatly crosses $\pi^{*}$ if there a witnessing prefix $\pi^{*}_{0}$ of $\pi^{*}$ such that $\pi^{*}_{0}\cap V(M)$ and $(\pi^{*}\setminus\pi^{*}_{0})\cap V(M)$ form a neat partition of $M$.
Example 5.
Consider a graph $G$ with four vertices $u^{\prime},u^{\prime\prime},v^{\prime},v^{\prime\prime}$ and two edges: $\{u^{\prime},u^{\prime\prime}\}$ and $\{v^{\prime},v^{\prime\prime}\}$. Let $M=\{\{u^{\prime}[1],u^{\prime\prime}[2]\},\{v^{\prime\prime}[1],v^{\prime}[2]\}\}$. Clearly, $M$ is a matching of $G^{*}$. The neat partition for $M$ is $\{\{u^{\prime}[1],v^{\prime\prime}[1]\},\{u^{\prime\prime}[2],v^{\prime}[2]\}\}$, the corresponding partition of $V(G)$ is $\{\{u^{\prime},v^{\prime\prime}\},\{u^{\prime\prime},v^{\prime}\}\}$.
Consider the linear order $\pi^{*}=(v^{\prime\prime}[2],u^{\prime}[1],u^{\prime}[2],u^{\prime\prime}[1],v^{\prime\prime}[1],v^{\prime}[1],u^{\prime\prime}[2],v^{\prime}[2])$. The order neatly crosses $M$ as witnessed by the prefix $\pi^{*}_{0}=(v^{\prime\prime}[2],u^{\prime}[1],u^{\prime}[2],u^{\prime\prime}[1],v^{\prime\prime}[1])$. On the other hand, if in $\pi^{*}$ we swap $v^{\prime\prime}[1]$ and $u^{\prime\prime}[2]$ then we obtain a linear order of $V(G^{*})$ that does not cross $M$ neatly.
In the next statement, we demonstrate that if each linear order of $G$ is crossed by a large induced matching then every linear order of $G^{*}$ is neatly crossed by a large induced matching. The statement is formulated in terms of parameter lsim-width defined earlier.
Lemma 8.
Every linear order of $G^{*}$ is neatly crossed by an induced matching of size at least $lsimw(G)/2$
Proof.
Let $\pi^{*}$ be a linear order of $V(G^{*})$. Let $\pi$ be a linear order of $V(G)$ obtained from $\pi^{*}$ as follows. For each $v\in V(G)$, let $ind(v)=1$ if $v[1]$ appears in $\pi^{*}$ before $v[2]$ and $2$ otherwise. In $\pi$, $u$ is ordered before $v$ if and only if $u[ind(u)]$ is ordered before $v[ind(v)]$.
Let $\pi_{0}$ be a prefix of $\pi$ witnessing a matching $M$ of size at least $lsimw(G)$ crossing $\pi$. Let $U=V(M)\cap\pi_{0}$. Then $U$ is can be weakly partitioned into $U_{1}$ and $U_{2}$ such that $U_{i}$ is the set of $u\in U$ such that $ind(u)=i$. By the pigeonhole principle, the size of either $U_{1}$ or $U_{2}$ is at least $|U|/2$. We assume the former w.l.o.g. Let $M_{1}$ be the subset of $M$ consisting of all $e\in M$ such that $U_{1}\cap e\neq\emptyset$. Let $W_{1}$ be $V(M_{1})\setminus U_{1}$. Clearly, $M_{1}$ is a matching between $U_{1}$ and $W_{1}$. Let $M^{*}_{1}=\{\{u[1],w[2]\}|\{u,v\}\in M_{1}$, $u\in U_{1},w\in W_{1}\}$. It is not hard to see that $M^{*}_{1}$ is an induced matching of $G^{*}$ between $U_{1}[1]$ and $W_{1}[2]$.
Let $\pi^{*}_{0}$ be the largest prefix of $\pi^{*}$ finishing with a vertex of $U_{1}[1]$ and let $\pi^{*}_{1}$ be the matching suffix. We claim that $\pi^{*}_{0}$ is a witnessing prefix for $\pi^{*}$ being neatly crossed by $M^{*}_{1}$ thus establishing the lemma. It is immediate from the definition that $U_{1}[1]\subseteq\pi^{*}_{0}$. It thus remains to observe that $W_{1}[2]\subseteq\pi^{*}_{1}$.
Let $u[1]$ be the last vertex of $\pi^{*}_{0}$. Let $w[2]\in W_{1}[2]$. Note that by definition of $\pi_{0}$, $u\in\pi_{0}$ while $w$ is not. It follows from $u[ind(u)]<w[ind(w)]$ according to $\pi^{*}$. Now $ind(u)=1$ by definition of $U_{1}$. That is $u[1]$ occurs before $w[ind(w)]$ and surely before w[3ind(w)]w[3-ind(w)] in $\pi^{*}$. As $u[1]$ is the last element of $\pi^{*}_{0}$, both occurrences of $w$ are outside of $\pi^{*}_{0}$. ∎
Lemma 9.
Let $G$ be a graph without isolated vertices. Let $E_{1},E_{2}$ be a partition of $E(G)$ so that $V(G[E_{1}])=V(G[E_{2}])=V(G)$. (Note that this means that $V(G^{*})=V(G[E_{1}]^{*})=V(G[E_{2}]^{*})$). Then the following statements hold.
  1. 1.
    Let $\pi$ be a linear order of $V(G)$. Then either $G[E_{1}]$ or $G[E_{2}]$ has an induced matching of size at least $lsimw(G)/2$ crossing $\pi$.
  2. 2.
    Let $\pi^{*}$ be a linear order of $V(G^{*})$, Then either $G[E_{1}]^{*}$ or $G[E_{2}]^{*}$ has an induced matching of size at least $lsimw(G)/4$ neatly crossing $\pi^{*}$.
Proof.
Let $M$ be a matching of $G$ of size at least $lsimw(G)$ crossing $\pi$. Clearly either $M\cap E_{1}$ or $M\cap E_{2}$ is of size at least $|M|/2$. Assume the former w.l.o.g. As $M\cap E_{1}\subseteq M$, clearly, $M\cap E_{1}$ crosses $\pi$. This proves the first statement.
For $i\in\{1,2\}$, Let $E^{*}_{i}=E(G[E_{i}]^{*})$. It is not hard to see that $E^{*}_{1},E^{*}_{2}$ partition $E(G^{*})$ and $G^{*}[E^{*}_{i}]=G[E_{i}]^{*}$ for each $i\in\{1,2\}$. Let $M^{*}$ be an induced matching of size at least $lsimw(G)/2$ neatly crossing $\pi^{*}$ existing by Lemma 8. For each $i\in\{1,2\}$, let $M^{*}_{i}=M^{*}\cap E^{*}_{i}$. Clearly, both $M^{*}_{1}$ and $M^{*}_{2}$ cross $\pi^{*}$ and least one of them is of size at least half the size of $M^{*}$ that is at least $lsimw(G)/4$. ∎
Remark 6.
Note that Lemma 9 does not make any statements regarding $lsimw(G[E_{1}])$ and $lsimw(G[E_{2}])$. Indeed, if $G$ is an $n\times n$ grid, $E_{1}$ is the set of ’horizontal’ edges and $E_{2}$ is the set of ’vertical’ edges then $lsimw((G[E_{1}])=lsimw(G[E_{2}])=1$, while $lsimw(G)=\Omega(n)$. Nevertheless, Lemma 9 states that for each linear order $\pi$ of $V(G)$ must be $\Omega(n)$ either for $G[E_{1}]$ or for $G[E_{2}]$. Continuing on the discussion in Remark 5, the above phenomenon is precisely the reason of rigidity of obdd-based models representing cnfs with junction variables. We will see in the proof of Theorems 6 and 7 below how this leads to exponential lower bounds.
Next, we state two theorems, whose proofs are provided in Section 3.3, that serve as engines for proving lower bounds for Theorems 5, 6, and 7. We will then use Lemmas 8 and Lemma 9 for ‘harnessing’ these engines in order to prove the theorems. For both of the theorems below, $G$ is a graph without isolated vertices.
Theorem 10.
Let $B$ be a $\wedge_{d}$-obdd representing $\psi(G)$. Let $\pi^{*}$ be a linear order of $\mathsf{var}(\psi(G))$ respected by $B$. Let $M$ be an induced matching of $G^{*}$ neatly crossing $\pi^{*}$. Then $|B|\geq 2^{\Omega(|M|)}$.
Theorem 11.
Let $B$ be an obdd representing $\varphi(G)$. Let $\pi$ be a linear order of $V(G)$ respected by $B$. Let $M$ be an induced matching of $G$ crossing $\pi$. Then $|B|\geq 2^{\Omega(|M|)}$.
Proof.
(Theorem 5) For a sufficiently large $k$, let ${\bf G}_{k}$ be a class of graphs $G_{k}$ of treewidth at most $k$ and lsim-width $\Omega(k\log n)$. Existence of such a class of graphs follows from Lemma 1. In order to prove the theorem it is sufficient to demonstrate that $\wedge_{d}$-obdd representing $\psi(G_{k})$ is of size $n^{\Omega(k)}$ and that $itw(\psi(G))=O(k)$.
Let $B$ be a $\wedge_{d}$-obdd representing $\psi(G_{k})$. Let $\pi^{*}$ be a linear order of $V(G^{*}_{k})$ obeyed by $B$. By Lemma 8, there is an induced matching $M$ of $G^{*}_{k}$ of size $\Omega(lsimw(G_{k}))$ neatly crossing $\pi^{*}$. Hence, by Theorem 10, the size of $B$ is $2^{\Omega(lsimw(G))}$ which is $n^{\Omega(k)}$ according to the previous paragraph.
For the proof that $itw(\psi(G))=O(k)$, let $(T,{\bf B})$ be a tree decomposition of $G$ of width at most $k$. For each $t\in V(T)$, replace each $v\in{\bf B}(t)$ with $v[1]$ and $v[2]$. It is not hard to see that we obtain a tree decomposition for the primal graph $\varphi(G^{*})$. As the largest bag size of $(T,{\bf B})$ is at most $k+1$, the largest bag size of the new tree decomposition is at most $2k+2$ and hence $ptw(\varphi(G^{*}))$ is at most $2k+1$. By Proposition 2 $itw(\varphi(G^{*}))$ is at most $2k+1$. Take a tree decomposition witnessing $itw(\varphi(G^{*}))$ and add to each bag two vertices corresponding to the long clauses. As a result we obtain a tree decomposition of the incidence graph $\psi(G)$ of width at most $2k+3$. ∎
We are now turning to the proof of Theorems 6 and 7. For both theorems, the target class of graphs are $n\times n$ grids that we denote by $G_{n}$. A formal definition is provided below.
Definition 14.
Let $n>0$ be an integer. A $n\times n$ grid, denoted by $G_{n}$, is a graph with $V(G_{n})=\{(i,j)|1\leq i\leq n,1\leq j\leq n\}$ and $E(G_{n})=E_{hor}\cup E_{vert}$ where $E_{hor}=\{\{(i,j),(i,j+1)\}|1\leq i\leq n,1\leq j\leq n-1\}$ and $E_{vert}=\{\{(i,j),(i+1,j)\}|1\leq i\leq n-1,1\leq j\leq n\}$. We also refer to $E_{hor}$ and $E_{vert}$ as, respectively, horizontal and vertical edges of $G_{n}$.
As the max-degree of $G_{n}$ is $4$, $lsimw(G_{n})$ and $lmmw(G)$ are linearly related. Next, it is known [20] that $pw(G_{n})$ and $lmmw(G_{n})$ are, in turn, linearly related. Finally, it is well know that $pw(G_{n})\geq tw(G_{n})\geq\Omega(n)$. Therefore, we conclude the following.
Proposition 3.
$lsimw(G_{n})=\Omega(n)$.
We also need an auxuliary lemma whose proof is provided in the Appendix.
Lemma 12.
Let $B$ be a $\wedge_{d}$-obdd respecting an order $\pi$. Let ${\bf g}=\{(x,i)\}$ where $x\in\mathsf{var}(B)$ and $i\in\{0,1\}$. Suppose that all the variables of $f(B)|_{\bf g}$ are essential.  [^1] A variable $x$ of a Boolean function $h$ is essential is there are two assignments ${\bf a}_{1}$ and ${\bf a}_{2}$ that differ only in their assignment of $x$ such that $h({\bf a}_{1})\neq h({\bf a}_{2})$. Let $B^{\prime}$ be obtained from $B$ by the following operations.
  1. 1.
    Removal of all the edges $(u,v)$ such that $u$ is a decision node associated with $x$ and $(u,v)$ is labelled with 1i1-i.
  2. 2.
    Contraction of all the edges $(u,v)$ such $u$ is a decision node associated with $x$ and $(u,v)$ is labelled with $i$ (meaning that $(u,v)$ is removed and $u$ is identified with $v$).
  3. 3.
    Removal of all the nodes not reachable from the source of the resulting graph.
Then $B^{\prime}$ is a $\wedge_{d}$-obdd respecting $\pi$ and $f(B^{\prime})=f(B)|_{\bf g}$.
Proof.
(Theorem 6) We demonstrate that the statement of the theorem holds for the family of cnfs $\psi(G_{n},E_{hor},E_{vert})$ with $n>0$. For the lower bound, let $B$ be a $\wedge_{d}$-obdd representing $\psi(G_{n},E_{hor},E_{vert})$. Let $\pi_{0}$ be a linear order respected by $B$. Let $\pi^{*}$ be the linear order on $V(G^{*})$ induced by $\pi_{0}$ (that is, two elements are ordered by $\pi^{*}$ in exactly the same way as they are ordered by $\pi_{0}$).
By Lemma 9, we can assume w.l.o.g, that $\pi^{*}$ is neatly crossed by an induced matching $M$ of $G_{n}[E_{hor}]^{*}$ of size $\Omega(lsimw(G_{n}))$. By Proposition 3, $|M|=\Omega(n)$. Then, by Theorem 10, we conclude that
Claim 1.
A $\wedge_{d}$-obdd representing $\psi(G_{n}[E_{hor}])$ and respecting $\pi^{*}$ is of size $2^{\Omega(n)}$.
Let ${\bf g}=\{(\textsc{jn},1)\}$ It is not hard to see that $\psi(G_{n},E_{hor},E_{vert})|_{\bf g}=\psi(G_{n}[E_{hor}])$ and the latter does not have non-essential variables. By Lemma 12, a $\wedge_{d}$-obdd representing $\psi(G_{n}[E_{hor}])$ and respecting $\pi^{*}$ can be obtained from $B$ by a transformation that does not increase its size By Claim 1, the size of $B$ is $2^{\Omega(n)}$ as required.
For the upper bound, we consider an fbdd whose source $rt$ is associated with jn. Let $u_{0}$ and $u_{1}$ be the children of $rt$ with respective edges $(rt,u_{0})$ and $(rt,u_{1})$ labelled with 0 and $1$. Then $u_{0}$ is the source of an fbdd representing $\psi(G_{n}[E_{vert}])$ and $u_{1}$ is the source of an fbdd representing $\psi(G_{n}[E_{hor}])$. It is not hard to see that the resulting fbdd indeed represents $\psi(G_{n},E_{hor},E_{vert})$. It remains to demonstrate that both $\psi(G[E_{hor}])$ $\psi(G_{n}[E_{vert}])$ can be represented by poly-size fbdds. We demonstrate that the incidence pathwidth of both $\psi(G_{n}[E_{hor}])$ and $\psi(G_{n}[E_{vert}])$ is at most $7$. It then follows [17] that both these cnfs can be represented by linear-size fbdds.
Let $\pi$ be the linear order over $V(G)=V(G[E_{hor}])$ where the vertices occur in the ’dictionary’ order: $(i_{1},j_{1})$ precedes $(i_{2},j_{2})$ if $i_{1}<i_{2}$ or $i_{1}=i_{2}$ and $j_{1}<j_{2}$ and let $\pi_{i}$ be the vertex in the position $i$ of $\pi$. That is, $\pi=(\pi_{1},\dots,\pi_{n^{2}})$. We consider a path decomposition $(P,{\bf B})$ where $P=\pi_{1},\dots,\pi_{n^{2}}$ and for each $u\in V(P)$, ${\bf B}(u)$ is defined as follows.
If $u=(i,1)$ for $1\leq i\leq n$ then ${\bf B}(u)=\{u[1],u[2],\neg V(G)[1],\neg V(G)[2]\}$ (recall that the last two elements correspond to the long clauses). If $u=(i,j)$ for $1\leq i\leq n$ and $2\leq j\leq n$, let v=(i,j1)v=(i,j-1). Then ${\bf B}(u)=\{u[1],u[2],v[1],v[2],(u[1]\vee v[2]),(u[2]\vee v[1]),\neg V(G)[1],\neg V(G)[2]\}$. It follows from a direct inspection that $(P,{\bf B})$ is a tree decomposition of the incidence graph of $\psi(G_{n}[E_{hor}])$ of width $7$. For $\psi(G_{n}[E_{vert}])$, the construction is analogous with rows and columns changing their roles. ∎
Proof.
(Theorem 7) We consider the class of cnfs $\varphi(G_{n},E_{hor},E_{vert})$. Let $B$ be an obdd representing $\varphi(G_{n},E_{hor},E_{vert})$. Let $\pi_{0}$ be an order respected by $B$ and let $\pi$ be a linear order of $V(G_{n})$ induced by $\pi_{0}$. By Lemma 9, we can assume w.l.o.g. the existence of an induced matching $M$ of $\varphi(G_{n}[E_{hor}])$ of size at least $\Omega(n)$ crossing $\pi$. By Theorem 11, an obdd respecting $\pi$ and representing $\varphi(G_{n}[E_{hor}])$ is of size $2^{\Omega(n)}$. As all the variables of $\varphi(G_{n}[E_{hor}])$ are essential, it follows from Lemma 12 that an obdd representing $\varphi(G_{n}[E_{hor}])$ can be obtained from $B$ by a transformation that does not increase its size (note that the transformation does not introduce conjunction nodes hence, being applied to an obdd produces another obdd). We thus conclude that the size of $B$ is at least $2^{\Omega(n)}$.
For the upper bound, we demonstrate existence of a poly-size $\wedge_{d}$-obdd representing $\varphi(G_{n}E_{hor},E_{vert})$ and respecting the order $\pi_{d}$ where the first element is jn followed by $V(G_{n})$ in the dictionary order, that is $(i_{1},j_{1})$ precedes $(i_{2},j_{2})$ if $i_{1}<i_{2}$ or $i_{1}=i_{2}$ and $j_{1}<j_{2}$.
Let $Hor_{i}$ be the subgraph of $G_{n}$ induced by row $i$, that is the path $(i,1),\dots(i,n)$ and $Vert_{i}$ be the subgraph of $G_{n}$ induced by the column $i$, that is the path $(1,i),\dots,(n,i)$. By direct layerwise inspection, we conclude that
Claim 2.
There are linear sized obdds respecting $\pi_{d}$ and representing $\varphi(Hor_{i})$ and $\varphi(Vert_{i})$ for each $1\leq i\leq n$.
Proof.
In order to represent $\mathsf{var}(Hor_{i})$, let us query the variables by the ascending order of their second coordinate, that is $(i,1),\dots,(i,n)$. For each $1\leq j\leq n-1$, let ${\bf H}_{j}$ be the set of all functions $\varphi(Hor_{i})|_{\bf g}$ where $\bf g$} is an assignment over $(i,1),\dots,(i,j)$. It is not hard to see that that there are at most $3$ such functions: constant zero, and two functions on $(i,j+1),\dots,(i,n)$ completely determined by the assignment of $(i,j)$. It is well known [25] that the resulting obdd can be seen as a dag with layers $1,\dots,n+1$, the last layer is reserved for the sink nodes, and each layer $1\leq j\leq n$ containing nodes labelled with $(i,j)$ and having $O(1)$ size.
The argument for $\varphi(Vert_{i})$ is similar with the first coordinate of its variables being used instead of the second one. ∎
Let $B^{h}_{i}$ and $B^{v}_{i}$ be obdds respecting $\pi_{d}$ and representing $Hor_{i}$ and $Vert_{i}$ as per Claim 2. By using $O(n)$ conjunction nodes, it is easy to create $\wedge_{d}$-obdds $B_{1}$ and $B_{0}$ that are, respectively conjunctions of all $B^{v}_{i}$ and all $B^{h}_{i}$ and hence, respectively represent $\wedge_{1\leq i\leq n}\varphi(Hor_{i})$ and $\wedge_{1\leq i\leq n}\varphi(Vert_{i})$. It is not hard to see that the former conjunction is $\varphi(G[E_{hor}])$ and the latter one is $\varphi(G[E_{vert}])$, hence they are, respectively represented by $B_{1}$ and $B_{0}$. The resulting $\wedge_{d}$-obdd is then constructed as follows. The source $rt$ is a decision node labelled with jn. Let $u_{1}$ and $u_{0}$ be the children of $rt$ labelled by $1$ and 0 respectively. Then $u_{1}$ is the source of $B_{1}$ and $u_{0}$ is the source of $B_{0}$. ∎
Remark 7.
Note that the variables of a single column of $G_{n}$ explored along the column satisfy the dictionary order. However, if we try to explore first one column and then another one then the dictionary order will be violated. Thus, the upper bound in the proof of Theorem 7 demonstrates how the conjunction nodes overcome the rigidity of obdd by splitting the set of all variables in chunks that do respect a fixed order.
3.3 Proofs of Theorem 10 and 11
In this subsection we prove Theorems 10 and 11 using the approach as presented in Section 3.1.
We first prove Theorems 10. For that purpose, we define a set $\mathcal{F}$ of assignments of $\psi(G)$ along with their $UB$ sets. Then we demonstrate that $\mathcal{F}$ is a fooling set and its size is exponential in $|M|$. The required lower bound will then immediately follow from Lemma 4. W.l.o.g. we assume existence of sets $U,W\subseteq V(G)$. such that $M$ is a matching between $U[1]$ and $W[2]$ and that each element of $U[1]$ is ordered by $\pi^{*}$ before each element of $W[2]$. We let $\pi_{0}$ be the prefix of $\pi^{*}$ whose last element is the last element of $U[1]$ in $\pi^{*}$.
Definition 15.
We define $\mathcal{F}$ as the set of all assignments over $\pi_{0}$ satisfying the following two conditions.
  1. 1.
    At least one variable of $U[1]$ is assigned with 0 and at least one variable of $U[1]$ is assigned with $1$.
  2. 2.
    All the variables of $\pi_{0}\setminus U[1]$ are assigned with $1$.
Example 6.
Let $G$ be a matching consisting of three edges $\{u_{1},w_{1}\},\{u_{2},w_{2}\},\{u_{3},w_{3}\}$ Let $\pi^{*}=(u_{1}[1],u_{2}[1],u_{1}[2],u_{2}[2],w_{1}[1],w_{2}[1],u_{3}[1],w_{1}[2],w_{2}[2],w_{3}[2],u_{3}[2],w_{3}[1])$ Let $M=\{\{u_{1}[1],w_{1}[2]\},\{u_{2}[1],w_{2}[2]\},\{u_{3}[1],w_{3}[2]\}$. In this specific case, $U=\{u_{1}[1],u_{2}[1],u_{3}[1]\}$ and $W=\{w_{1}[2],w_{2}[2],w_{3}[2]\}$. Further on $\pi_{0}=(u_{1}[1],u_{2}[1],u_{1}[2],u_{2}[2],w_{1}[1],w_{2}[1],u_{3}[1])$. Now, $\mathcal{F}$ consists of all the assignments that map all of $u_{1}[2],u_{2}[2],w_{1}[1],w_{1}[2]$ to $1$ and assign $u_{1}[1],u_{2}[1],u_{3}[1]$ so that exactly two of them are assigned with $1$.
Before we proceed, we extend the notation by letting $U=\{u_{1},\dots,u_{q}\}$, $W=\{w_{1},\dots,w_{q}\}$ and $M=\{\{u_{1}[1],w_{1}[2]\},\dots,\{u_{q}[1],w_{q}[2]\}\}$. Further on, for $I\subseteq\{1,\dots,q\}$, we let $U_{I}=\{u_{i}|i\in I\}$, $U[1]_{I}=\{u_{i}[1]|i\in I\}$, and $U[2]_{I}=\{u_{i}[2]|i\in I\}$. The sets $W_{I}$, $W[1]_{I}$, and $W[2]_{I}$ are defined accordingly.
Definition 16.
Let $\bf g$$\in\mathcal{F}$. We denote by $I({\bf g})$ the set of all $1\leq i\leq q$ such that $u_{i}[1]$ is assigned with $1$ by $\bf g$}. We refer to $W[2]_{I({\bf g})}$ as $UB({\bf g})$.
We are now going to prove that the assignments $\bf g$$\in\mathcal{F}$ equipped with the sets $UB({\bf g})$ form a fooling set. For this purpose, we will use assignments of a special form as defined below.
Definition 17.
Let $\bf g$$\in\mathcal{F}$. Let $J\subseteq I({\bf g})$. We denote by ${\bf h}_{J}[{\bf g}]$ the assignment over $V(G^{*})$ such that ${\bf g}\subseteq{\bf h}_{J}[{\bf g}]$, for each $u\in V(G^{*})\setminus\mathsf{var}({\bf g})$, ${\bf h}_{J}[{\bf g}]=0$ if $u\in W[2]_{J}$ and 11 otherwise.
Example 7.
Continuing on Example 6, let ${\bf g}=\{(u_{1}[1],0),(u_{2}[1],1),(u_{3}[1],1),\\ (u_{1}[2],1),(u_{2}[2],1),(w_{1}[1],1),(w_{2}[1],1)\}$. Then $I({\bf g})=\{2,3\}$ and $UB({\bf g})$ is $\{w_{2}[2],w_{3}[2]\}$. For $J=I({\bf g})$, ${\bf h}_{J}({\bf g})$ is the extension of $\bf g$} assigning $w_{2}[2]$ and $w_{3}[2]$ with 0 and assigning with $1$ the remaining variables that are not assigned by $\bf g$}.
We are now proving an auxiliary lemma about the assignments ${\bf h}_{J}[{\bf g}]$.
Lemma 13.
Let $\bf g$$\in\mathcal{F}$. Let $J\subseteq I({\bf g})$. ${\bf h}_{J}[{\bf g}]$ satisfies $\psi(G)$ if and only if $J\neq\emptyset$. In particular, ${\bf h}_{I}[{\bf g}]$ satisfies $\psi(G)$.
Proof.
First of all, if $J=\emptyset$ then ${\bf h}_{J}[{\bf g}]$ falsifies the clause $\neg V(G)[2]$. Indeed, the only variables of $V(G)[2]$ assigned by ${\bf h}_{J}[{\bf g}]$ with 0 are those of $W[2]_{J}$. But $J=\emptyset$ implies that $W[2]_{J}=\emptyset$.
Otherwise, we observe that $\neg V(G)[1]$ is satisfied because one of $U[1]$ is assigned negatively and $\neg V(G)[2]$ is satisfied because a variable of $W[2]_{J}$ is assigned negatively. Hence, $\psi(G)$ may be falsified only if both variables of a clause $(u[1]\vee w[2])$ are assigned with zeroes. By assumption $w[2]=w_{j}[2]$ for some $j\in J$. Further on, $u_{j}[1]$ assigned with $1$ by $\bf g$}, but on the other hand, the only variables of $V(G)[1]$ assigned by zeroes belong to $U[1]$. Hence $u[1]=u_{i}[1]$ for some $i\neq j$. By definition of $\psi(G)$, $G^{*}$ has an edge $\{u_{i}[1],w_{j}[2]\}$ in contradiction to $M$ being an induced matching.
For the second statement, we note that $I({\bf g})\neq\emptyset$ because, by definition of $\bf g$}, at least one variable of $U[1]$ is assigned with $1$. Then we apply the first statement, ∎
Now we are ready to prove that the $UB$ sets are indeed unbreakable.
Lemma 14.
For each $\bf g$$\in\mathcal{F}$, $\mathcal{S}(B)|_{\bf g}$ does not break $UB({\bf g})=W[2]_{I({\bf g})}$.
Proof.
Assume the opposite. Then $\mathcal{S}(B)|_{\bf g}=\mathcal{S}_{1}\times\mathcal{S}_{2}$ such that there is a partition $I_{1},I_{2}$ of $I({\bf g})$ so that $\mathsf{var}(\mathcal{S}_{j})\cap W[2]_{I({\bf g})}=W[2]_{I_{j}}$ for each $j\in\{1,2\}$.
By Lemma 13, ${\bf h}_{2}=Proj({\bf h}_{I_{1}}[{\bf g}],\mathsf{var}(\mathcal{S}_{2}))\in\mathcal{S}_{2}$ and ${\bf h}_{1}=Proj({\bf h}_{I_{2}}[{\bf g}],\mathsf{var}(\mathcal{S}_{1}))\in\mathcal{S}_{1}$. Then ${\bf g}\cup{\bf h}_{1}\cup{\bf h}_{2}$ is a satisfying assignment of $\psi(G)$. However, this is a contradiction since ${\bf g}\cup{\bf h}_{1}\cup{\bf h}_{2}$ sets all the variables of $V(G)[2]$ to $1$ thus falsifying $\neg V(G)[2]$. ∎
We next prove that the distinct projections property holds for $\mathcal{F}$.
Lemma 15.
Let ${\bf g}_{1},{\bf g}_{2}$ be two distinct elements of $\mathcal{F}$. Then $Proj(\mathcal{S}(B)|_{{\bf g}_{1}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))\neq Proj(\mathcal{S}(B)|_{{\bf g}_{2}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))$.
Proof.
Assume the opposite and let ${\bf g}_{1},{\bf g}_{2}$ be two distinct elements of $\mathcal{F}$ such that $Proj(\mathcal{S}(B)|_{{\bf g}_{1}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))=Proj(\mathcal{S}(B)|_{{\bf g}_{2}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))$.
Let $u_{i}[1]$ be a variable assigned differently by ${\bf g}_{1}$ and ${\bf g}_{2}$. Assume w.l.o.g. that ${\bf g}_{1}(u_{i}[1])=0$ while ${\bf g}_{2}(u_{i}[1])=1$. That is, $i\in I({\bf g}_{2})$. Let $I=\{i\}$. By Lemma 13, ${\bf h}_{I}[{\bf g}_{2}]$ satisfies $\psi(G)$. Let ${\bf h}_{0}=Proj({\bf h}_{I}[{\bf g}_{2}],UB({\bf g}_{1})\cup UB({\bf g}_{2}))$ As, by assumption, ${\bf h}_{I}[{\bf g}_{2}]$ is an extension of ${\bf g}_{2}$, ${\bf h}_{0}\in Proj(\mathcal{S}(B)|_{{\bf g}_{2}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))$ and hence, by assumption in the first paragraph, ${\bf h}_{0}\in Proj(\mathcal{S}(B)|_{{\bf g}_{1}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))$.
This means that ${\bf g}_{1}\cup{\bf h}_{0}$ can be extended to a satisfying assignment of $\psi(G)$. However, ${\bf g}_{1}$ maps $u_{i}[1]$ to 0 simply by definition. On the other hand, ${\bf h}_{0}$, being a subset of ${\bf h}_{I}[{\bf g}_{2}]$, assigns $w_{i}[2]$ with zero thus falsifying the clause $(u_{i}[1]\vee w_{i}[2])$ of $\psi(G)$. ∎
Proof.
(of Theorem 10) It follows from the combination of Lemmas 14 and 15 that $\mathcal{F}$ is a fooling set. It thus follows from Lemma 4 that $|B|\geq|\mathcal{F}|$.
It remains to notice that $\{Proj({\bf g},U[1])|{\bf g}\in\mathcal{F}\}$ includes all possible assignments to $U_{1}$ but all zeroes and all ones. It thus follows that $|\mathcal{B}|\geq 2^{q}-2$. ∎
Proof.
(of Theorem 11) Let $\pi$ be the order respected by $B$. Let $M$ be a matching crossing $\pi$. Let $U$ and $W$ be subsets of $V(G)$ such that $M$ is a matching between $U$ and $W$ and all the vertices of $U$ occur in $\pi$ before all the vertices of $W$. Let $\pi_{0}$ be the prefix of $\pi$ ending with the last vertex of $U$ in $\pi$. Let $\mathcal{F}_{0}$ be the set of all the assignments that assign at least one variable of $U$ with zero and assign all the variables of $\pi_{0}\setminus U$ with $1$. Let ${\bf g}\in\mathcal{F}_{0}$. We observe that $\bf g$} can be extended to a satisfying assignment of $\varphi(G)$. Indeed, let $\bf h$} be an extension of $\bf g$} to $Var(\varphi)$ just assigning with $1$ all the variables unassigned by $\bf g$}. Then a clause can be falsified by $\bf h$} only if it includes two variables of $U$. But such a clause does not exist due to $M$ being an induced matching. On the other hand, there is an extension of $\bf g$} that falsifies a clause of $\varphi(G)$. Indeed, let $u$ be a variable of $U$ assigned with 0 by $\bf g$}. Let $v$ be such that $\{u,v\}\in M$. Extension of $\bf g$} by assigning $v$ with 0 falsifies the clause $(u\vee v)$. It follows from Lemma 3 that $L({\bf g})\neq\emptyset$. Further on, it follows from Lemma 2 (since $B$ does not have conjunction nodes) that $|L({\bf g})|=1$. Denote the only element of $L({\bf g})$ by $u({\bf g})$. We prove that the mapping of $\mathcal{F_{0}}$ from $\bf g$} to $u({\bf g})$ is injective thus implying that $|B|\geq|\mathcal{F}_{0}|$. As $|\mathcal{F}_{0}|\geq 2^{|M|}-1$, the theorem will immediately follow.
So, assume that there are ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}_{0}$ such that $u=u({\bf g}_{1})=u({\bf g}_{2})$. Let $X=\mathsf{var}(\varphi)\setminus(\pi_{0}\cup\mathsf{var}(u))$. It follows from Lemma 3 that $\mathcal{S}(B)|_{{\bf g}_{1}}=\mathcal{S}(B)|_{{\bf g}_{2}}=X^{\{0,1\}}\times\mathcal{S}(B_{u})$. We prove that the restrictions of $\mathcal{S}(B)$ to ${\bf g}_{1}$ and ${\bf g}_{2}$ are in fact distinct. Indeed, let $u\in U$ be such that ${\bf g}_{1}(u)\neq{\bf g}_{2}(u)$. Assume w.l.o.g. that $g_{1}(u)=1$ and $g_{2}(u)=0$. Let $v$ be such that $\{u,v\}\in M$. Let ${\bf h}_{1}$ be an assignment to $\mathsf{var}(\varphi)\setminus\pi_{0}$ that assigns $v$ with 0 and the remaining variables with $1$. We observe that ${\bf g}_{1}\cup{\bf h}_{1}$ is a satisfying assignment of $\varphi(G)$ meaning that ${\bf h}_{1}\in\mathcal{S}(B)|_{{\bf g}_{1}}$. Indeed, by the same argument that we use in the previous paragraph for assignment $\bf h$}, ${\bf g}_{1}\cup{\bf h}_{1}$ can only falsity a clause $(u^{\prime}\cup v)$ such that $u^{\prime}\in U$ and $u^{\prime}\neq u$. But such a clause does not exist because $M$ is an induced matching.
It thus follows from our assumption that ${\bf h}_{1}\in\mathcal{S}(B)|_{{\bf g}_{2}}$. Hence, ${\bf g}_{2}\cup{\bf h}_{1}$ is a satisfying assignment of $\varphi(G)$. However, this is a contradiction since ${\bf g}_{2}\cup{\bf h}_{1}$ falsifies the clause $(u\vee v)$. ∎
4 The Apply operation for $\wedge_{d}$-obdd
4.1 Upper and Lower bounds
An important property of obdd is efficiency of conjunction (Apply operation) of two obdds obeying the same order. In particular, if $B_{1}$ and $B_{2}$ are two obdds over the same set of variables and obeying the same order $\pi$ then the function $f(B_{1})\wedge f(B_{2})$ can be represented by an obdd obeying $\pi$ of size $|B_{1}|\cdot|B_{2}|$ and can be computed at time $O(|B_{1}|\cdot|B_{2}|)$.
In this section we will first demonstrate that, in general $\wedge_{d}$-obdd does not allow efficient $Apply$ operation. We will then identify a restricted case where $Apply$ can be carried out efficiently. The corresponding main result is stated in Theorem 19, the theorem is proved in Subsection 4.2. The corresponding lower bound is stated in Theorem 20, and we prove the theorem in Subsection 4.3.
In the Conclusion section, we argue that this case gives rise to an interesting theory and pose the related open questions.
Recall that $G_{n}$ is an $n\times n$ grid. Let $V_{n}=V(G_{n})$. We let $E_{vert}$ and $E_{hor}$ be the sets of vertical and horizontal edges of $G_{n}$. Let $G_{vert}=(V_{n},E_{vert})$ and $G_{hor}=(V_{n},E_{hor})$. Let $\pi=\pi_{n}$ be a dictionary order over $V_{n}$. That is $(i_{1},j_{1})<_{\pi}(i_{2},j_{2})$ if and only if either $i_{1}<i_{2}$ or $i_{1}=i_{2}$ and $j_{1}<j_{2}$.
Lemma 16.
  1. 1.
    There is linear size $\wedge_{d}$-obdd respecting $\pi$ and representing $\varphi(G_{vert})$.
  2. 2.
    There is linear size $\wedge_{d}$-obdd respecting $\pi$ and representing $\varphi(G_{hor})$.
Proof.
We reuse the proof of Theorem 7. In particular, the $\wedge_{d}$-obdd $B_{0}$ and $B_{1}$ considered in the proof are the desired representations of $\varphi(G_{hor})$ and $\varphi(G_{vert})$ respectively. ∎
Theorem 17.
The Apply operation for $\wedge_{d}$-obdd respecting the same order may lead to exponential explosion.
Proof.
It is not hard to see that $\varphi(G_{n})=\varphi(G_{vert})\wedge\varphi(G_{hor})$. By Lemma 16, both $\varphi(G_{vert})$ and $\varphi(G_{hor})$ are polynomially representable by $\wedge_{d}$-obdds.However, since the treewidth of $G_{n}$ is $\Omega(n)$, $\varphi(G_{n})$ requires an exponential representation as a dnnf [2, 12] let alone an $\wedge_{d}$-obdd. ∎
Thus we have demonstrated that, in general, $\wedge_{d}$-obdd does not support efficient Apply operation. Nevertheless, the operation can be carried out efficiently in special cases. We are going to consider one such case.
First, let us make several notational conventions. For a linear order $\pi$ and $u\in\pi$, we denote by $\pi^{>u}$ the suffix of $\pi$ starting at the immediate successor of $u$.
Definition 18.
Let $B$ be a $\wedge_{d}$-fbdd and let $\pi$ be a linear order such that $\mathsf{var}(B)\subseteq\pi$. We say that $B$ is embeddable in $\pi$ if the following recursively defined conditions hold.
  1. 1.
    $B$ consists of a single sink node.
  2. 2.
    Suppose that the source $u$ of $B$ is a variable node labelled with a variable $x$. Let $u_{0}$ and $u_{1}$ be the children of $u$. Then both $B_{u_{0}}$ and $B_{u_{1}}$ are embeddable into $\pi^{>x}$.
  3. 3.
    Suppose that the source $u$ of $B$ is a $\wedge_{d}$ node and let $u_{1}$ and $u_{2}$ be the children of $u$. Then there are disjoint intervals $\pi_{1}$ and $\pi_{2}$ of $\pi$ such that $B_{u_{1}}$ is embeddable in $\pi_{1}$ and $B_{u_{2}}$ is embeddable in $\pi_{2}$.
It is not hard to demonstrate by induction that if $B$ is embeddable $\pi$ then $B$ is a $\wedge_{d}$-obdd respecting $\pi$. The opposite however is not true in general. Moreover, embeddability does not imply structurudness as different $\wedge_{d}$-nodes can break the same sequence of variables into different intervals. This intuition is illustrated in the following example.
Example 8.
Figure 4 illustrates two $\wedge_{d}$-obdds over variables $(x_{1},\dots,x_{9})$ and respecting the order as listed. The ellipses correspond to obdds and the sequences of variables provided near each ellipses are their corresponding respected orders.
We observe that the $\wedge_{d}$-obdd on the left is embeddable into $(x_{1},\dots,x_{9})$. Indeed, the sink is a decision node labelled by the first variable in the order. Each of the two $\wedge_{d}$-nodes break the remaining sequence $(x_{2},\dots,x_{9})$ into intervals and the orders of variables in the intervals are respected by the ‘bottom layer’ obdds. On the other hand, the $\wedge_{d}$-obdd on the left-hand side is not structured [16]. Indeed, one $\wedge_{d}$-node breaks the variables into intervals $(x_{2},\dots,x_{6})$ and $(x_{7},\dots,x_{9})$, whole the other $\wedge_{d}$-node breaks the variables into intervals $(x_{2},\dots,x_{4})$ and $(x_{5},\dots,x_{9})$. It is not hard to see that, as a result, there is no vtree respected by the model.
Finally, we observe that the $\wedge_{d}$-obdd on the right-hand side is not embeddable into $(x_{1},\dots,x_{9})$ since the $\wedge_{d}$-node decomposes the set of variables into two classes that are not intervals.
$(x_{2},\dots,x_{6})$$(x_{7},\dots,x_{9})$$(x_{2},\dots,x_{4})$$(x_{5},\dots,x_{9})$$(x_{2},x_{4},x_{6},x_{8})$$(x_{1},x_{3},x_{5},x_{7},x_{9})$$\wedge_{d}$$\wedge_{d}$$x_{1}$$\wedge_{d}$
Figure 4: Embeddability vs non-embeddability
We remark that it is not hard to see that for ordinary obdds, the notions of respecting an order and being embeddable into that order coincide. In fact, a stronger statement (Lemma 18) can be made in presence of the notion of irregularity index.
The embeddability restriction allows efficient simulation of $\wedge_{d}$-obdd by an ordinary obdd, The degree of efficiency is regulated by a parameter defined below that we refer to as the irregularity index.
Definition 19.
Suppose that $\pi$ is embeddable into $\pi$. Then the irregularity index of $B$ w.r.t. $\pi$ denoted by $ir_{\pi}(B)$ is defined as follows.
  1. 1.
    If $B$ consists of a single sink node then $ir_{\pi}(B)=0$.
  2. 2.
    Suppose that the source $u$ of $B$ is a variable node labelled with a variable $x$. Let $u_{0}$ and $u_{1}$ be the children of $u$. Then $ir_{\pi}(B)=max(ir_{\pi^{>x}}(B_{0}),ir_{\pi^{>x}}(B_{1}))$.
  3. 3.
    Suppose that the source $u$ of $B$ is a $\wedge_{d}$ node and let $u_{1}$ and $u_{2}$ be the children of $u$. Let $\pi_{1}$ and $\pi_{2}$ be disjoint intervals of $\pi$ such that $B_{u_{1}}$ is embeddable in $\pi_{1}$ and $B_{u_{2}}$ is embeddable in $\pi_{2}$. Assume that $\pi_{1}$ occurs before $\pi_{2}$ in $\pi$. Then $ir_{\pi}(B)=max(ir_{\pi_{1}}(B_{u_{1}})+1,ir_{\pi_{2}}(B_{u_{2}}))$.
As the irregularity index greater than 0 requires presence of conjunction gates, we observe the following.
Lemma 18.
Let $\pi$ be an obdd respecting an order $\pi$ Then $B$ is embeddable into $\pi$ with the irregularity index 0.
In fact, a much more general statement holds. In particular, it turns out that a $\wedge_{d}$-fbddembeddable into a linear order $\pi$ can be polynomially simulated by an obdd with the iiregularity index plus one being the degree of the polynomial. The formal statement is provided below.
Theorem 19.
Let $B$ be a $\wedge_{d}$-fbdd embeddable into a linear order $\pi$. Then $B$ can be represented as an obdd respecting $\pi$ and having size at most $|B|^{k+1}$ where $k=ir_{\pi}(B)$. Moreover, there is an algorithm that, given $B$ and $\pi$ returns such an obdd in time polynomial in $|B|$.
This simulation is best possible in the following sense.
Theorem 20.
For each $n\geq 1$, there is a $\wedge_{d}$-fbdd $B$ of $O(n)$ variables and size $O(n)$ embeddable into an order $\pi_{n}$ with irregularity index $n$, such that any obdd respecting $\pi_{n}$ that simulates $B$ must be of size at least $2^{\Omega(n)}$.
Theorems 19 and 20 are proved in Subsections 4.2 and 4.3 respectively. In order to prove Theorem 19, we use the same approach as was used in [3] and then in [18] to simulate variants of dnnf by corresponding variants of read-once branching programs.
It follows from Theorem 19 that if two $\wedge_{d}$-fbdds are embeddable into the same order with low irregularity indices, the $Apply$ operation for them can be carried out efficiently. The formal statement is provided below.
Theorem 21.
Let $B_{1}$ and $B_{2}$ be two $\wedge_{d}$-fbdds embeddable into the same order $\pi$. Suppose that $\mathsf{var}(B_{1})=\mathsf{var}(B_{2})$. Then $f(B_{1})\wedge f(B_{2})$ can be represented as an obdd $B$ respecting $\pi$ of size $|B_{1}|^{ir_{\pi}(B_{1})}\cdot|B_{2}|^{ir_{\pi}(B_{2})}$.
Moreover, there is an algorithm that, provided $B_{1},B_{2}$ and $\pi$ as input, returns $B$ in time polynomial in the size of $B$.
Proof.
According to Theorem 19, there is an algorithm that, for each $i\in[2]$, given $B_{i}$ and $\pi$ returns an obdd $B^{\prime}_{i}$ respecting $\pi$ with $f(B^{\prime}_{i})=f(B_{i})$ and of size at most $|B_{i}|^{ir_{\pi}(B_{i})}$. As $B^{\prime}_{1}$ and $B^{\prime}_{2}$ respect the same order, it is well known [25] that the $Apply$ operation for them can be carried out in time $O(|B^{\prime}_{1}|\cdot|B^{\prime}_{2}|)$ and the size of the resulting obdd has the same upper bound. ∎
In light of Theorem 21, it is natural to ask whether the $Apply$ operation can be efficiently carried out for two $\wedge_{d}$-obdds respecting the same order $\pi$ but not necessarily embeddable into it. In the Conclusion section, we conjecture that this is indeed the case and provide supporting evidence. We also argue that embeddability property is of a broad interest in the knowledge compilation context. In particular, the embeddability can be formulated for dnnfs and it is a wekaer restriction than structuredness.
4.2 Proof of Theorem 19
The following statements are not hard to derive from the embeddability definition.
Lemma 22.
Let $B$ be a $\wedge_{d}$-fbdd embeddable into a linear order $\pi$. Let $u$ be a $\wedge_{d}$-node of $B$ and let $u_{1}$ and $u_{2}$ be the children of $u$. Then there are disjoint intervals $\pi_{1}$ and $\pi_{2}$ of $\pi$ such that $Var((B_{u_{i}})\subseteq\pi_{i}$ for each $i\in[2]$.
Lemma 23.
Let $B$ be a $\wedge_{d}$-fbdd embeddable into a linear order $\pi$. Let $u\in V(B)$. Then $B_{u}$ is embeddable into $\pi$ $ir_{\pi}(B_{u})\leq ir_{\pi}(B)$.
Definition 20.
Let $B$ be a $\wedge_{d}$-fbdd embeddable into a linear order $\pi$. Let $u$ be a $\wedge_{d}$-node of $B$ and let $u_{1}$ and $u_{2}$ be the children of $u$. Let $\pi_{1}$ and $\pi_{2}$ be the intervals as in Lemma 22. Assume w.l.o.g. that $\pi_{1}$ occurs on $\pi$ before $\pi_{2}$. Then we call $(u,u_{1})$ a top edge and $(u,u_{2})$ a bottom edge. Each edge whose tail is a decision node is a neutral edge.
Example 9.
Consider the left-hand side $\wedge_{d}$-obdd on Figure 4 (see Example 8 for the description of this model). The left-hand side children of the $\wedge_{d}$-nodes (as shown on the picture) are the heads of top edges and the right-hand side children of these nodes are the heads of bottom edges.
The notion of top edges gives rise to the irregularity indices of paths and nodes as defined below.
Definition 21.
Let $B$ be a $\wedge_{d}$-fbdd embeddable into a linear order $\pi$. Let $P$ be a path of $B$. The irregularity index of $P$ w.r.t. $B$ denoted by $ir_{\pi,B}(P)$ is the number of top edges in $P$.
Further on, for $u\in V(B)$, the irregularity index of $u$ denoted by $ir_{\pi,B}(u)$ is the largest $ir_{\pi,B}(P)$ over all paths $P$ from the source of $B$ to $u$.
Lemma 24.
With the notation as in Definition 21, $ir_{\pi,B}(P)\leq ir_{\pi}(B)$. In particular, for each $u\in V(B)$, $ir_{\pi,B}(u)\leq ir_{\pi}(B)$.
Proof.
By induction on $ir_{\pi,B}(P)$. If $ir_{\pi,B}(P)=0$ then the statement trivially holds since $ir_{\pi}(B)\geq 0$ by definition. Assume now that $ir_{\pi,B}(P)>0$. Let $P^{\prime}$ be the shortest suffix of $P$ such that $ir_{\pi,B}(P)=ir_{\pi,B}(P^{\prime})$. Let $u$ be the first node of $P^{\prime}$. Then $u$ must be a $\wedge_{d}$-node due to the minimality of $P^{\prime}$. Let $v$ be the immediate successor of $u$ on $P^{\prime}$. Then $(u,v)$ must be a top edge of $B$ due to minimality of $P^{\prime}$. Let $P^{\prime\prime}$ be the suffix of $P$ starting from $v$. We conclude that $ir_{\pi,B}(P^{\prime\prime})=ir_{\pi,B}(P^{\prime})-1$. Next, $ir_{\pi,B}(P^{\prime\prime})=ir_{\pi,B_{v}}(P^{\prime\prime})$. Hence, by the induction assumption combined with Lemma 23, $ir_{\pi,B}(P^{\prime\prime})\leq ir_{\pi}(B_{v})$. By definition of the irregularity index, $ir_{\pi}(B_{u})\geq ir_{\pi}(B_{v})+1\geq ir_{\pi,B}(P)$. By another application of Lemma 23 $ir_{\pi}(B)\geq ir_{\pi}(B_{u})$ hence the lemma holds. ∎
We are now going to present the transformation from $\wedge_{d}$-fbdd to obdd that we will then use to prove Theorem 19. The key part of the transformation is the following operation of combining two obdds into one.
Definition 22.
Let $B_{1}$ and $B_{2}$ be two obdds such that $\mathsf{var}(B_{1})\cap\mathsf{var}(B_{2})=\emptyset$. Then $conj(B_{1},B_{2})$ is an obdd obtained from $B_{1}$ and $B_{2}$ by the following two operations.
  1. 1.
    Identify the positive sink of $B_{1}$ and the source of $B_{2}$
  2. 2.
    Join the negative sinks of $B_{1}$ and $B_{2}$ into a single vertex.
It is not hard to observe the following.
Lemma 25.
Let $B=conj(B_{1},B_{2})$. Then $f(B)=f(B_{1})\wedge f(B_{2})$.
The transformation also requires preprocessing of the input $\wedge_{d}$-fbdd $B$ that is described below.
  1. 1.
    The name of each decision node $u$ changes into $(u,())$. Throughout the transformation the first component remains invariant and the names of eliminated $\wedge_{d}$ nodes may be appended to the second component.
  2. 2.
    Throughout the transformation the resulting DAG $B^{\prime}$ may have several sources. We identify the main one and refer to it as $source(B^{\prime})$. Initially, $B$ has only one source and, naturally, it is $source(B)$. We thus adapt the definition of $\wedge_{d}$-fbdd as having several sources. In particular, we set $f(B^{\prime})=f(B^{\prime}_{source(B^{\prime})})$.
The transformation is presented in Algorithm 1 as procedure $MakeOBDD(B,\pi)$ The transformation uses the procedure described in Algorithm 2 as an auxiliary subroutine.
Algorithm 1 $MakeOBDD(B,\pi)$
$B_{0}\leftarrow B,i\leftarrow 0$
while $B_{i}$ has $\wedge_{d}$-nodes do
  $i\leftarrow i+1$
  Let $u$ be a lowest $\wedge_{d}$-node (meaning that $B_{u}$ does not have $\wedge_{d}$-node but $u$)
  $B_{i}\leftarrow ConjElim(B_{i-1},u,\pi)$
end while
 Return $B_{i}$
Algorithm 2 $ConjElim(B^{\prime},u,\pi)$
 Let $(u_{1},s_{1})$ and $(u_{2},s_{2})$ be the children of $u$ and assume that the edge $(u,(u_{1},s_{1}))$ is the top one.
 Let ${\bf u}=(u_{1},s_{1})$ and let $D$ be a copy of $B^{\prime}_{\bf u}$ where each name $(u^{\prime},s^{\prime})$ of a node is replaced $(u^{\prime},s^{\prime}+u)$.
$B^{\prime\prime}\leftarrow B^{\prime}$
$D^{*}\leftarrow conj(D,B_{(u_{2},s_{2})})$
 Reconnect edges of $B^{\prime\prime}$ ending at $u$: replace each edge $(y,u)$ with $(y,{\bf u})$.
 Remove $u$ from $B^{\prime}$.
if $u=source(B^{\prime})$ then
  $source(B^{\prime\prime})\leftarrow(u_{1},s_{1}+u)$
else
  $source(B^{\prime\prime})\leftarrow source(B^{\prime})$
end if
The lemma below states several properties of the output of the $ConjElim$ procedure. The first two properties will be needed for proving validity of the $MakeOBDD$ transformation, the last two properties will be needed to establish the required size upper bound of the output of $MakeOBDD$.
Lemma 26.
Let $B^{\prime}$ be a $\wedge_{d}$-fbdd embeddable into an order $\pi$. Suppose that each variable node of $B^{\prime}$ is of the form $(u^{\prime},s^{\prime})$ where $u^{\prime}$ is a name of variable node of $B$ and $s^{\prime}$ is a sequence of $\wedge_{d}$ nodes of $B$. Let $u$ be a lowest $\wedge_{d}$ node of $B^{\prime}$. Let $B^{\prime\prime}=ConjElim(B^{\prime},u,\pi)$. Then the following statements hold.
  1. 1.
    $B^{\prime\prime}$ is a $\wedge_{d}$-fbdd respecting $\pi$
  2. 2.
    $f(B^{\prime\prime})=f(B^{\prime})$.
  3. 3.
    For each ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$, $ir_{\pi,B^{\prime\prime}}({\bf v})\leq ir_{\pi,B^{\prime}}({\bf v})$.
  4. 4.
    For each $(u^{\prime},s^{\prime}+u)\in V(D)$ $ir_{\pi,B^{\prime\prime}}((u^{\prime},s^{\prime}+u))\leq ir_{\pi,B^{\prime}}((u^{\prime},s^{\prime}))-1$ ($D$ is as defined in step 2 of $ConjElim$ algorithm).
Proof.
Let $em$ (standing for ’elimination mapping’) be a bijection mapping each ${\bf v}\in V(B^{\prime})\setminus\{u\}$ to itself and $u$ to $(u_{1},s_{1}+u)$. Then the following statement is immediate by construction.
Claim 3.
Let ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$ be a non-sink node and let ${\bf v}_{1}$ and ${\bf v}_{2}$ be the children of $\bf v$} in $B^{\prime}$. Then $em({\bf v}_{1})$ and $em({\bf v}_{2})$ are the children of $em({\bf v})$. Equivalently, if ${\bf v}_{1}$ and ${\bf v}_{2}$ are the children of $\bf v$} in $B^{\prime\prime}$ then $em^{-1}({\bf v}_{1})$ and $em^{-1}({\bf v}_{2})$ are the children of $\bf v$} in $B^{\prime}$.
The following statement is also immediate by construction.
Claim 4.
$\mathsf{var}(B^{\prime\prime}_{(u_{1},s_{1}+u)})=\mathsf{var}(B^{\prime}_{u})$.
Next, we observe that the $Var$ set remains the same for the subgraphs rooted by the nodes whose names are not changed by the transformation.
Claim 5.
Let ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$. Then $\mathsf{var}(B^{\prime}_{\bf v})=\mathsf{var}(B^{\prime\prime}_{\bf v})$
Proof.
By bottom up induction. The claim is trivially true for the sinks. Suppose that $\bf v$} is not a sink and let ${\bf v}_{1}$ and ${\bf v}_{2}$ be the children of $\bf v$} in $B^{\prime}$. By construction, if $\bf v$} is a variable node that the label of $\bf v$} is the same in $B^{\prime}$ and in $B^{\prime\prime}$. Hence the claim follows from the combination of the induction assumption and Claims 3 and 4. ∎
Now, we begin our reasoning towards establishing the embeddability.
Claim 6.
$B^{\prime\prime}_{(u_{1},s_{1}+u)}$ is a $\wedge_{d}$-fbdd embeddable into $\pi$.
Proof.
By construction and assumption about $B^{\prime}$, there are intervals $\pi_{1}$ and $\pi_{2}$ of $\pi$ such that such that $\pi_{1}$ precedes $\pi_{2}$ in $\pi$ and $B^{\prime}_{(u_{j},s_{j})}$ respects $\pi_{j}$ for each $j\in[2]$. is the same as being embeddable into $\pi$ according to Lemma 18
Claim 7.
For each ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$, $B^{\prime\prime}_{\bf v}$ is a $\wedge_{d}$-fbdd embeddable into $\pi$.
Proof.
The statement is trivially true for sinks so suppose that $\bf v$} with has children ${\bf v}_{1}$ and ${\bf v}_{2}$ in $B^{\prime\prime}$. It follows from the combination Claims 3 and 6 and the induction assumption that each $B^{\prime\prime}_{{\bf v}_{j}}$ is a $\wedge_{d}$-fbdd embeddable into $\pi$.
It follows from the combination of Claims 3, 4, and 5 that for each $j\in[2]$, $\mathsf{var}(B^{\prime}_{em^{-1}({\bf v}_{j})})=\mathsf{var}(B^{\prime\prime}_{{\bf v}_{j}})$. With this in mind, assume first that $\bf v$} is a decision node. Then its label $\mathsf{var}({\bf v})$ is not changed in $B^{\prime\prime}$. As in $B^{\prime}$, all of $\mathsf{var}(B^{\prime}_{em^{-1}({\bf v}_{j})})$ for each $j\in[2]$ are below $\mathsf{var}(B^{\prime}_{\bf v})$, the same dependency preserves for $B^{\prime\prime}$ for all $\mathsf{var}(B^{\prime\prime}_{{\bf v}_{j}})$ for each $j\in[2]$. Combined with the emeddability of as specified in the end of the previoys paragraph, we conclude that $B^{\prime\prime}_{\bf v}$ is embeddable in $\pi$.
It remains to assume that $\bf v$} is a $\wedge_{d}$-node. Assume w.l.o.g. that the edge $({\bf v},em^{-1}({\bf v}_{1}))$ is the top one in $B^{\prime}$. It follows that there are non-overlapping intervals $\pi_{1}$ and $\pi_{2}$ of $\pi$ with $\pi_{1}$ preceding $\pi_{2}$ and such that for each $j\in[2]$ $\mathsf{var}(B^{\prime}_{em^{-1}({\bf v}_{j})})\subseteq\pi_{j}$. We conclude that for each $j\in[2]$ $\mathsf{var}(B^{\prime\prime}_{{\bf v}_{j}})\subseteq\pi_{j}$ and hence $B^{\prime\prime}_{\bf v}$ is embeddable into $\pi$. ∎
Claim 8.
$B^{\prime\prime}_{source(B^{\prime\prime})}$ is a $\wedge_{d}$-fbdd embeddable into $\pi$.
Proof.
[…]
Claim 9.
$f(B^{\prime\prime}_{(u_{1},s_{1}+u)})=f(B^{\prime}_{u})$.
Proof.
Immediate by construction and from Lemmma 25. ∎
Claim 10.
For each ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$, $f(B^{\prime})_{\bf v}=f(B^{\prime\prime})_{\bf v}$
Proof.
We argue by bottom up induction. The statement is trivially true for sinks so suppose that $\bf v$} is a non-sink node of $B^{\prime\prime}$ with children ${\bf v}_{1}$ and ${\bf v}_{2}$. It follows from the combination Claims 3 and 9 and the induction assumption that for each $j\in[2]$ $f(B^{\prime\prime}_{{\bf v}_{j}})=f(B^{\prime}_{em^{-1}({\bf v}_{j})})$.
Assume that $\bf v$} is a variable node labelled with a variable $x$. Assume further that in $B^{\prime}$ the edge $({\bf v},em^{-1}({\bf v}_{1}))$ is labelled with $1$ and the edge $({\bf v},em^{-1}({\bf v}_{2}))$ is labelled with 0. Clearly, in $B^{\prime\prime}$, $\bf v$} remains a variable node labelled with $x$, $({\bf v},{\bf v}_{1})$ is labelled with $1$ and $({\bf v},{\bf v}_{2})$ is labelled with 0. Hence, the statement is immediate from the previous paragraph. If $\bf v$} is a $\wedge_{d}$-node in $B^{\prime}$ it remains so in $B^{\prime\prime}$ and hence the statement is again immediate from the previous paragraph. ∎
The following is immediate by construction
Claim 11.
$B^{\prime\prime}[V(B^{\prime})\cap V(B^{\prime\prime})]=B^{\prime}[V(B^{\prime})\cap V(B^{\prime\prime})]$, in particular the roles of the nodes, the labels on the nodes (if any) and the labels on the edges (if any) are all preserved.
We continue to observe that the roles of outgoing edges of $\wedge_{d}$ notes (the top and the bottom ones) are preserved.
Claim 12.
Let ${\bf v}\in V(B^{\prime})\cap V(B^{\prime\prime})$ be a conjunction node. Let ${\bf w}_{1}$ be a child of $\bf v$} in $B^{\prime\prime}$. Then the role of $({\bf v},{\bf w}_{1})$ in $B^{\prime\prime}$ is the same as the one of $({\bf v},em^{-1}({\bf w}_{1}))$ in $B^{\prime}$
Proof.
Let ${\bf w}_{0}$ be the other child of $\bf v$} in $B^{\prime\prime}$. Due to the embeddability in $\pi$, there are non-overlapping intervals $\pi_{1}$ and $\pi_{0}$ of $\pi$ such that $\mathsf{var}(B^{\prime}_{em^{-1}({\bf w}_{1})})\subseteq\pi_{1}$ and $\mathsf{var}(B^{\prime}_{em^{-1}({\bf w}_{0})})\subseteq\pi_{0}$. The roles of the edges $({\bf v},em^{-1}({\bf w}_{1}))$ and $({\bf v},em^{-1}({\bf w}_{0}))$ are completely determined by the relative location of $\pi_{1}$ and $\pi_{0}$. It follows from the combination of Claims 3, 4, and 5 that $\mathsf{var}(B^{\prime}_{em^{-1}({\bf w}_{1})})=\mathsf{var}(B^{\prime\prime}_{{\bf w}_{1}})$ and $\mathsf{var}(B^{\prime}_{em^{-1}({\bf w}_{0})})=\mathsf{var}(B^{\prime\prime}_{{\bf w}_{0}})$. Hence the role of $({\bf v},{\bf w}_{1})$ is completely determined by the location of $\pi_{1}$ on $\pi_{0}$ relative to $\pi$ and the dependency is the same as in $B^{\prime}$. ∎
The first statement of the lemma follows from the combination of Claims 6 and 8, the second statement follows from the combination of Claims 9 and 10.
For the third statement, let $P$ be a path in $B^{\prime\prime}$ from $source(B^{\prime\prime})$ to $\bf v$} such that $ir_{\pi,B^{\prime\prime}}({\bf v})=ir_{\pi,B^{\prime\prime}}(P)$. If $P$ is also a path in $B^{\prime}$ then the statement is immediate from Claim 12. Otherwise, by Claim 11, $P$ must contain vertices of $D$. By construction, we can identify nodes ${\bf w}_{0}$ and ${\bf w}_{1}$ of $D$ so that $P=P_{1}+P_{2}+P_{3}$, where $P_{1}$ is the prefix of $P$ ending at ${\bf w}_{0}$, $P_{2}$ is the subpath of $P$ between ${\bf w}_{0}$ and ${\bf w}_{1}$, $P_{3}$ is the suffix of $P$ starting at ${\bf w}_{1}$ and $V(P)\cap V(D)=V(P_{2})$. We note that neither $P_{2}$ nor $P_{3}$ contain conjunction nodes. Hence, $ir_{\pi,B^{\prime\prime}}(P)=ir_{\pi,B^{\prime\prime}}(P_{1})$. Further on, by construction $B^{\prime}$ has a path $P^{\prime}=P^{\prime}_{1}+P^{\prime}_{3}$ where $P^{\prime}_{1}$ obtained from $P_{1}$ by replacing its last node with $u$ and $P^{\prime}_{3}$ is obtained from $P_{3}$ by replacing its first node with $u$. We observe that the outgoing edge of $u$ in $P^{\prime}$ is the bottom one and that $ir_{\pi,B^{\prime\prime}}(P_{1})=ir_{\pi,B^{\prime}}(P^{\prime}_{1})$ by Claim 12. We conclude that $ir_{\pi,B^{\prime}}(P^{\prime})=ir_{\pi,B^{\prime\prime}}(P)$. As $P^{\prime}$ is a path in $B^{\prime}$ from $source(B^{\prime})$ to $\bf v$}, this proves the third statement.
For the fourth statement, let $P$ be a path in $B^{\prime\prime}$ from $source(B^{\prime\prime})$ to $(u^{\prime},s^{\prime}+u)$ such that $ir_{\pi,B^{\prime\prime}}(P)=ir_{\pi,B^{\prime\prime}}((u^{\prime}s^{\prime}+u))$. By construction $P=P_{1}+P_{2}$ where $P_{1}$ is the prefix of $P$ ending at $(u_{1},s_{1}+u)$ and $P_{2}$ is the suffix of $P$ starting at $(u_{1},s_{1}+u)$. We note that $P_{2}$ does not contain conjunction nodes and hence $ir_{\pi,B^{\prime\prime}}(P)=ir_{\pi,B^{\prime\prime}}(P_{1})$. Further on, we observe that $P$ can be transformed by a path $P^{\prime}=P^{\prime}_{1}+P^{\prime}_{2}$ from $source(B^{\prime})$ where $P^{\prime}_{1}$ is obtained from $P_{1}$ by replacing the last node with $u$ and $P^{\prime}_{2}$ is obtained from $P_{2}$ by first replacing each node $(u^{*},s^{*}+u)$ with $(u^{*},s)$ and then adding $u$ at the beginning. We note that by Claim 12, $ir_{\pi,B^{\prime\prime}}(P_{1})=ir_{\pi,B^{\prime}}(P^{\prime}_{1})$. Furthermore, by construction, the outgoing edge of $u$ in $P^{\prime}_{2}$ is a top one, hence $ir_{\pi,B^{\prime}}(P^{\prime}_{2})\geq 1$. We conclude that $ir_{\pi,B^{\prime}}(P^{\prime})\geq ir_{\pi,B^{\prime\prime}}(P)+1$ thus establishing the fourth statement. ∎
Proof.
( of Theorem 19) Let $B^{*}$ be the output of $MakeOBDD(B,\pi)$. We are going to prove that $B^{*}$ is an ondd respecting $\pi$ of size at most $|B|^{ir_{\pi}(B)+1}$ and such that $f(B^{*})=f(B)$. As $B^{*}$ is obtained in a constructive way, the second statement of the theorem will be immediate.
It is clear from the description of the algorithm that the algorithm runs $q$ iterations where $q$ is at most the number of conjunction gates of $B$. The output of the algorithm is $B_{q}=B^{*}$. Applying the first two statements of lemma 26 inductively to each $B_{i}$, we conclude that $B_{q}$ is a $\wedge_{d}$-obdd embeddable into $\pi$ and $f(B_{q})=f(B)$. Since by construction, $B_{q}$ contains no conjunction gates, we conclude that $B^{*}$ is an $\wedge_{d}$-fbdd embeddable into $\pi$ and hence an obdd respecting $\pi$.
To establish an upper bound on the size of $B^{*}$, let us first replace $B_{i}$ with $i$ in the subscript for $ir$. That is, instead of $ir_{\pi,B_{i}}({\bf u})$, we write $ir_{\pi,i}({\bf u})$. Also for a decision node ${\bf u}=(u^{\prime},s^{\prime})$, let $len({\bf u})$ be the length of $s^{\prime}$. With this in mind, let us first prove the following claim.
Claim 13.
For each $0\leq i\leq q$ and a decision node ${\bf u}\in V(B_{i})$, $ir_{\pi,i}({\bf u})\leq ir_{\pi}(B)-len({\bf u})$.
Proof.
By induction on $i$. For each decision node ${\bf u}\in B_{0}$, $len({\bf u})=0$. Then, by Lemma 24, $ir_{\pi,0}({\bf u})\leq ir_{\pi}(B)=ir_{\pi}(B)-len({\bf u})$. Hence, the statement holds for $i=0$. Assume that $i>0$. Let ${\bf u}\in V(B_{i-1})\cap V(B_{i})$ be a decision node. By the third statement of Lemma 26 and the induction assumption, $ir_{\pi,i}({\bf u})\leq ir_{\pi,i-1}({\bf u})\leq ir_{\pi}(B)-len({\bf u})$. Let ${\bf u}\in V(B_{i})\setminus V(B_{i-1})$ be a decision node. By the fourth statement of Lemma 26, there is ${\bf u}_{0}\in{\bf B}_{i-1}$ such that $len({\bf u})=len({\bf u}_{0})+1$ and $ir_{\pi,i}({\bf u})\leq ir_{\pi,i-1}({\bf u}_{0})-1$. Hence, by the induction assumption, $ir_{\pi,i}({\bf u})\leq ir_{\pi}(B)-len({\bf u}_{0})-1=ir_{\pi}(B)-len({\bf u})$ as required. ∎
It follows that for each decision node ${\bf u}\in V(B_{q})$, $len({\bf u})\leq ir_{\pi}(B)$. Indeed, assume that this is not the case for some ${\bf u}\in V(B_{q})$. Then, by Claim 13, $ir_{\pi,i}({\bf u})\leq ir_{\pi}(B)-len({\bf u})<0$ in contradiction to the definition of the irregularity index of a node. We conclude that the names of the nodes are uniquely determined by sequences of at most $ir_{\pi}(B)+1$ elements of $B$, (the first coordinate plus a sequence of at most $ir_{\pi}(B)$ conjunction nodes). Clearly, the number of such distinct names is upper-bounded by $n^{ir_{\pi}(B)+1}$. ∎
4.3 Proof of Theorem 20
Let $f_{0}(s,x_{1},x_{2},y_{1},y_{2})$ be a Boolean function such that an assignment $\bf a$} to $\{s,x_{1},x_{2},y_{1},y_{2}\}$ is satisfying if and only if the following holds.
  1. 1.
    If ${\bf a}(s)=0$ then $\bf a$} does not map both $x_{1}$ ad $x_{2}$ to zeroes and $\bf a$} does not map both $y_{1}$ and $y_{2}$ to zeroes.
  2. 2.
    If ${\bf a}(s)=1$ then $\bf a$} does not map both $x_{1}$ ad $x_{2}$ to $1$. and $\bf a$} does not map both $y_{1}$ and $y_{2}$ to $1$.
For each $i\in\{0,1\}$, we denote by $D[x,y,i]$ an obdd querying variables $x$ and $y$ in the order listed. $f(D[x,y,i])$ is true on all the assignments but $\{(x,i),(y,i)\}$. Further on, let $D[s,x_{1},x_{2},y_{1},y_{2}]$ be defined as a $\wedge_{d}$-fbdd whose source $u$ is a variable node labelled with $s$. Let $u_{0}$ and $u_{1}$ be the children of $u$ such that $(u,u_{0})$ is labelled with 0 and $(u,u_{1})$ is labelled with $1$. Each $u_{i}$ is a $\wedge_{d}$ node whose children are $D[x_{1},x_{2},i]$ and $D[y_{1},y_{2},i]$. All the $True$ sinks of the latter two are contracted into a single $True$ sink and all the $False$ sinks are contracted into a single $False$ sink. We observe that
$f_{0}(s,x_{1},x_{2},y_{1},y_{2})=\bigvee_{i\in\{0,1\}}s^{i}\wedge D[x_{1},x_{2},i]\wedge D[y_{1},y_{2},i]$ (1)
where $s^{1}=s$ and $s^{0}=\neg s$. The next lemma follows from (1) and by construction.
Lemma 27.
$f(D[s,x_{1},x_{2},y_{1},y_{2}])=f_{0}(x_{1},x_{2},y_{1},x_{2})$
We now consider a variation $D^{-}[s,x_{1},x_{2},y_{1},y_{2}]$ of $D[s,x_{1},x_{2},y_{1},y_{2}]$ that differs from the former is that $True$ sinks are not all contacted together. In particular, we have the xTruex-True sink obtained by contracting $True$ sinks of graphs $D[x_{1},x_{2},i]$ and the yTruey-True sink obtained by contracting together $True$ sinks of the graphs $D[y_{1},y_{2},i]$.
We now define a $\wedge_{d}$-fbdd $B_{n}$ over variables $s_{1},x_{1,1},x_{1,2},y_{1,1},y_{1,2},\dots s_{n}.x_{n,1},x_{n,2},y_{n,1},y_{n,2}$. The building blocks for the graph are $D^{-}[s_{i},x_{i,1},x_{i,2},y_{i,1},y_{i,2}]$ for each $i\in[n-1]$ and $D[s_{n},x_{n,1},x_{n,2},y_{n,1},y_{n,2}]$. For each $i\in[n-1]$, we identify the xTruex-True sink of $D[s_{i},x_{i,1},x_{i,2},y_{i,1},y_{i,2}]$ with the source of $D[s_{i+1},x_{i+1,1},x_{i+1,2},y_{i+1,1},y_{i+1,2}]$. The remaining $True$ sinks are contacted into a single $True$ sink and the remaining $False$ sinks are contracted into a single $False$ sink.
In order to proceed, we need to upgrade our notation. We denote the set $\{s_{1},\dots,s_{n}\}$ by $S_{n}$ and the set $\{s_{i},\dots,s_{n}\}$ by $S^{i}_{n}$. Further on, we denote the sets $\{x_{1,1},x_{1,2},\dots,x_{n,1},x_{n,2}\}$ and $\{y_{1,1},y_{1,2},\dots,y_{n,1},y_{n,2}\}$ by $X_{n}$ and $Y_{n}$ respectively and the sets $\{x_{i,1},x_{i,2},\dots,x_{n,1},x_{n,2}\}$ and $\{y_{i,1},y_{i,2},\dots,y_{n,1},y_{n,2}\}$ by $X^{i}_{n}$ and $Y^{i}_{n}$ respectively. Next $\wedge_{i\in[n]}f_{0}(s_{i},x_{i,1},x_{i,2},y_{i,1},y_{i,2})$ is denoted by $f_{n}(S_{n},X_{n},Y_{n})$. Finally, $\wedge_{j\in\{i,\dots,n\}}f_{0}(s_{j},x_{j,1},x_{j,2},y_{j,1},y_{j,2})$ is denoted by $f^{i}_{n}(S^{i}_{n},X^{i}_{n},Y^{i}_{n})$.
Lemma 28.
$f(B_{n})=f_{n}(S_{n},X_{n},Y_{n})$.
Proof.
Let $u[1],\dots,u[n]$ be the respective sources of
$D^{-}[s_{1},x_{1,1},x_{1,2},y_{1,1},y_{1,2}],\dots D^{-}[s_{n-1},x_{n-1,1},x_{n-1,2},y_{n-1,1},y_{n-1,2}]$ and $D[s_{n},x_{n,1},x_{n,2},y_{n,1},y_{n,2}]$.
We prove by induction on $i$ going from $n$ down to $1$ that $f([B_{n}]_{u[i]})=f^{i}_{n}(S^{i}_{n},X^{i}_{n},Y^{i}_{n})$. Since $B_{n}]_{u[1]}=B_{n}$, this will imply the lemma. For $i=n$, this is just Lemma 27. So, we assume that $i<n$.
For $a\in\{0,1\}$, let $w_{a}$ be the source of $D[x_{i,1},x_{i,2},a]$. It is immediate from the induction assumption that
$f([B_{n}]_{w_{a}})=f(D[x_{i,1},x_{i,2},a])\wedge f^{i+1}_{n}(S^{i+1},X^{i+1}_{n},Y^{i+1}_{n})$ (2)
Further on, for each $a\in\{0,1\}$, let $v_{a}$ be the child of $u[i]$ such that $(u[i],v_{a})$ is labelled with $a$. It follows by construction and from (2) that
$f([B_{n}]_{v_{a}})=f(D[x_{i,1},x_{i_{2}},a])\wedge f(D[y_{i,1},y_{i,2},a])\wedge f^{i+1}_{n}(S^{i+1}_{n},X^{i+1}_{n},Y^{i+1}_{n})$ (3)
It then follows from (3) and by construction that
$f([B_{n}]_{u[i]}=(\bigvee_{a\in\{0,1\}}s_{i}^{a}\wedge f(D[x_{i,1},x_{i,2},a])\wedge f(D[y_{i,1},y_{i,2},a]))\wedge f^{i+1}_{n}(S^{i+1}_{n},X^{i+1}_{n},Y^{i+1}_{n})$ (4)
Combining (1), (3), and (4) we observe that
$f([B_{n}]_{u[i]}=f_{0}(s_{i},x_{i,1},x_{i,2},y_{i,1},y_{i,2})\wedge f^{i+1}_{n}(S^{i+1}_{n},X^{i+1}_{n},Y^{i+1}_{n})$ (5)
It only remains to observe that $f_{0}(s_{i},x_{i,1},x_{i,2},y_{i,1},y_{i,2})\wedge f^{i+1}_{n}(S^{i+1}_{n},X^{i+1}_{n},Y^{i+1}_{n})=f^{i}_{n}(S^{i}_{n},X^{i}_{n},Y^{i}_{n})$. ∎
In order to proceed, we need an easily verifiable lemma addressing the situation where an obdd is placed on top of $\wedge_{d}$-fbdd.
Lemma 29.
Let $\pi$ be a linear order. Let $B_{1}$ be an obdd with two sinks respecting $\pi$. Further on, let $B_{2}$ be a $\wedge_{d}$-fbdd embeddable into $\pi$. Assume that all of $\mathsf{var}(B_{1})$ occur in $\pi$ before all of $\mathsf{var}(B_{2})$. Let $B$ be a $\wedge_{d}$-fbdd obtained by identifying the positive sink of $B_{1}$ with the source of $B_{2}$ and contracting the negative sinks of $B_{1}$ and $B_{2}$. Then $B$ is embeddable into $\pi$ with $ir_{\pi}(B)=ir_{\pi}(B_{2})$.
Let $\pi_{n}=s_{1},x_{1,1},x_{1,2}\dots s_{n},x_{n,1},x_{n,2},y_{n,1},y_{n,2},\dots y_{1,1},y_{1,2}$ and let $\pi^{i}_{n}=s_{i},x_{i,1},x_{i,2}\dots s_{n},x_{n,1},x_{n,2},y_{n,1},y_{n,2},\dots y_{i,1},y_{i,2}$
Lemma 30.
$B_{n}$ is embeddable in $\pi_{n}$ with irregularity index $n$.
Proof.
Let $u[1],\dots,u[n]$ be as in the proof of Lemma 28. For each $i$ from $n$ down to $1$ we prove that $B_{n}]_{u[i]}$ is embeddable in $\pi_{n}$ with the irregularity index ni+1n-i+1. The lemma will then follow as $B_{n}]_{u[1]}=B_{n}$.
Assume first that $i=n$. For each $a\in\{0,1\}$, let $w_{x,a}$ be the source of $D[x_{n,1},x_{n,2},a]$ and let $w_{y,a}$ be the source of $D[y_{n,1},y_{n,2},a]$. We observe that since $(x_{n,1},x_{n,2})$ and $(y_{n,1},y_{n,2})$ are both intervals of $\pi_{n}$, both $B_{n}]_{w_{x,a}}$ and $B_{n}]_{w_{y,a}}$ are embeddable into $\pi_{n}$ with irregularity index 0 by Lemma 18. Next, let $v_{a}$ be the $\wedge_{d}$-node whose children are $w_{x,a}$ and $w_{y,a}$. Since $(x_{n,1},x_{n,2})$ and $(y_{n,1},y_{n,2})$ are disjoint intervals of $\pi_{n}$, we conclude that $B_{n}]_{v_{a}}$ is embeddable into $\pi_{n}$ with irregularity index $1$. We node that $\mathsf{var}([B_{n}]_{v_{a}})$ form a subinterval of $\pi_{n}$ occurring after $s_{n}$ hence a subinterval of $\pi_{n}^{>s_{n}}$. We conclude that, by construction $B_{n}]_{u[n]}$ is embeddable into $\pi_{n}$ with irregularity index $1$.
Assume now that $i<n$. We observe first that the order of $\mathsf{var}([B_{n}]_{u[i+1]})]$ under $\pi$ is $\pi_{n}^{i+1}$, clearly an interval of $\pi_{n}$. For each $a\in\{0,1\}$, we now reuse the notation for $w_{x,a}$ to the source of $D[x_{i,1},x_{i,2},a]$ and let $w_{y,a}$ be the source of $D[y_{i,1},y_{i,2},a]$. Since $(x_{i,1},x_{i,2})$ occurs in $\pi_{n}$ before $\pi^{i+1}_{n}$ we argue by construction, the induction assumption and Lemma 29 that $B_{n}]_{w_{x,a}}$ is embeddable into $\pi_{n}$ with irregularity index nin-i. Further on, we reuse $v_{a}$ to be the $\wedge_{d}$ node connecting $w_{x,a}$ and $w_{y,a}$ . We observe that since $(y_{i,1},y_{i,2})$ occurs in $\pi_{n}$ after $(x_{i,1},x_{i_{2}})+\pi_{i+1}$ $B_{n}]_{v_{a}}$ is embeddable into $\pi_{n}$ with irregularity index ni+1n-i+1. We further note that $Var([B_{n}]_{v_{a}})$ form an interval of $\pi_{n}$ occurring after $s_{i}$ and hence an interval of $\pi^{>s_{i}}_{n}$. Therefore, by construction $B_{n}]_{u[i]}$ is embeddable into $\pi_{n}$ with the irregularity index ni+1n-i+1 as required. ∎
Lemma 31.
Let $B$ be an obdd representing $f_{n}=f_{n}(S_{n},X_{n},Y_{n})$ and respecting $\pi_{n}$. Then $|B|\geq 2^{n}$.
Proof.
Let $\bf S$} be the set of all the assignments to $S_{n}\cup X_{n}$. that map each $x_{i,1}$ to 0 and each $x_{i,2}$ to $1$.
Claim 14.
For any two ${\bf a}_{1},{\bf a}_{2}\in{\bf S}$, $f_{n}|_{{\bf a}_{1}}\neq f_{n}|_{{\bf a}_{2}}$.
Since $\pi_{n}$ places all the variables in $\bf S$} before all the variables of $\mathsf{var}(f_{n})\setminus{\bf S}$, it follows from the claim and a well known facts about obdds that $|B|\geq|{\bf S}|$. By construction, $|{\bf S}|=2^{n}$ so the lemma follows. It this remains to prove the claim.
Since ${\bf a}_{1}\neq{\bf a}_{2}$, there is $i\in[n]$ such that say ${\bf a}_{1}(s_{i})=0$ whereas ${\bf a}_{1}(s_{1})=1$. It follows that there is no satisfying assignment for $f_{n}|_{{\bf a}_{1}}$ that assigns both $y_{i,1}$ and $y_{i,2}$ with zeroes while there is such an assignment for $f_{n}|_{{\bf a}_{2}}$. ∎
Proof.
(Of Theorem 20) For each $n$ we consider the function $f_{n}$ and the order $\pi_{n}$ of its variables. We note that, by construction, $B_{n}$ is of size linear in $n$. Furthermore, it follows from combination of Lemmas 30 and 31 that $f(B_{n})=f_{n}$ and $B_{n}$ respects $\pi_{n}$ with irregularity index $n$. On the other hand, it follows from Lemma 31 that if $f_{n}$ is represented by an obdd respecting $\pi_{n}$ then such an obdd must be of size at least $2^{n}$. ∎
5 The effect of structuredness on Decision dnnf
dnnf, being a generalization of fbdd, does not enable efficient $Apply$ (conjunction) operation. A restriction on dnnf that enables carrying $Apply$ efficiently is structuredness [16]. The purpose of this section is to define two different forms of structuredness for Decision dnnfs and discuss the power of these models.
Recall that a dnnf is just a deMorgan circuit with all the $\wedge$ gates being decomposable. Assume that both $\wedge$ and $\vee$ gates have fan-in $2$. To define the structuredness, we need to define first a variable tree or a vtree.
Definition 23.
A vtree $VT$ over a set $V$ of variables is a rooted binary tree with every non-leaf node having two children and whose leaves are in a bijective correspondence with $V$. The function $\mathsf{var}$ naturally extends to vtrees. In particular, for each $x\in V(VT)$, $\mathsf{var}(VT_{x})$ is the set of variables labelling the leaves of $VT_{x}$.
Definition 24.
Let $D$ be a dnnf A $\wedge_{d}$-node $u$ of $D$ with inputs $u_{1}$ and $u_{2}$ respects a vtree $VT$ if there is a node $x$ of $VT$ with children $x_{1}$ and $x_{2}$ such that $\mathsf{var}(D_{u_{1}})\subseteq\mathsf{var}(VT_{x_{1}})$ and $\mathsf{var}(D_{u_{2}})\subseteq\mathsf{var}(VT_{x_{2}})$. A dnnf $D$ over $V$ respects $VT$ if each $\wedge_{d}$-node of $D$ respects $VT$. A dnnf $D$ is structured if it respects a vtree.
To apply the structuredness to Decision dnnf, we need to be aware of two equivalent definitions of the model: as a special case of dnnf and as a generalization of fbdd. So far, we have used the latter definition which we referred to as $\wedge_{d}$-fbdd. The former definition states that a Decision dnnf is a dnnf where each $\vee$ node $u$ obeys the following restriction. Let $u_{0}$ and $u_{1}$ be the inputs of $u$. Then they both are $\wedge_{d}$-nodes and there is a variable $x\in\mathsf{var}(D)$ such that one of inputs of $u_{0}$ is an input gate labelled with $\neg x$ and one of the inputs of $u_{1}$ is an input gate labelled with $x$. Then a structured decision dnnf is just a $\wedge_{d}$-fbdd whose representation as a Decision dnnf respects some vtree. As the fpt upper bound of [14], in fact, applies to structured Decision dnnfs, the following holds.
Theorem 32.
A cnf $\varphi$ of primal treewidth at most $k$ can be represented as a structured decision dnnf of size at most $2^{k}\cdot|\varphi|$.
Yet the model is quite restrictive and can be easily ‘fooled’ into being xp-sized even with one long clause. In particular, the following has been established in [5].
Theorem 33.
Let $G$ be a graph without isolated vertices. Let $\varphi^{*}(G)$ be the cnf obtained from $\varphi(G)$ by adding a single clause $\neg V(G)=\{\neg v|v\in V(G)\}$. Then the size of structured Decision dnnf representing $\varphi^{*}(G)$ is at least $\Omega(\textsc{fbdd}(\varphi(G)))$.
In particular, there is an infinite set of positive integers $k$ for each of them there is a class ${\bf G}_{k}$ of graphs of treewidth at most $k$ (and hence the $itw(\varphi^{*}(G))\leq k+1$) and the structured Decision dnnf representation of $\varphi^{*}(G)$ being of size at least. $n^{\Omega(k)}$ where $n=|V(G)|$.
In order to get an additional insight into the ability of $\wedge_{d}$-fbdd to efficiently represent cnfs of bounded incidence treewidth, we introduce an alternative definition of structuredness, this time based on $\wedge_{d}$-fbdd.
Definition 25.
A $\wedge_{d}$-fbdd $B$ respects a vtree $VT$ if each $\wedge_{d}$-node of $B$ respects $VT$. That is, unlike in the case of the ‘proper’ structuredness, we do not impose any restrictions on the decision nodes of $B$. We say that $B$ is structured if $B$ respects a vtree $VT$.
Remark 8.
Note that, even though Decision dnnf and $\wedge_{d}$-fbdd are the same model, their structured versions are not. In particular, a structured $\wedge_{d}$-fbdd is a generalization of fbdd as it does not impose any restriction on a $\wedge_{d}$-fbdd in the absence of $\wedge_{d}$ nodes. As a result the property of efficient conjunction held for structured dnnf does not hold for structured $\wedge_{d}$-fbdds.
We consider structured $\wedge_{d}$-fbdd a very interesting model for further study because it is a powerful special case of $\wedge_{d}$-fbdd which may lead to insight into size lower bounds of $\wedge_{d}$-fbdds for cnf of bounded incidence treewidth. In particular, we do not know whether such cnfs can be represented by fpt-sized structured $\wedge_{d}$-fbdds. We know, however, that the instances that we used so far for deriving xp lower bounds admit an fpt-sized representation as structured $\wedge_{d}$-fbdds. In fact, a much more general result can be established.
Theorem 34.
Let $\varphi$ be a cnf that can be turned into a cnf of primal treewidth at most $k$ by removal of at most $p$ clauses. Then $\varphi$ can be represented as a structured $\wedge_{d}$-fbdd of size at most $O(n^{p}\cdot 2^{k})$ where $n=|\mathsf{var}(\varphi)|$.
In the rest of this section, we provide a proof of Theorem 34.
Let $\bf g$} be an assignment to a subset of variables of a cnf $\psi$. Let $\psi[{\bf g}]$ be the cnf obtained from $\psi$ by removal of all the clauses satisfied by $\bf g$} and removal of the occurrences of $\mathsf{var}({\bf g})$ from the remaining clauses.
We note that $\psi[{\bf g}]$ is not necessarily the same function as $\psi|_{\bf g}$ as the set of the variables of the former may be a strict subset of the set of variables of the latter. Yet, it is not hard to demonstrate the following.
Proposition 4.
If $\mathsf{var}(\psi[{\bf g}])=\mathsf{var}(\psi|_{\bf g})$ then $\mathcal{S}(\psi[{\bf g}])=\mathcal{S}(\psi|_{\bf g})$. Otherwise, $\mathcal{S}(\psi|_{\bf g})=\mathcal{S}(\psi[{\bf g}])\times Y^{\{0,1\}}$ where $Y=\mathsf{var}(\psi|_{\bf g})\setminus\mathsf{var}(\psi[{\bf g}])$.
In order to prove Theorem 34, we identify a set $\bf C$} of at most $p$ ’long’ clauses whose removal from $\varphi$ results in a cnf of primal treewidth at most $k$. An important part of the resulting structured $\wedge_{d}$-fbdd is an $O(n^{p})$-sized decision tree representation of $\bf C$}. We are now going to define a decision tree and to demonstrate the possibility of such a representation.
Definition 26.
A decision tree is a rooted binary tree where each leaf node is labeled by either $True$ or $False$ and each non-leaf node is labelled with a variable. Moreover, for each non-leaf node one outgoing edge is labelled with $1$, the other with 0. A decision tree must be read-once, that is the same variable does not label two nodes along a directed path. Then each directed path $P$ corresponds to the assignment ${\bf a}(P)$ in the same way as we defined it for $\wedge_{d}$-fbdd.
For a decision tree $DT$, $\mathsf{var}(DT)$ is the set of variables labeling the non-leaf nodes of $DT$. We say that $DT$ represents a function $f$ is $\mathsf{var}(DT)\subseteq\mathsf{var}(f)$ and the following two statements hold.
  1. 1.
    Let $P$ be a root-leaf path of $DT$ ending with a node labeled by $True$. Then $f({\bf g})=1$ for each $\bf g$} such that $\mathsf{var}({\bf g})=\mathsf{var}(f)$ and ${\bf a}(P)\subseteq{\bf g}$.
  2. 2.
    Let $P$ be a root-leaf path of $DT$ ending with a node labeled by $False$. Then $f({\bf g})=0$ for each $\bf g$} such that $\mathsf{var}({\bf g})=\mathsf{var}(f)$ and ${\bf a}(P)\subseteq{\bf g}$.
Lemma 35.
A cnf of $p$ clauses can be represented as a decision tree $DT(\varphi)$ of size $O(n^{p})$ where $n=|\mathsf{var}(\varphi)|$.
Proof.
If $\varphi$ has no clauses then the resulting decision tree consists of a single truetrue node. If $\varphi$ has an empty clause then the decision tree consists of a single falsefalse node.
Otherwise, pick a clause $C$ with variables $x_{1},\dots,x_{r}$. We define $DT(\varphi)$ as having a root-leaf path $P$ consisting of vertices $u_{1},\dots,u_{r+1}$ where $u_{1},\dots,u_{r}$ are respectively labelled with $x_{1},\dots,x_{r}$. Furthermore for each $1\leq i\leq r$, the outgoing edge of $u_{i}$ is labelled with 0 if $x_{i}\in C$ and with $1$ if $\neg x_{i}\in C$. Clearly, ${\bf a}(P)$ falsifies $C$, hence we label $u_{r+1}$ with $False$.
Next, for each $1\leq i\leq r$, we introduce a new neighbour $w_{i}$ of $u_{i}$ and, of course, label the edge $(u_{i},w_{i})$ with the assignment of $x_{i}$ as in $C$. Let $P_{i}$ be the path from the root to $w_{i}$ and let ${\bf g}_{i}={\bf g}(P_{i})$ Clearly, ${\bf g}_{i}$ satisfies $C$. Therefore $\varphi[{{\bf g}_{i}}]$ can be expressed as a cnf of at most p1p-1 clauses. Make each $w_{i}$ the root of $DT(\varphi[{{\bf g}_{i}}])$. Clearly, the resulting tree represents $\varphi$. By the induction assumption, the trees $DT(\varphi[{{\bf g}_{i}}])$ are of size at most $n^{p-1}$ and there are at most $|C|\leq n$ such trees. Hence the lemma follows. ∎
Proof.
(of Theorem 34) Let $\bf C$} be at most $p$ clauses of $\varphi$ whose removal makes the remaining cnf to be of primal treewidth at most $k$.
Let $DT({\bf C})$ be the decision tree of size $O(n^{p})$ representing $\bf C$} guaranteed to exist by Lemma 35. Let $\psi=\varphi\setminus{\bf C}$. Since structured Decision dnnf is a special case of structured $\wedge_{d}$-fbdd, Theorem 32 holds for structured $\wedge_{d}$-dnnf. In fact, it is not hard to establish the following strengthening.
Claim 15.
There is a vtree $VT$ such that the following holds. Let $P$ be a root-leaf path of $DT({\bf C})$. Then there is a structured $\wedge_{d}$-fbdd $B_{P}$ respecting $VT$ and representing $\psi[{\bf a}(P)]$ and of size $O(2^{k}\cdot|var(\psi)|)$.
We construct a structured $\wedge_{d}$-fbdd $B$ representing $\varphi$ of size at most $O(n^{p}2^{k}\cdot|\mathsf{var}(\varphi))$ in the following three stages.
  1. 1.
    For each root-leaf path $P$ of $DT({\bf C})$ that ends at a $True$ sink, remove the $True$ label off the sink and identify the sink with the source of $B_{P}$ as per Claim 15.
  2. 2.
    Contract all the $True$ sinks of the resulting dag into a single $True$ sink and all the $False$ sinks into a single $False$ sink.
  3. 3.
    Let $B^{\prime}$ be the dag obtained as a result of the above two stages. Let $X=\mathsf{var}(\varphi)\setminus\mathsf{var}(B^{\prime})$. If $X=\emptyset$ we set $B=B^{\prime}$. Otherwise, let $x_{1},\dots,x_{a}$ be the variables of $X$. Then $B$ is obtained from $B^{\prime}$ by introducing new nodes $u_{1},\dots,u_{a}$ and setting $u_{a+1}$ to be the source of $B^{\prime}$. For each $1\leq i\leq a$, label $u_{i}$ with $x_{i}$ and introduce two edges from $u_{i}$ to $u_{i+1}$, one labelled with $1$, the oth er labelled with 0.
It is not hard to see that $B$ is indeed a $\wedge_{d}$-fbdd representing $\varphi$ and with the required upper bound on its size. We further note that each $\wedge_{d}$ node of $B$ is a $\wedge_{d}$ node of some $B_{P}$ and hence respects $VT$ by construction.
6 Conclusion
The main purpose of this paper is obtaining better understanding of the power of $\wedge_{d}$-fbdd in terms of the representation of cnfs of bounded incidence treewidth. We believe that the next important step forward is understanding the complexity of structured $\wedge_{d}$-fbdd representing cnfs of bounded incidence treewidth. Recall that Theorem 34 present an upper bound that is fpt in $k$ but xp in $p$. A natural question for further consideration is whether we can improve the upper bound so that it is fpt. We conjecture that this is not the case, a formal statement of the conjecture is provided below.
Conjecture 1.
There are constants $c$ and $d$ such that the following holds. For each sufficiently large $k$, there is an infinite class ${\bf\Phi}_{k}$ of cnfs $\varphi$ so that $\varphi$ can be turned into a cnf of primal treewidth $ck$ by removal of $k$ clauses and such that a representation of $\varphi$ as a structured $\wedge_{d}$-fbdd is of size at least $n^{d\cdot k}$ where $n=|Var(\varphi)|$.
We believe a recent result [21] will be useful for resolving Conjecture 1. In [21], we consider $\wedge_{d}$-fbdd with imbalanced gates. Roughly speaking, each $\wedge_{d}$-gate has only one input depending on more than $n^{\alpha}$ inputs for some constant $0<\alpha<1$ fixed for a specific $\wedge_{d}$-fbdd. We establish a lower bound $n^{\Omega(1-\alpha)\cdot k}$ for cnf of primal treewidth. Intuitively, this means that Conjecture 1 can be considered only for balanced vtrees. In particular, we believe that if an xp lower bound is obtained for balanced vtrees the lower bound can be combined with the lower bound in [21] to form a ‘win-win’ argument.
We believe that the xp lower bound for $\wedge_{d}$-obdd has potential to be ‘transferred to the area of proof complexity. Indeed, a similar ‘primal vs incidence treewidth’ dilemma is known for the resolution proof system: it is easy to demonstrate existence of fpt sized resolution refutation parameterized by the primal treewidth of the input cnf. However, the existence of fpt sized resolution refutation parameterized by the incidence treewidth is a well known open question. Importantly, the same dilemma persists for the regular resolution.
It is well known that a regular resolution refutation of a cnf $\varphi$ can be represented as a read-once branching program whose sinks are labelled with clauses of $\varphi$. The constraint is that the assignment carried by each source-sink path must falsify the clause associated with the sink. Let us say that the a regular resolution refutation is oblivious if the corresponding read-once branching program respects a linear order of its variables. It is known that although being similar to fbdd, regular resolution has a greater intuitive similarity to $\wedge_{d}$-fbdd: in order to refute a conjunction of two variable-disjoint cnfs, it is enough to refute just one of them. Continuing this similarity to the oblivious case, we observe an intuitive similarity between $\wedge_{d}$-obdd and oblivious regular resolution. The xp lower bound for the former naturally leads to an open question whether a similar lower bound exists for the latter.
Open question 2.
What is the complexity of oblivious resolution refutation parmeterized by the incidence treewidth of the input cnf? In particular, can an xp lower bound be established?
In the rest of the section, we will consider possible directions of further research stemming from our results related to the Apply operation. Recall, that we demonstrated a possibility of efficient Apply operation for two $\wedge_{d}$-obdd embeddable into the same order if they are both ‘close’ to an obdd respecting the same order. The obdd can be seen as a ‘common denominator’ of two $\wedge_{d}$-obdds. A natural direction is to consider closeness of two models to each other rather than to some third model. This leads us to stating the following ‘open-ended’ open question.
Open question 3.
Introduce a parameter measuring closeness of two $\wedge_{d}$-obdds $B_{1}$ and $B_{2}$ respecting the same order but not necessarily embeddable into it so that $f(B_{1})\wedge f(B_{2})$ can be represented as a $\wedge_{d}$-obdd of size $|B_{1}|\times|B_{2}|$ to the power proportional to the value of the parameter.
Question 3 naturally leads to a similar question being asked about structured dnnf.
Open question 4.
Introduce a parameter $k=k(T_{1},T_{2})$ measuring closeness of two vtrees $T_{1}$ and $T_{2}$ so that if $B_{1}$ and $B_{2}$ are two structured dnnfs respecting $T_{1}$ and $T_{2}$ respectively then $f(B_{1})\wedge f(B_{2})$ can be represented as a structured dnnf of size $(|B_{1}|\cdot|B_{2}|)^{\Theta(k)}$.
It is known that the notion of embeddability, introduced in this paper in the context of $\wedge_{d}$-obdds is applicable to dnnfs and even to Boolean circuits in general [1]. This leads us to the following open question consisting of two parts.
Open question 5.
  1. 1.
    Introduce the notion of the irregularity index of a dnnf embeddable into a specific linear order.
  2. 2.
    A nondeterministic read-once branching program is oblivious if there is a linear order of variables, so that along each source-sink path, the variables are queried according to this order. We refer to such a branching program as a Nondeterministic obdd (nobdd) (the model is discussed in Section 10.1 of [25]). Let $B$ be a dnnf embeddable into an order $\pi$ with irregularity index $k$. Similarly to Theorem 19, can $B$ be simulated by a nobdd of size $|B|^{O(k)}$?
It is known that the embeddability can be seen as a weaker form of structuredness [1]. The goal of the next conjecture is to make this connection ‘quantified’.
Conjecture 2.
A structured dnnf is also a dnnf embeddable into a specific order with the irregularity index at most $\log n$.
Specifically, we believe that, for a vtree $T$ a linear order $\pi$ of variables and the irregularity index $k$ of $T$ w.r.t. this order can be naturally defined so that $k\leq\log n$ where $n$ is the number of variables of $T$. Furthermore, we believe that if $B$ is a structured dnnf respecting $T$ then $B$ is embeddable into $\pi$ with irregularity index at most $k$.
Conjecture 2 needs to be accompanied with demonstration of succinctness of embeddable dnnfs as compared to structured ones. This connection is formally specified in the next two open questions. The first question seeks to establish succinctness of embeddable dnnfs as compared to structured ones. The second question is a refinement of the first one subject to a specific irregularity index.
Open question 6.
Is there an infinite set $\bf F$} of Boolean functions so that each $f\in{\bf F}$ can be represented by an embeddable dnnf of size polynomial in $|\mathsf{var}(f)|$ but requires a structured dnnf of size exponential in $|\mathsf{var}(f)|$ for its representation?
Open question 7.
Does the following hold for each sufficiently large $k$. There is an infinite class ${\bf F}_{k}$ of Boolean functions so that each $f\in{\bf F}_{k}$ is representable as an embeddable dnnf with irregularity index $k$ and of size polynimial in $|\mathsf{var}(f)|$ but requires for its representation an exponential sized structured dnnf respecting a vtree with irregularity index at most $k$.
The final open question that we propose in this paper facilitates the primary purpose of the embeddability notion, namely efficient implementation of Apply. The open question below can be seen as an upgrade of Open Question 4.
Open question 8.
Introduce a parameter $k=k(B_{1},B_{2})$ measuring closeness of two embeddable dnnfs $B_{1}$ and $B_{2}$ so that $f(B_{1})\wedge f(B_{2})$ can be represented as an embeddable dnnf of size at most $(|B_{1}|\cdot|B_{2}|)^{\Theta(k)}$. In case $B_{1}$ and $B_{2}$ are structured respecting trees $T_{1}$ and $T_{2}$, respectively, the parameter should be equal $k(T_{1},T_{2})$ as in Open Question 4.
Acknowledgement
We would like to thank Florent Capelli for useful feedback.
Appendix A Proof of Lemma 12
Claim 16.
Let $X=(\mathsf{var}(B)\setminus\{x\})\setminus\mathsf{var}(B^{\prime})$. Then $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B^{\prime})\times X^{\{0,1\}}$.
It is immediate from the essentiality assumption that $X=\emptyset$. Therefore, the statement of the lemma is immediate from the claim. It thus remains to prove the claim. The proof is by induction on $|B|$. For $|B|=1$, $\mathsf{var}(B)=\emptyset$, hence the claim is true in a vacuous way.
For $|B|>1$, we assume first that the source $u$ is a decision node labelled by a variable $y$. Further on, assume first that $y=x$. Let $(u,v)$ be the outgoing edge of $u$ labelled with $i$. It is clear by construction that $B^{\prime}=B_{v}$. But then the claim is immediate from Lemma 36. We thus assume that $y\neq x$.
Let ${\bf h}\in\mathcal{S}(B)|_{\bf g}$. Let ${\bf h}_{0}=Proj({\bf h},\mathsf{var}(B^{\prime}))$. We need to demonstrate that ${\bf h}_{0}\in\mathcal{S}(B^{\prime})$.
Let ${\bf a}={\bf h}\cup\{(x,i)\}$. Let $j$ be the assignment of $y$ by $\bf a$}. Let ${\bf b}={\bf a}\setminus\{(y,j)\}$.
Let $(u,v)$ be the outgoing edge of $u$ labelled by $j$. Let ${\bf b}_{0}=Proj({\bf b},\mathsf{var}(B_{v}))$. By Lemma 36, ${\bf b}_{0}\in\mathcal{S}(B_{v})$. Let ${\bf c}_{0}={\bf b}_{0}\setminus\{(x,i)\}$. Let ${\bf c}_{1}=Proj({\bf c}_{0},\mathsf{var}(B^{\prime}_{v}))$. By the induction assumption, ${\bf c}_{1}\in\mathcal{S}(B^{\prime}_{v})$.
Let ${\bf h}_{1}={\bf c}_{1}\cup\{(y,j)\}$. On the one hand ${\bf h}_{1}\subseteq{\bf h}$. On the other hand $\mathsf{var}(B^{\prime}_{v})\subseteq\mathsf{var}(B^{\prime})$ simply by construction. We conclude that ${\bf h}_{1}\subseteq{\bf h}_{0}$. By Lemma 36 applied to $B^{\prime}$ and $B^{\prime}_{v}$, we conclude that an arbitrary extension of ${\bf h}_{1}$ to $\mathsf{var}(B^{\prime})$ is an element of $\mathcal{B^{\prime}}$. It follows that ${\bf h}_{0}\in\mathcal{S}(B^{\prime})$.
For the opposite direction, pick $j\in\{0,1\}$ Let $v$ be the outgoing edge of $u$ in $B^{\prime}$ such that $(u,v)$ is labelled with $j$. Let ${\bf q}=\{(x,i),(y,j)\}$. Let $Y=(\mathsf{var}(B)\setminus\{x,y\})\setminus\mathsf{var}(B^{\prime}_{v})$. Let ${\bf c}\in\mathcal{S}(B^{\prime}_{v})$ and let ${\bf y}\in Y^{\{0,1\}}$. It is sufficient to prove that ${\bf c}\cup{\bf y}\in\mathcal{S}(B)|_{\bf q}$.
The set $Y$ can be weakly partitioned (meaning that some of partition classes may be empty) into $Y_{0}=(\mathsf{var}(B_{v})\setminus\mathsf{var}({\bf g}))\setminus\mathsf{var}(B^{\prime}_{v})$ and $Y_{1}=(\mathsf{var}(B)\setminus\{y\})\setminus var(B_{v})$. Let ${\bf y}_{0}=Proj({\bf y},Y_{0})$ and ${\bf y}_{1}=Proj({\bf y},Y_{1})$. By the induction assumption, ${\bf c}\cup{\bf y}_{0}\in\mathcal{S}(B_{v})|_{\bf g}$. Let ${\bf c}_{0}={\bf c}\cup{\bf y}_{0}\cup{\bf g}$. By Lemma 36, ${\bf c_{0}}\cup{\bf y}_{1}\in\mathcal{S}(B)|_{\{(y,j)\}}$. We note that ${\bf c}_{0}\cup{\bf y}_{1}={\bf c}\cup{\bf g}$. Hence ${\bf c}\in\mathcal{S}(B)|_{\bf q}$ as required.
It remains to assume that $u$ is a conjunction node. Let $u_{1}$ and $u_{2}$ be the children of $u$. For the sake of brevity, we denote $B_{u_{1}}$ and $B_{u_{2}}$ by $B_{1}$ and $B_{2}$ respectively. Assume w.l.o.g. that $x\in\mathsf{var}(B_{1})$. By Lemma 37
$\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B_{1})|_{\bf g}\times\mathcal{S}(B_{2})$ (6)
By applying the induction assumption to $B_{1}$, (6) is transformed into
$\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}\times\mathcal{S}(B^{\prime}_{1})\times\mathcal{S}(B_{2})$ (7)
By construction, $B^{\prime}$ is obtained by transforming $B_{1}$ int $B^{\prime}_{1}$. Therefore, by Lemma 37 used with the empty assignment
$\mathcal{S}(B^{\prime})=\mathcal{S}(B^{\prime}_{1})\times\mathcal{S}(B_{2})$ (8)
The claim now follows from combination of (7) and (8).
The proof of the lemma is now complete.
Appendix B Proof of Lemma 3
Lemma 36.
Let $B$ be a $\wedge_{d}$-obdd. Let $\bf g$} be an assignment. Suppose that the source node $u$ of $B$ is a decision node associated with a variable $x$. Let $i\in\{1,2\}$ be such that $(x,i)\in{\bf g}$. Let $(u,v)$ be the outgoing edge of $u$ labelled with $i$. Let ${\bf g}_{0}={\bf g}\setminus\{(x,i)\}$. Let $X=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\mathsf{var}(B_{v})$. Then $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B_{v})|_{{\bf g}_{0}}\times X^{\{0,1\}}$
Proof.
Let ${\bf h}\in\mathcal{S}(B)|_{\bf g}$. Let ${\bf a}={\bf g}\cup{\bf h}$. Then ${\bf a}\in\mathcal{S}(B)$.
By construction, $(x,i)\in{\bf a}$. Let ${\bf a}_{0}={\bf a}\setminus\{(x,i)\}$. Clearly, $\mathsf{var}({\bf b}_{0})\subseteq\mathsf{var}(B_{v})\subseteq\mathsf{var}({\bf a}_{0})$. Let ${\bf c}_{0}=Proj({\bf a}_{0},\mathsf{var}(B_{v}))$. We conclude that ${\bf c}_{0}\in\mathcal{S}(B_{v})$. Let ${\bf d}={\bf c}_{0}\setminus{\bf g}_{0}$. It follows that ${\bf d}\in\mathcal{S}(B_{v})|_{{\bf g}_{0}}$.
We note that ${\bf h}={\bf a}_{0}\setminus{\bf g}_{0}$ and hence ${\bf d}\subseteq{\bf h}$. Let ${\bf x}={\bf h}\setminus{\bf d}$. Then $\mathsf{var}({\bf x})=(\mathsf{var}(B)\setminus\mathsf{var}(g))\setminus(\mathsf{var}(B_{v})\setminus\mathsf{var}({\bf g}_{0})=(\mathsf{var}(B)\setminus\mathsf{var}(g))\setminus\mathsf{var}(B_{v})=X$. We conclude that ${\bf h}\in\mathcal{S}(B_{v})|_{{\bf g}_{0}}\times X^{\{0,1\}}$.
Conversely, let ${\bf d}\in\mathcal{S}(B_{v})|_{{\bf g}_{0}}$ and let ${\bf x}\in X^{\{0,1\}}$. Let ${\bf h}={\bf d}\cup{\bf x}$ and let ${\bf a}={\bf g}\cup{\bf h}$. We need to prove that ${\bf a}\in\mathcal{S}(B)$. Let ${\bf g}^{\prime}_{0}=Proj({\bf g}_{0},\mathsf{var}(B_{v}))$. Let ${\bf c}_{0}={\bf g}^{\prime}_{0}\cup{\bf d}$. Then ${\bf c}_{0}\in\mathcal{S}(B_{v})$. Consequently, there is ${\bf b}_{0}\subseteq{\bf c}_{0}$ such that ${\bf b}_{0}\in\mathcal{A}(B_{v})$. Let ${\bf b}={\bf b}_{0}\cup\{(x,i)\}$. By construction, ${\bf b}\in\mathcal{A}(B)$. As ${\bf b}\subseteq{\bf c}_{0}\cup\{(x,i)\}\subseteq{\bf a}$, we conclude that ${\bf a}\in\mathcal{S}(B)$. ∎
Lemma 37.
Let $B$ be a $\wedge_{d}$-obdd. Let $\bf g$} be an assignment. Assume that the source $u$ of $B$ is a conjunction node. Let $u_{1}$ and $u_{2}$ be children of $u$. Let $V=\mathsf{var}(B)$, $V_{1}=\mathsf{var}(B_{u_{1}})$ and $V_{2}=\mathsf{var}(B_{u_{2}})$. Further on, let ${\bf g}_{1}=Proj({\bf g},V_{1})$, ${\bf g}_{2}=Proj({\bf g},V_{2})$. Then $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B_{u_{1}})|_{\bf g_{1}}\times\mathcal{S}(B_{u_{2}})|_{\bf g_{2}}$.
Proof.
Let ${\bf h}\in\mathcal{S}(B)|_{\bf g}$. Let ${\bf a}={\bf g}\cup{\bf h}$. It follows that there is ${\bf b}\subseteq{\bf a}$ such that ${\bf b}\in\mathcal{A}(B)$.
For each $i\in\{1,2\}$, let ${\bf a}_{i}=Proj({\bf a},V_{i})$ and let ${\bf b}_{i}=Proj({\bf b},V_{i})$. By construction, for each $i\in\{1,2\}$, ${\bf b}_{i}\in\mathcal{A}(B_{u_{i}})$ and hence ${\bf a}_{i}\in\mathcal{S}(B_{u_{i}})$. For each ${\bf a}_{i}$, let ${\bf h}_{i}={\bf a}_{i}\setminus{\bf g}_{i}$. Clearly, ${\bf h}={\bf h}_{1}\cup{\bf h}_{2}$ and ${\bf h}_{i}\in\mathcal{S}(B_{u_{i}})|_{{\bf g}_{i}}$ for each $i\in\{1,2\}$. Hence, ${\bf h}\in\mathcal{S}(B_{u_{1}})|_{\bf g_{1}}\times\mathcal{S}(B_{u_{2}})|_{\bf g_{2}}$.
Conversely, let ${\bf h}_{i}\in\mathcal{S}(B_{u_{i}})|_{{\bf g}_{i}}$ for each $i\in\{1,2\}$. Let ${\bf h}={\bf h}_{1}\cup{\bf h}_{2}$ and let ${\bf a}={\bf g}\cup{\bf h}$. We need to demonstrate that ${\bf a}\in\mathcal{S}(B)$. For each $i\in\{1,2\}$, let ${\bf a_{i}}={\bf g}_{i}\cup{\bf h}_{i}$. By assumption, ${\bf a}_{i}\in\mathcal{S}(B_{u_{i}})$ for each $i\in\{1,2\}$ implying existence of ${\bf b}_{1}\subseteq{\bf a}_{1}$ and ${\bf b}_{2}\subseteq{\bf a}_{2}$ such that ${\bf b}_{1}\in\mathcal{A}(B_{u_{1}})$ and ${\bf b}_{2}\in\mathcal{A}(B_{u_{2}})$. Let ${\bf b}={\bf b}_{1}\cup{\bf b}_{2}$. By construction, ${\bf b}\in\mathcal{A}(B)$. As clearly, ${\bf b}\subseteq{\bf a}$, we conclude that ${\bf a}\in\mathcal{S}(B)$. ∎
Lemma 38.
Let $B$ be a $\wedge_{d}$-obdd respecting an order $\pi$. Let $\bf g$} be an assignment to a prefix of $\pi$. Suppose that the source node $u$ of $B$ is a decision node associated with a variable $x$. Let $i\in\{0,1\}$ be such that $(x,i)\in{\bf g}$ Let $(u,v)$ be the outgoing edge of $u$ labelled with $i$. Let ${\bf g}_{0}={\bf g}\setminus\{(x,i)\}$. Then $B_{v}$ respects $\pi_{0}=\pi\setminus\{x\}$, ${\bf g}_{0}$ is an assignment to a prefix of $\pi_{0}$ and $B[{\bf g}]$ is obtained from $B_{v}[{\bf g_{0}}]$ by adding the edge $(u,v)$ labelled with $i$
Proof.
$B_{v}$ obeys $\pi$ but does not contain $x$ hence clearly $B_{v}$ respects $\pi_{0}$. Clearly, removal of $x$ from a prefix of $\pi$ results in a prefix of $\pi_{0}$. Therefore, ${\bf g}_{0}$ assigns a prefix of $\pi_{0}$.
It is not hard to see that $B[{\bf g}]$ can be obtained as follows.
  1. 1.
    Remove the outgoing edge of $u$ labelled with 2i2-i.
  2. 2.
    Remove all the nodes besides $u$ and those in $B_{v}$.
  3. 3.
    Inside $B_{v}$ remove all the edges whose tails are labelled with variables of $\bf g$} but the labels on the edges are opposite to their respective assignments in $\bf g$}. Then remove all the nodes all the nodes of $B_{v}$ that are not reachable from $v$ in the resulting graph.
It is not hard to see that the graph resulting from the above algorithm is the same as when we first compute $B_{v}[{\bf g}_{0}]$ and then add $(u,v)$. ∎
Corollary 1.
With the notation and premises as of Lemma 38, $L_{B}({\bf g})=L_{B_{v}}({\bf g}_{0})$
Proof.
By construction of $B[{\bf g}]$ as specified in Lemma 38, it is immediate that $B[{\bf g}]$ and $B_{v}[{\bf g}_{0}]$ have the same set of complete decision nodes. Moreover, as the construction does not introduce new paths between the complete decision nodes nor removes an existing one, minimal complete decision node of $B_{v}[{\bf g_{0}}]$ remains minimal in $B[{\bf g}]$ and no new minimal complete decision node appears in $B[{\bf g}]$. ∎
We now introduce special notation for the set $X$ as in Lemma 3. In particular, we let $X_{B}({\bf g})=(\mathsf{var}(B)\setminus\mathsf{var}(g))\setminus\bigcup_{w\in L_{B}({\bf g})}\mathsf{var}(B_{w})$.
Corollary 2.
With the notation and premises as of Lemma 38, $X_{B}({\bf g})=((\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\mathsf{var}(B_{v}))\cup X_{B_{v}}({\bf g}_{0})$.
Proof.
By Corollary 1
$X_{B}({\bf g})=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\bigcup_{w\in L_{B_{v}}({\bf g}_{0})}\mathsf{var}(B_{w})$ (9)
As $\mathsf{var}(B_{v})\setminus\mathsf{var}({\bf g}_{0})\subseteq\mathsf{var}(B)\setminus\mathsf{var}({\bf g})$, we conclude that $X_{B_{v}}({\bf g}_{0})\subseteq X_{B}({\bf g})$. But what are $X_{B}({\bf g})\setminus X_{B_{v}}({\bf g}_{0})$? They are all the variables of $\mathsf{var}(B)\setminus\mathsf{var}(B_{v})$ but $\mathsf{var}({\bf g})$. Hence the corollary follows.
Lemma 39.
Let $B$ be a $\wedge_{d}$-obdd respecting an order $\pi$. Let $\bf g$} be an assignment to a prefix of $\pi$. Suppose that the source $u$ of $B$ is a conjunction node with children $u_{1}$ and $u_{2}$. Let $V=\mathsf{var}(B)$, $V_{1}=\mathsf{var}(B_{u_{1}})$ and $V_{2}=\mathsf{var}(B_{u_{2}})$. Further on, let $\pi_{1}=\pi[V_{1}]$, $\pi_{2}=\pi[V_{2}]$ (where $\pi[V_{i}]$ is the order of $V_{i}$ with the elements ordered as in $\pi$), ${\bf g}_{1}=Proj({\bf g},V_{1})$, ${\bf g}_{2}=Proj({\bf g},V_{2})$. Then, for each $i\in\{1.2\}$, $B_{u_{i}}$ is a $\wedge_{d}$-obdd respecting $\pi_{i}$ and ${\bf g}_{i}$ s an assignment over a prefix of $\pi_{i}$. Furthermore, $B[{\bf g}]$ is obtained from $B_{u_{1}}[{\bf g}_{1}]$ and $B_{u_{2}}[{\bf g}_{2}]$ by adding a conjunction node $u$ with children $u_{1}$ and $u_{2}$.
Proof.
For each $i\in\{1,2\}$, $B_{u_{i}}$ respects $\pi$. Hence, $B_{u_{i}}$ obeys the suborder of $\pi$ induced by $\mathsf{var}(B_{u_{i}})$. Assume that say $\mathsf{var}({\bf g}_{1})$ do not form a prefix of $\pi_{1}$. Then there is $x\in\mathsf{var}({\bf g}_{1})$ and $y\in\pi_{1}\setminus\mathsf{var}({\bf g}_{1})$ such that $y<_{\pi_{1}}x$. Then $y<_{\pi}x$ and $y\in\pi\setminus\mathsf{var}({\bf g})$ in contradiction to $\mathsf{var}({\bf g})$ being a prefix of $\pi$.
It is not hard to see that, in the considered case, the second stage of obtaining the alignment can be reformulated as removal of nodes reachable from both $u_{1}$ and $u_{2}$. Now, let us obtain $B_{u_{1}}[{\bf g}_{1}]\cup B_{u_{2}}[{\bf g}_{2}]$ as follows.
  1. 1.
    Carry out the edge removal in both $B_{u_{1}}$ and $B_{u_{1}}$. Clearly, this is equivalent to edges removal from $B$ to obtain $B[{\bf g}]$.
  2. 2.
    Carry out node removal separately for $B_{u_{1}}[{\bf g_{1}}]$ and $B_{u_{2}}[{\bf g_{2}}]$ and then perform the graph union operation. Clearly, the nodes removed (apart from the source) will be precisely those that are not reachable from $u_{1}$ in $B_{u_{1}}$ and in $u_{2}$ in $B_{u_{2}}$ after the removal of edges.
Then, adding the source as specified will result in $B[{\bf g}]$ as required. ∎
Corollary 3.
With the notation and premises as in Lemma 39, $L_{B}({\bf g})=L_{B_{1}}({\bf g}_{1})\cup L_{B_{2}}({\bf g}_{2})$.
Proof.
It is clear from the construction of $B[{\bf g}]$ by Lemma 39 that the set of complete decision nodes of $B[{\bf g}]$ is the union of the sets of such nodes of $B_{1}[{\bf g}_{1}]$ and $B_{2}[{\bf g}_{2}]$. Let $w\in L_{B}({\bf g})$. Assume w.l.o.g. that $w$ is a node of $B_{1}[{\bf g}_{1}]$ as the minimality cannot be destroyed by moving from a graph to a subgraph, we conclude that $w\in L_{B_{1}}({\bf g}_{1})$.
Conversely, let $w\in L_{B_{1}}({\bf g}_{1})\cup L_{B_{2}}({\bf g}_{2})$. We assume w.l.o.g. that $w\in L_{B_{1}}({\bf g}_{1})$ We note that the construction of $B[{\bf g}]$ by Lemma 39 does not add new edges between vertices of $B_{1}[{\bf g}_{1}]$ and a path from a decision node of $B_{2}[{\bf g}_{2}]$ to a decision node of $B_{1}[{\bf g}_{1}]$ is impossible due to the decomposability of $u$. Therefore, we conclude that $w\in L_{B}({\bf g})$. ∎
Corollary 4.
With the notation and premises as in Lemma 39, $X_{B}({\bf g})=X_{B_{1}}({\bf g}_{1})\cup X_{B_{2}}({\bf g}_{2})$
Proof.
Due to the decomposability of $u$, $X_{B}({\bf g})$ is the disjoint union of $X_{1}=X_{B}({\bf g})\cap\mathsf{var}(B_{1})$ and $X_{2}=X_{B}({\bf g})\cap\mathsf{var}(B_{2})$. We demonstrate that for each $i\in\{1,2\}$ $X_{i}=X_{B_{i}}({\bf g}_{i})$. The argument is symmetric for both possible values of $i$, so we assume w.l.o.g that $i=1$.
By definition $X_{1}=Y_{0}\setminus Y_{1}$ where $Y_{0}=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\cap\mathsf{var}(B_{1})=\mathsf{var}(B_{1})\setminus\mathsf{var}({\bf g}_{1})$ and $Y_{1}=(\bigcup_{w\in L_{B}({\bf g})}\mathsf{var}(B_{w}))\cap\mathsf{var}(B_{1})$ By decomposability of $u$ and Corollary 3, $Y_{1}=\bigcup_{w\in L_{B_{1}}({\bf g}_{1})}\mathsf{var}(B_{w})$. Substituting the obtained expressions fr $Y_{0}$ and $Y_{1}$ into $Y_{0}\setminus Y_{1}$, we obtain the definition of $X_{B_{1}}({\bf g}_{1})$. ∎
Proof.
(of Lemma 3) By induction of $|\mathsf{var}(B)|$. Suppose that $|\mathsf{var}(B)|=0$. Then $B$ consists of a singe sink node. The sink cannot be a $False$ node because otherwise, we get a contradiction with $|\mathcal{S}(B)|_{\bf g}\neq\emptyset$. Thus the sink is a $True$ node and hence $\mathcal{S}(B)=\{\emptyset\}$, any alignment of $B$ is $B$ itself and for any assignment $\bf g$}, $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B)$. Clearly, $X_{B}({\bf g})=\emptyset$ and $\mathcal{S}(B)|_{\bf g}=\emptyset^{0,1}$ as required by the lemma.
Assume now that $\mathsf{var}(B)|>0$ and $u$ be the source node of $B$.
Assume that $u$ is a decision node. Let $x$ be the variable associated with $u$. If $\mathsf{var}({\bf g})\cap\mathsf{var}(B)=\emptyset$ then $B[{\bf g}]=B=B_{u}$, that is $\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B_{u})$. As $u$ is the only minimal decision node of $B$, the lemma holds in this case.
Otherwise, we observe that $x\in\mathsf{var}({\bf g})$. Indeed, let $y\in\mathsf{var}({\bf g})\cap\mathsf{var}(B)$. This means that $B$ has a decision node $w$ associated with a variable $y$. If $w=u$ then $y=x$ and we are done. Otherwise, as $u$ is the source node $B$ has a path from $u$ to $w$. It follows that $x$ precedes $y$ in $\pi$. As $\pi_{0}$ is a prefix including $y$, it must also include $x$.
Let $(x,i)$ be an element of $\bf g$}. Let $(u,v)$ be the outgoing edge of $u$ labelled with $i$. Let ${\bf g}_{0}={\bf g}\setminus\{x,i\}$. Let $Y=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\mathsf{var}(B_{v})$. We note that by Lemma 36, $\mathcal{S}(B)|_{\bf g}\neq\emptyset$ implies that $\mathcal{S}(B_{v})({\bf g}_{0})\neq\emptyset$. Therefore, we may apply the induction assumption to $B_{v}[{\bf g}_{0}]$.
Assume that $L_{B_{v}}({\bf g}_{0})\neq\emptyset$. Combining Lemma 36 and the induction assumption, we observe that
$\mathcal{S}(B)|_{\bf g}=\prod_{w\in L_{B_{v}}({\bf g}_{0})}\mathcal{S}(B_{w})\times(X_{B_{v}}({\bf g}_{0})\cup Y)^{0,1}$ (10)
The lemma is immediate from application of Corollaries 1 and 2 to (10).
If $L_{B_{v}}({\bf g}_{0})=\emptyset$ then the same reasoning applies with $\mathcal{S}(B)|_{\bf g}=(X_{B_{v}}({\bf g}_{0})\cup Y)^{0,1}$ used instead (10).
It remains to assume that $u$ is a conjunction node. Let $u_{1}$ and $u_{2}$ be the children of $u$. For $i\in\{1,2\}$, let $B_{i}=B_{u_{i}}$, $V_{i}=\mathsf{var}(B)\cap\mathsf{var}(B_{i})$, ${\bf g}_{i}=Proj({\bf g},V_{i})$, $\pi_{i}=\pi[V_{i}]$. By Lemma 37, for each $i\in\{1,2\}$, $B_{i}$ respects $\pi_{i}$, ${\bf g}_{i}$ is an assignment over a prefix of $\pi_{i}$, and $\mathcal{S}(B_{i})|_{{\bf g}_{i}}\neq\emptyset$. Therefore, we can apply the induction assumption to each $B_{i}[{\bf g_{i}}]$.
Assume that both $L_{B_{1}}({\bf g}_{1})$ and $L_{B_{2}}({\bf g}_{2})$ are non-empty. Therefore, by Lemma 37 combined with the induction assumption, we conclude that
$\mathcal{S}(B)=\prod_{i\in\{1,2\}}(\prod_{w\in L_{B_{i}}({\bf g}_{i})}(\mathcal{S}(B_{w}))\times X_{B_{i}}({\bf g}_{i})^{0,1})$ (11)
The lemma is immediate by applying Corollary 3 and Corollary 4 to (11).
Assume that, say $L_{B_{1}}({\bf g}_{1})\neq\emptyset$ while $L_{B_{2}}({\bf g}_{2})=\emptyset$. Then the same reasoning applies with $\mathcal{S}(B)=\prod_{w\in L_{B_{1}}({\bf g}_{1})}(\mathcal{S}(B_{w}))\times X_{B_{1}}({\bf g}_{1})^{0,1}\times X_{B_{2}}({\bf g}_{2})$ being used instead (11). Finally, if $L_{B_{1}}({\bf g}_{1})=L_{B_{2}}({\bf g}_{2})=\emptyset$ then we use the same argumentation with $\mathcal{S}(B)=X_{B_{1}}({\bf g}_{1})\times X_{B_{2}}({\bf g}_{2})$ instead (11). ∎